### v3

# Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities

Abstract
We present new concentration inequalities for either martingale dependent or exchangeable random symmetric matrices under a variety of tail conditions, encompassing now-standard Chernoff bounds to self-normalized heavy-tailed settings. These inequalities are often randomized in a way that renders them strictly tighter than existing deterministic results in the literature, are typically expressed in the Loewner order, and are sometimes valid at arbitrary data-dependent stopping times. Along the way, we explore the theory of positive semidefinite supermartingales and maximal inequalities, a natural matrix analog of scalar nonnegative supermartingales that is potentially of independent interest.
1 Introduction
Bounding the tail probability of the sum of independent and identically distributed (i.i.d.) real-valued random variables is one of the most crucial tasks in the nonasymptotic theory of statistics, for which a copious collection of methods has flourished [Boucheron et al., 2013]. This paper concerns concentration inequalities for sums of random symmetric matrices under martingale dependence or exchangeability (both generalizing the i.i.d. assumption).
Notable previous results include the study of Rademacher and random graph adjacency matrices by Oliveira [2009, 2010]; those in a series of seminal papers by Tropp [2012, 2015], who famously ports the techniques of cumulant generating functions (CGFs) and Chernoff-type bounds, which are central in scalar concentration, into the matrix regime; some new bounds by Mackey et al. [2014] using exchangeability; and a recent study on “anytime-valid” concentration inequalities by Howard et al. [2020, 2021] which improves and unifies several matrix Chernoff-type bounds. We return to several of these works later in the current paper.
To set the stage for our exposition, we quote a celebrated result by Mackey et al. [2014, Corollary 4.2] (and also Tropp [2012, Theorem 1.3]) which states that, if $d\times d$ random i.i.d. symmetric matrices $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ satisfy $(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})^{2}\preceq\mathbf{B}$ (meaning that $\mathbf{B}-(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})^{2}$ is positive semidefinite) almost surely, then
$\mathbb{P}\left(\lambda_{\max}\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}-\mathbb{E}\mathbf{X}_{1}\right)\geqslant\sqrt{\frac{2\lambda_{\max}(\mathbf{B})\log(d/\alpha)}{n}}\right)\leqslant\alpha.$ (MHI)
Above, $\lambda_{\max}(\,\cdot\,)$ refers to the largest eigenvalue of a symmetric matrix, a typical scalarization of matrix-valued tails. When $d=1$, it recovers, with the same optimal constant, the scalar Hoeffding’s inequality as a special case; hence its name “matrix Hoeffding’s inequality”.
In this paper, we present numerous new techniques to strengthen this type of matrix concentration inequalities. Each of our new bounds enjoys a combination of some of the following virtues:
  1. 1.
    Anytime validity. We derive new inequalities that hold not only for a fixed number ($n$) of matrices but also a stopping time number of them [Howard et al., 2020].
  2. 2.
    Randomization. We losslessly tighten many of the previously known inequalities by randomly decreasing the tail threshold (decreasing the $a$ in (MHI) on the left hand side but not the right hand side), extending a technique that was recently explored by Ramdas and Manole [2023] for scalars.
  3. 3.
    Exchangeability. We go beyond the i.i.d. or martingale difference assumptions on random matrices and derive bounds that allow them to just be exchangeable instead.
  4. 4.
    Tighter (or no) scalarization. In some bounds we use the trace functional $\mathrm{tr}(\,\cdot\,)$ which is larger than the usual $\lambda_{\max}(\,\cdot\,)$ (e.g. in (MHI)), and introduce some bounds completely in the matrix order $\preceq$ without any scalarization, which is partly achieved by studying the maximal inequalities of matrix-valued sub- and super-martingales.
  5. 5.
    Heavy tails. Instead of assuming the matrices to be bounded or (sub)Gaussian, we sometimes allow them to only have a finite variance or $p$th moment, or be symmetric.
To showcase these virtues, let us pick two of the bounds we shall prove in the exposition that follows. Under the assumptions of the (MHI) above, our Theorem 5.5 states that
$\mathbb{P}\left(\lambda_{\max}\left(\frac{\gamma_{1}\mathbf{X}_{1}+\dots+\gamma_{\tau}\mathbf{X}_{\tau}}{\gamma_{1}+\dots+\gamma_{\tau}}-\mathbb{E}\mathbf{X}_{1}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{2}\lambda_{\max}(\mathbf{B})\sum_{i=1}^{\tau}\gamma_{i}^{2}}{\gamma_{1}+\dots+\gamma_{\tau}}\right)\leqslant\alpha$ (1)
for any predictable positive tuning parameters $\{\gamma_{n}\}$, stopping time $\tau$, and an independent random variable $U\sim\text{Unif}_{(0,1)}$. The above equips (MHI) with virtues 1 and 2 without compromises, since replacing $U$ with 1 and taking $\tau=n$, $\gamma_{1}=\dots=\gamma_{n}=\sqrt{\frac{2\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B})}}$ in above, one obtains the exact deviation bound $\sqrt{\frac{2\log(d/\alpha)\lambda_{\max}(\mathbf{B})}{n}}$ as in (MHI). A bound of very different flavor is that of our Theorem 4.17, which states that for exchangeable matrices $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ with variance $\mathbf{V}$,
$\mathbb{P}\left(\exists n\geqslant 1:\operatorname{abs}\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}-\mathbb{E}\mathbf{X}_{1}\right)\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ (2)
for any positive definite $\mathbf{A}$. This bound enjoys virtues 3, 4, and 5, sequentializing the matrix Chebyshev inequality $\mathbb{P}\left(\operatorname{abs}\left(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1}\right)\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ by Ahlswede and Winter [2002, Appendix C] to an infinite time horizon.
Outline.
The rest of the paper is organized as follows. In Section 2, we briefly review the technical background (and challenges) of matrix concentration bounds as well as a collection of classical and recent scalar bounds that anticipate our matrix extensions. We present in Section 3 our first kind of bounds, derived via randomized matrix Markov, Chebyshev, and Chernoff inequalities. We develop and apply the theory of matrix-valued supermartingales and backward submartingales in Section 4, further obtaining bounds for i.i.d. and exchangeable matrices. Finally, in Section 5, we revisit scalar supermartingales (and the scalar processes they dominate) constructed via matrices which yield some other randomized bounds as well, concluding with a discussion on the respective advantages of matrix- and scalar- based sequential tests in both theoretical and empirical dimensions.
2 Preliminaries
2.1 Matrix Analysis
Let $\mathcal{S}_{d}$ denote the set of all $d\times d$ real-valued symmetric matrices, which is the only class of matrices considered in this paper. [^1] We only briefly discuss the case with non-symmetric and non-square matrices in Appendix D. These matrices are denoted by bold upper-case letters $\mathbf{A},\mathbf{B}$, etc; whereas $d$ dimensional (column) vectors by bold lower-case letters $\mathbf{u},\mathbf{v}$, etc. For $I\subseteq\mathbb{R}$, we denote by $\mathcal{S}_{d}^{I}$ the set of all real symmetric matrices whose eigenvalues are all in $I$. $\mathcal{S}_{d}^{[0,\infty)}$, the set of positive semidefinite and $\mathcal{S}_{d}^{(0,\infty)}$, the set of positive definite matrices are simply denoted by $\mathcal{S}_{d}^{+}$ and $\mathcal{S}_{d}^{++}$ respectively. The Loewner partial order is denoted $\preceq$, where $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite, and $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite.
Any function $f:I\to J$ (where $I,J\subseteq\mathbb{R}$) is identified canonically with a function $f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$ in the following way: for a diagonal matrix $\mathbf{\Sigma}\in\mathcal{S}_{d}^{I}$, $f(\mathbf{\Sigma})$ is obtained by applying $f$ entrywise; otherwise it is obtained by applying $f$ after orthogonal diagonalization, $f(\mathbf{U}^{\mathsf{T}}\mathbf{\Sigma}\mathbf{U})=\mathbf{U}^{\mathsf{T}}f(\mathbf{\Sigma})\mathbf{U}$. Notable matrix-to-matrix operators that recur in this paper defined this way include the matrix exponential $\exp:\mathcal{S}_{d}\to\mathcal{S}^{++}_{d}$, the matrix logarithm $\log:\mathcal{S}^{++}_{d}\to\mathcal{S}_{d}$, the matrix absolute value [^2] We use $\operatorname{abs}(\,\cdot\,)$ instead of ||$\cdot$|| ⋅ | to distinguish it from the determinant and other norms. $\operatorname{abs}:\mathcal{S}_{d}\to\mathcal{S}^{+}_{d}$, and powers $\mathbf{X}\mapsto\mathbf{X}^{k}$ for any $k\in\mathbb{R}$. Take heed that usually $\exp(\mathbf{A}+\mathbf{B})\neq\exp\mathbf{A}+\exp\mathbf{B}$ and $\log\mathbf{A}\mathbf{B}\neq\log\mathbf{A}+\log\mathbf{B}$, but a trace equality holds on the latter:
Fact 1 (Trace Logarithm Identity).
$\mathrm{tr}\log\mathbf{A}\mathbf{B}=\mathrm{tr}\log\mathbf{A}+\mathrm{tr}\log\mathbf{B}$.
[…]
It is rare to find matrix-to-matrix functions that are both monotone (i.e. preserving the $\preceq$ order) and convex (i.e. $f$ such that $f(\sum\lambda_{i}\mathbf{X}_{i})\preceq\sum{\lambda_{i}}f(\mathbf{X}_{i})$ if $\sum\lambda_{i}=1$ and each $\lambda_{i}\geqslant 0$; concavity is defined similarly). Scalar-to-scalar monotonicity ($I\to J$) does not imply matrix-to-matrix monotonicity ($\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$), and the same holds for convexity. For example, the matrix square $\mathbf{X}\mapsto\mathbf{X}^{2}$ is not monotone on the domain $\mathcal{S}_{d}^{+}$ and the matrix 4th power $\mathbf{X}\mapsto\mathbf{X}^{4}$ is neither monotone nor convex. The following are some of the few well-known nontrivial “operator-monotone” and “operator-convex” functions in the matrix analysis literature (cf. Bhatia [2013, Chapter V]).
Fact 2 (Operator-monotone and -concave Functions).
The matrix logarithm $\mathbf{X}\mapsto\log\mathbf{X}$ ($\mathcal{S}_{d}^{++}\to\mathcal{S}_{d}$) is monotone and concave. For $k\in(0,1)$, the matrix power $\mathbf{X}\mapsto\mathbf{X}^{k}$ ($\mathcal{S}_{d}^{+}\to\mathcal{S}_{d}^{+}$) is monotone and concave. For $k\in(1,2]$, the matrix power $\mathbf{X}\mapsto\mathbf{X}^{k}$ ($\mathcal{S}_{d}^{+}\to\mathcal{S}_{d}^{+}$ if $k\neq 2$; $\mathcal{S}_{d}\to\mathcal{S}_{d}^{+}$ if $k=2$) is convex.
[…]
Fact 3 (Trace Monotonicity).
Let $f:I\to\mathbb{R}$ be monotone increasing ($I\subseteq\mathbb{R}$). Then so is $\mathrm{tr}\circ f:\mathcal{S}_{d}^{I}\to\mathbb{R}$, i.e.,
$\mathbf{A}\preceq\mathbf{B}\implies\mathrm{tr}f(\mathbf{A})\leqslant\mathrm{tr}f(\mathbf{B}).$ (3)
Fact 4 (Trace Jensen).
For an interval $I\subseteq\mathbb{R}$, assume $f:I\to\mathbb{R}$ is convex. Then so is $\mathrm{tr}\circ f:\mathcal{S}_{d}^{I}\to\mathbb{R}$, i.e.,
$\mathrm{tr}f\left(\sum\lambda_{i}\mathbf{X}_{i}\right)\leqslant\mathrm{tr}\left(\sum\lambda_{i}f(\mathbf{X}_{i})\right),$ (4)
if $\sum\lambda_{i}=1$ and each $\lambda_{i}\geqslant 0$.
2.2 Concentration of Scalars
We briefly review some classical and recent scalar tail bounds in this subsection. Markov inequality, the foundation of a large family of scalar concentration inequalities, states that if $X$ is a nonnegative random variable,
$\mathbb{P}(X\geqslant a)\leqslant(\mathbb{E}X)a^{-1}$ (MI)
for any $a>0$. There have been many generalizations of (MI). Ville’s inequality [Ville, 1939] states that a similar tail bound holds uniformly over time for a nonnegative supermartingale $\{M_{n}\}$,
$\mathbb{P}(\exists n,\ M_{n}\geqslant a)\leqslant(\mathbb{E}M_{0})a^{-1},$ (VI)
or equivalently, $\mathbb{P}(M_{\tau}\geqslant a)\leqslant(\mathbb{E}M_{0})a^{-1}$ for any stopping time $\tau$ [Howard et al., 2020]. Ville’s inequality (VI) has been the workhorse for many recent time-uniform concentration inequalities (“confidence sequences”) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b] that have been favored by, for example, the multi-armed bandit literature.
Applying (MI) to the square of the sum of i.i.d. random variables $X_{1},\dots,X_{n}$ of mean $\mu$ and variance $\sigma^{2}$, we have the famous Chebyshev inequality
$\mathbb{P}\left(\left|\overline{X}_{n}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{na^{2}},$ (CI)
where $\overline{X}_{n}$ is the arithmetic mean of $X_{1},\dots,X_{n}$. A perhaps less-known extension of (CI) is that if $X_{1},\dots,X_{n}$ have central $p$th moment $v_{p}$ ($1\leqslant p\leqslant 2$) instead of variance,
$\mathbb{P}\left(\left|\overline{X}_{n}-\mu\right|\geqslant a\right)\leqslant\frac{2^{2-p}v_{p}}{n^{p-1}a^{p}},$ (5)
which we prove in Section A.1 for completeness. In the other direction, if $X_{i}$’s are light-tailed and have moment generating function (MGF) $G(\gamma)=\mathbb{E}\mathrm{e}^{\gamma(X_{1}-\mu)}$, (MI) applied to the exponential of their mean yields the Chernoff-Hoeffding inequality
$\mathbb{P}(\overline{X}_{n}-\mu\geqslant a)\leqslant G(\gamma/n)^{n}\mathrm{e}^{-\gamma a}.$ (6)
Recently, Ramdas and Manole [2023] discovered some interesting strengthened forms of the bounds above. First, uniform randomization. It is found that if $U$ is a $\mathrm{Unif}_{(0,1)}$-distributed random variable independent from $X$ or $\{M_{n}\}$, (MI) and (VI) and be losslessly tightened into
$\mathbb{P}(X\geqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (UMI)
and
$\mathbb{P}(M_{\tau}\geqslant Ua)\leqslant(\mathbb{E}M_{0})a^{-1}$ (UVI)
respectively. The downstream bounds including (CI), (5) and (6) can thus also be randomized into strictly tighter forms. Second, Ramdas and Manole [2023] introduce exchangeable generalizations of (CI) which partly shares the anytime virtue of (VI)-type bounds, stating that as long as the infinite sequence $X_{1},X_{2},\dots$ is exchangeable,
$\mathbb{P}\left(\exists n,\ \left|\overline{X}_{m}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{a^{2}};$ (XCI)
and further, if i.i.d., (XCI) implies that for any fixed $N$,
$\mathbb{P}\left(\exists n\geqslant N,\ \left|\overline{X}_{m}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{Na^{2}}.$ (7)
Extensions of (MI) and (VI) to random symmetric matrices are proved by Ahlswede and Winter [2002, Appendix C]; of (6) by Tropp [2012, 2015]. We will bring up these previous extensions again soon in the following. Our current paper is partly animated by the question if these novel bounds (UMI), (UVI), (XCI), and (7) by Ramdas and Manole [2023] have matrix siblings as well; and if yes, to what extent of generality. We shall show that there is indeed a plethora of matrix tail bounds that recover as $d=1$ special cases the results mentioned above, using classical matrix analysis tools plus some new probability techniques.
2.3 Templates of Our Bounds
We now sketch out the skeletons that our bounds may have, and discuss their relation. Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random symmetric matrices whose common distribution belongs to a family $\mathcal{P}$. Denote by $\mathbf{M}$ their common mean and $\overline{\mathbf{X}}_{n}$ their empirical average $\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$. The following are the concentration events that we may prove to be of high probability. $\mathbf{A}\in\mathcal{S}^{+}_{d}$ denotes a constant matrix of deviation, and $r(\cdot)$ is a “randomization” map producing an $r(\mathbf{A})$ that is random but always smaller than $\mathbf{A}$ in Loewner order (the $Ua$ term in (UMI) is a one-dimensional example of $r(a)$). An arrow $A\to B$ denote the inclusion of events $A\subseteq B$, and hence also that $A$ is a stronger claim than $B$.
{tikzcd}
Most results in the past, for example (MHI), have been spectral bounds, which are weaker than many of the bounds we shall present in this paper.
3 Randomized Fixed-Time Matrix Tail Bounds
3.1 Randomized Matrix Markov Inequality
We begin our exposition with a matrix version of uniformly randomized Markov inequality (UMI) by Ramdas and Manole [2023]; a randomized version of the matrix Markov inequality by Ahlswede and Winter [2002, Theorem 12],
$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}),$ (MMI)
where $\mathbf{X}$ is an $\mathcal{S}_{d}^{+}$-valued random matrix and $\mathbf{A}\in\mathcal{S}_{d}^{++}$ is deterministic.
To achieve this, let us first define what it means for a random matrix to be uniformly and super-uniformly distributed. Note that it is easy to see that the randomizer $U$ in (UMI) and (UVI) can be replaced by any random variable that is stochastically larger; that is, one satisfying the following definition:
Definition 3.1.
We call a positive random scalar $U$ super-uniform if $\mathbb{P}(U\leqslant x)\leqslant x$ for all $x\geqslant 0$; that is, $U$ is stochastically greater than or equal to the uniform distribution $\operatorname{Unif}_{(0,1)}$.
This has two viable extensions to matrices, one in terms of the trace and the other of the spectral norm.
Definition 3.2.
We call an $\mathcal{S}_{d}^{+}$-valued random matrix $\mathbf{U}$ spectrally super-uniform if $\mathbb{P}(\mathbf{U}\nsucceq\mathbf{Y})\leqslant\lambda_{\max}(\mathbf{Y})$ for all $\mathbf{Y}\in\mathcal{S}_{d}^{+}$, and $\mathcal{Y}$-spectrally uniform if, further, equality holds for all $\mathbf{Y}\in\mathcal{Y}\subseteq\mathcal{S}_{d}^{+}$. We say $\mathbf{U}$ is trace super-uniform if $\mathbb{P}(\mathbf{U}\nsucceq\mathbf{Y})\leqslant\mathrm{tr}\mathbf{Y}$ for all $\mathbf{Y}\in\mathcal{S}_{d}^{+}$, and $\mathcal{Y}$-trace uniform if, further, equality holds for all $\mathbf{Y}\in\mathcal{Y}\subseteq\mathcal{S}_{d}^{+}$.
Note that every spectrally super-uniform matrix is a trace super-uniform matrix since $\lambda_{\max}(\mathbf{Y})\leqslant\mathrm{tr}\mathbf{Y}$. If $U$ is a super-uniform random scalar, then $U\mathbf{I}$ is a spectrally (hence trace) super-uniform random matrix, because $\mathbb{P}(U\mathbf{I}\nsucceq\mathbf{Y})=\mathbb{P}(U\leqslant\lambda_{\max}(\mathbf{Y}))\leqslant\lambda_{\max}(\mathbf{Y}).$
Straightforwardly the identity matrix $\mathbf{I}$ is spectrally (hence trace) super-uniform. If $U\sim\operatorname{Unif}_{(0,1)}$, $U\mathbf{I}$ is $\mathcal{S}_{d}^{[0,1]}$-spectrally uniform, and $\left\{\mathbf{v}\mathbf{v}^{\mathsf{T}}:\|\mathbf{v}\|\leqslant 1\right\}$-trace uniform. Having defined trace super-uniformity, we can now state a matrix extension of the uniformly randomized Markov’s inequality (UMI).
Theorem 3.3 (Uniformly Randomized Matrix Markov Inequality).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{X}\subseteq\mathcal{S}_{d}^{+}$, and $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}).$ (UMMI)
The above holds with equality if and only if $\mathbf{U}$ is $\mathbf{A}^{-1/2}\mathcal{X}\mathbf{A}^{-1/2}$-trace uniform.
Proof.
By our assumptions,
$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})=\mathbb{P}(\mathbf{A}^{-1/2}\mathbf{X}\mathbf{A}^{-1/2}\npreceq\mathbf{U})=\mathbb{E}(\mathbb{P}(\mathbf{A}^{-1/2}\mathbf{X}\mathbf{A}^{-1/2}\npreceq\mathbf{U}|\mathbf{X}))$ (8)
[…]
concluding the proof. ∎
[…]
$\mathbb{P}(\lambda_{\max}(\mathbf{X})>\mathrm{tr}(\mathbf{A}\mathbf{U}))\leqslant\mathbb{P}(\mathrm{tr}\mathbf{X}>\mathrm{tr}(\mathbf{A}\mathbf{U}))\leqslant\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2}),$ (10)
subject to the same upper bound $\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1})$.
The following is an example where the inequality of (UMMI) holds with equality. We take $\mathbf{U}=U\mathbf{I}$ where $U\sim\operatorname{Unif}_{(0,1)}$, and $\mathbf{A}\in\mathcal{S}_{d}^{++}$. Let $\mathbf{x}$ be a random vector in the ellipsoid $\{\mathbf{x}:\mathbf{x}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{x}\leqslant 1\}\subseteq\mathbb{R}^{d}$. Then,
$\mathbb{P}(\|\mathbf{A}^{-1/2}\mathbf{x}\|^{2}\leqslant U)=\mathbb{P}(\mathbf{x}\mathbf{x}^{\mathsf{T}}\npreceq U\mathbf{A})=\mathrm{tr}(\mathbf{A}^{-1}\mathbb{E}(\mathbf{x}\mathbf{x}^{\mathsf{T}}))=\mathbb{E}\|\mathbf{A}^{-1/2}\mathbf{x}\|^{2},$ (11)
where the second equality above is (UMMI) applied on the random matrix $\mathbf{x}\mathbf{x}^{\mathsf{T}}$. It is not hard to see here the condition for equality is met. Intuitively, this hints that (UMMI) is tighter when the random matrix $\mathbf{X}$ is approximately of lower rank, a point that we shall return to later in this paper with statistical examples.
Concluding this section, we remark on trace super-uniformity that clearly $U\mathbf{I}+\mathbf{Y}$ is trace super-uniform when $U\sim\operatorname{Unif}_{(0,1)}$ and $\mathbf{Y}\succeq 0$, and this provides an interesting class for our purposes. It may be interesting to fully characterize the set of all trace super-uniform matrices, but we leave it as future work.
3.2 Randomized Matrix Chebyshev Inequalities
Recall that the variance of a random symmetric matrix $\mathbf{X}$ is defined simply as $\mathbf{Var}(\mathbf{X})=\mathbb{E}(\mathbf{X}-\mathbb{E}\mathbf{X})^{2}$. Applying Theorem 3.3 to the square of one square-integrable matrix and the square of the sum of several i.i.d. square-integrable matrices, we have the following randomization of the one-matrix Chebyshev inequality by Ahlswede and Winter [2002, Theorem 14],
$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2})$ (12)
and the implied law of large numbers for the average $\overline{\mathbf{X}}_{n}$ of $n$ i.i.d. symmetric matrices with mean $\mathbf{M}$ and variance $\mathbf{V}$ by Ahlswede and Winter [2002, Theorem 15],
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (MCI)
Lemma 3.4 (Uniformly Randomized One-Matrix Chebyshev Inequality).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$, $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2}).$ (13)
Proof.
By Fact 2, $\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant\mathbb{P}((\mathbf{X}-\mathbb{E}\mathbf{X})^{2}\npreceq\mathbf{A}\mathbf{U}\mathbf{A})$. By Theorem 3.3, we have $\mathbb{P}((\mathbf{X}-\mathbb{E}\mathbf{X})^{2}\npreceq\mathbf{A}\mathbf{U}\mathbf{A})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2})$. ∎
We can immediately apply Lemma 3.4 to multiple matrices, obtaining the following weak-law-of-large-numbers style inequality that shrinks when $n$ gets larger. Recall that the anticommutator between two sqaure matrices $\mathbf{A},\mathbf{B}$ is defined as $\{\mathbf{A},\mathbf{B}\}=\mathbf{A}\mathbf{B}+\mathbf{B}\mathbf{A}$.
Theorem 3.5 (Uniformly Randomized $n$-Matrix Chebyshev Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{i}=\mathbf{M}$ and variance matrix upper bound $\mathbf{Var}(\mathbf{X}_{i})\preceq\mathbf{V}$, such that $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$; $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from them all. Then, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$, we have
$\mathbf{Var}(\overline{\mathbf{X}}_{n})\preceq n^{-1}\mathbf{V}.$ (14)
Consequently, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$.
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}),$ (UMCI)
Proof.
Note that
$\small\mathbf{Var}(\overline{\mathbf{X}}_{n})=n^{-2}\mathbb{E}\left(\sum_{i=1}^{n}(\mathbf{X}_{i}-\mathbf{M})\right)^{2}=n^{-2}\left(\sum_{i=1}^{n}\mathbf{Var}(\mathbf{X}_{i})+\sum_{i<j}\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}\right)\preceq n^{-1}\mathbf{V},$ (15)
and apply Lemma 3.4 to conclude the proof. ∎
It is easy to see that (UMCI) recovers as special cases (MCI) and (CI).
Remark 3.6.
The assumption $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for $i\neq j$ is satisfied when $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ are pairwise uncorrelated, which is in turn satisfied under martingale difference assumption (i.e., $\{\mathbf{X}_{i}\}$ is adapted to some filtration $\{\mathcal{F}_{i}\}$ with $\mathbb{E}(\mathbf{X}_{i}|\mathcal{F}_{i-1})=\mathbf{M}$ and $\mathbf{Var}(\mathbf{X}_{i}|\mathcal{F}_{i-1})\preceq\mathbf{V}$), which is in turn implied if they are i.i.d.
For $p\in[1,2)$, a $p$th moment bound can lead to an infinite variance version of (13), whose non-randomized ($\mathbf{U}=\mathbf{I}$) version is mentioned by Tu [2017].
Lemma 3.7 (Uniformly Randomized One-Matrix $p$-Chebyshev Inequality).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$ that has a finite $p$th central moment $\mathbf{V}_{p}:=\mathbb{E}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}$ and $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (16)
Proof.
By Fact 2, $\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p})\leqslant\mathbb{P}((\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}\npreceq\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})$. By Theorem 3.3, we have $\mathbb{P}((\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}\npreceq\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p})$. ∎
However, an $L^{p}$ analogy of (UMCI) (i.e. a matrix version of the $p$-Chebyshev inequality (5)) is not straightforward. This is largely because while the contraction of variance rule (14) exists for matrices, we do not know of any matrix version of $p$th moment contraction rule. In fact, the scalar $p$th moment contraction rule was only recently established (by Cherapanamjeri et al. [2022, Lemma 4.2] and Wang et al. [2021, Lemma 7], see the proof of (5) in Section A.1). Section A.2 contains a conjectured result along these lines, as well as a looser result that is the best we can currently achieve.
3.3 Randomized Matrix Chernoff Bounds
We now establish the classical Chernoff-type bounds with randomization for light-tailed random matrices that have finite moment generating functions (MGFs).
Lemma 3.8 (Uniformly Randomized One-Matrix Chernoff Bound).
Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$, $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$. For any $\mathbf{A}\in\mathcal{S}_{d}$ and real number $\gamma$,
$\mathbb{P}\left(\mathbf{X}\npreceq\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right)\leqslant\mathrm{tr}(\mathrm{e}^{-2\gamma\mathbf{A}}\,\mathbb{E}\mathrm{e}^{2\gamma\mathbf{X}}).$ (17)
Proof.
By Fact 2, we have $\mathbb{P}\left(2\gamma\mathbf{X}\npreceq\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right)\leqslant\mathbb{P}(\mathrm{e}^{2\gamma\mathbf{X}}\npreceq\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})$. By Theorem 3.3, we have $\mathbb{P}(\mathrm{e}^{2\gamma\mathbf{X}}\npreceq\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\leqslant\mathrm{tr}(\mathrm{e}^{-2\gamma\mathbf{A}}\,\mathbb{E}\mathrm{e}^{2\gamma\mathbf{X}})$, concluding the proof. ∎
The seemingly indigestible tail threshold $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})$ in (17) can be largely simplified in any of the following scenarios. First, one can simply take the trace on the tail event above, obtaining due to Fact 1 and the monotonicity of the trace,
$\mathbb{P}\left(\mathrm{tr}\mathbf{X}>\mathrm{tr}\mathbf{A}+\frac{1}{2\gamma}\mathrm{tr}\log\mathbf{U}\right)\leqslant\mathbb{P}\left(\mathbf{X}\npreceq\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right),$ (18)
without any additional assumptions; Second, if $\mathbf{U}=U\mathbf{I}$ where $U$ is scalar super-uniform, $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})=\mathbf{A}+\frac{1}{2\gamma}\mathbf{I}\log U$; Third, if $\mathbf{A}=a\mathbf{I}$, $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})=a\mathbf{I}+\frac{1}{2\gamma}\log\mathbf{U}$.
It is known that, unlike the scalar case, the one-matrix MGF is not linearly additive. Once again, the complication is partially resolved when the trace is applied.
Fact 5 (Lemma 3.4 of Tropp [2012]).
Let $\mathbf{Z}_{1},\dots,\mathbf{Z}_{n}$ be independent random matrices taking values in $\mathcal{S}_{d}$. Then,
$\mathrm{tr}\left\{\mathbb{E}\exp\left(\sum\mathbf{Z}_{i}\right)\right\}\leqslant\mathrm{tr}\exp\left(\sum\log\mathbb{E}\mathrm{e}^{\mathbf{Z}_{i}}\right).$ (19)
This implies the following bound that recovers as special cases both the scalar Chernoff-Hoeffding inequality (6) when taking $d=\mathbf{U}=1$, and the “master tail bound” on $\lambda_{\max}(\overline{\mathbf{X}}_{n}-\mathbf{M})$ by Tropp [2012, Theorem 3.6] when taking $\mathbf{U}=\mathbf{I}$.
Theorem 3.9 (Uniformly Randomized Matrix Chernoff-Hoeffding Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and matrix-valued MGF $\mathbf{G}(\gamma)=\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}$; $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from them all. Let $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$. For any $a,t>0$,
$\mathbb{P}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\npreceq a\mathbf{I}+\frac{\log\mathbf{U}}{\gamma}\right)\leqslant\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})\mathrm{e}^{-\gamma a}.$ (20)
Proof.
By Fact 5, $\mathrm{tr}(\mathbb{E}\mathrm{e}^{\gamma(\overline{\mathbf{X}}_{n}-\mathbf{M})})\leqslant\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})$. Using Lemma 3.8 on $\gamma(\overline{\mathbf{X}}_{n}-\mathbf{M})$ with $\mathbf{A}=a\mathbf{I}$ concludes the proof. ∎
Again, the bound remains valid if only an upper bound of the matrix MGF $\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}\preceq\mathbf{G}(\gamma)$ or (more commonly) the weaker log-MGF bound $\log\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}\preceq\log\mathbf{G}(\gamma)$ is known. Table 1 contains some important known matrix (log-)MGFs (see Tropp [2012, 2015]). An asymmetric Hoeffding bound is missing here but shall be discussed in Section 5 later since it requires some further techniques beyond the matrix log-MGF.
Name Condition $\mathbf{G}(\gamma)$ $\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})$
Rademacher $\mathbf{X}_{1}-\mathbf{M}=R_{1}\mathbf{C},R_{1}\sim\operatorname{rad}_{1/2}$ $\preceq\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$ $\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$
Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $=\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$ $=\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$
Bennett I $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})=\mathbf{V}$ $\preceq\mathrm{e}^{(\mathrm{e}^{\gamma}-\gamma-1)\mathbf{V}}$ $\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$
Bennett II $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})\preceq\mathbf{V}$ $\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$
Symmetric Hoeffding $(\mathbf{X}_{1}-\mathbf{M})^{2}\preceq\mathbf{B},\ \mathbf{X}_{1}-\mathbf{M}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}-\mathbf{X}_{1}$ $\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{B}/2n})$
Table 1: Some important matrix MGFs and log-MGFs. Here $\mathbf{B},\mathbf{C},\mathbf{V}$ are deterministic matrices, and =$\stackrel{{\scriptstyle d}}{{=}}$ means two random matrices having the same distribution. The proofs are in Section B.1.
4 Maximal Inequalities of Matrix (Super-/Sub-)Martingales
Some interesting tail bounds are obtained if we consider matrix-valued supermartingales and submartingales and their maximal inequalities. These processes can be defined on forward or backward filtrations and in both these cases, well-known scalar-valued maximal inequalities can be extended to matrices.
4.1 Matrix-Valued Supermartingales
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued stochastic process adapted to the filtration $\{\mathcal{F}_{n}\}$. [^3] We shall often omit the filtration $\{\mathcal{F}_{n}\}$ when unambiguous from the context. We say that $\{\mathbf{Y}_{n}\}$ is a supermartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}$ holds for every $n$. Before we proceed, note that the order $\preceq$ is preserved by addition, limit, and Lebesgue integration (hence including convergent infinite sums, expectations, conditional expectations, etc., essentially because $\mathcal{S}^{+}=\{\mathbf{A}:\mathbf{A}\succeq 0\}$ is closed both topologically and under $+$). See also a “dominated integrability” lemma we prove in Section B.2.
Matrix supermartingales are related to scalar supermartingales in the following ways. First, the monotonicity of trace implies that if $\{\mathbf{Y}_{n}\}$ is an $\mathcal{S}_{d}$-valued supermartingale, its trace process $\{\mathrm{tr}(\mathbf{Y}_{n})\}$ is a real-valued supermartingale. Second, we have the following lemma.
Lemma 4.1.
An $\mathcal{S}_{d}$-valued process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every deterministic vector $\mathbf{v}\in\mathbb{R}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale.
Proof.
$\{\mathbf{Y}_{n}\}$ is a supermartingale (21)
$\iff$ For all $n$, $\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq 0$ (22)
$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}(\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1}))\mathbf{v}\geqslant 0$ (23)
$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n-1}\mathbf{v}-\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}|\mathcal{F}_{n-1})\geqslant 0$ (24)
$\iff$ For all $\mathbf{v}$, $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale. (25)
This completes the proof. ∎
This leads to an analog of Doob’s martingale convergence theorem for positive semidefinite matrix supermartingales.
Theorem 4.2 (Matrix Supermartingale Convergence Theorem).
Any $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ converges almost surely to an $\mathcal{S}_{d}^{+}$-valued random matrix $\mathbf{Y}_{\infty}$ and $\mathbb{E}\mathbf{Y}_{\infty}\preceq\mathbb{E}\mathbf{Y}_{0}$.
Proof.
For any $\mathbf{v}\in\mathbb{R}^{d}$, the process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a scalar-valued nonnegative supermartingale, which converges a.s. to an integrable random scalar due to Doob’s martingale convergence theorem (see e.g. Corollary 11.5 in Klenke [2013]). First, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that $\mathbf{Y}_{n}$’s diagonal converges to integrable scalars. Then, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that the rest of $\mathbf{Y}_{n}$’s entries also converge to integrable scalars. The limit $\mathbf{Y}_{\infty}=\lim_{n\to\infty}\mathbf{Y}_{n}$ is in $\mathcal{S}_{d}^{+}$ because $\mathcal{S}_{d}^{+}$ is a closed set. Finally, for any $\mathbf{v}\in\mathbb{R}^{d}$,
$\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{\infty})\mathbf{v}=\lim_{n\to\infty}\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{0}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{0})\mathbf{v},$ (26)
so $\mathbb{E}\mathbf{Y}_{\infty}\preceq\mathbb{E}\mathbf{Y}_{0}$. ∎
Next, we see that using a similar trick, any scalar version of the optional stopping (sampling) theorems extends to symmetric matrices as well. For example:
Lemma 4.3 (Matrix Optional Stopping).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Then,
  1. 1.
    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale on $\{\mathcal{F}_{n}\}$.
  2. 2.
    If $\tau\leqslant N$ almost surely for some $N$, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
  3. 3.
    If $\{\mathbf{Y}_{n}\}$ is $\mathcal{S}_{d}^{+}$-valued, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
Proof.
1. For any $\mathbf{v}\in\mathbb{R}^{d}$, the scalar process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale on $\{\mathcal{F}_{n}\}$ and therefore so is the stopped process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n\wedge\tau}\mathbf{v}\}$ due to the optional stopping theorem (e.g. Theorem 10.15 of Klenke [2013]). This implies that $\{\mathbf{Y}_{n\wedge\tau}\}$ is a matrix supermartingale on $\{\mathcal{F}_{n}\}$. 2. For any $\mathbf{v}\in\mathbb{R}^{d}$, we apply the optional sampling theorem (e.g. Theorem 10.11(i) of Klenke [2013]) to the scalar-valued supermartingale $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ to obtain that $\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{v}=\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\tau}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{0}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{0})\mathbf{v}$. Therefore $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. 3. Apply Theorem 4.2 to the $\mathcal{S}_{d}^{+}$-valued stopped supermartingale $\{\mathbf{Y}_{\tau\wedge n}\}$, we have $\mathbb{E}\mathbf{Y}_{\tau\wedge\infty}\preceq\mathbb{E}\mathbf{Y}_{\tau\wedge 0}$, i.e., $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. ∎
This gives rise to the matrix versions of (VI) and (UVI) via some manipulation of crossing events.
Theorem 4.4 ((Uniformly Randomized) Matrix Ville’s Inequality).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale on $\{\mathcal{F}_{n}\}$, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Let $\mathbf{U}$ be a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathcal{F}_{\infty}$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (UMVI)
Consequently,
$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
Proof.
First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Lemma 4.3. By Theorem 3.3 and the monotinicity of the trace,
$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1/2}),$ (27)
which is (UMVI). Next, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$. Setting $\mathbf{U}$ to $\mathbf{I}$ and $\tau$ to $\nu\wedge N$ in (UMVI), we have
$\mathbb{P}(\nu\leqslant N)=\mathbb{P}(\mathbf{Y}_{\nu\wedge N}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1})$ (28)
for any $N\in\mathbb{N}$. Letting $N\to\infty$, we have
$\mathbb{P}(\nu<\infty)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1})$ (29)
which concludes the proof of (MVI). ∎
We remark that while there are two equivalent ways of stating the Ville’s inequality, the “any-time” statement (VI) and the stopping time statement (stated after (VI)), it appears that randomization only applies to the latter [Ramdas and Manole, 2023]. The same applies to our matrix versions of Ville’s inequality. Finally, we remark that some other maximal inequalities for scalar sub- and super-martingales can be extended to matrices as well, of which we derive a few in Appendix C.
4.2 Matrix Generalizations of Classical Test Supermartingales
Recall that scalar-valued nonnegative supermartingales are often constructed incrementally by multiplying each time a nonnegative variable with conditional expectation $\leqslant 1$. Analogously, we may use the following straightforward fact to construct $\mathcal{S}_{d}^{+}$-valued supermartingales.
Lemma 4.5.
Let $\mathbf{E}_{n}$ be $\mathcal{S}_{d}^{+}$-valued and $\mathcal{F}_{n}$-measurable, $\mathbf{F}_{n}$ be $\mathcal{S}_{d}^{++}$-valued and $\mathcal{F}_{n-1}$-measurable, such that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ almost surely. Then, the following $\mathcal{S}_{d}^{+}$-valued process is a supermartingale on $\{\mathcal{F}_{n}\}$:
$\mathbf{Y}_{n}=\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}.$ (30)
In some cases, we can take $\mathbf{F}_{1}=\mathbf{F}_{2}=\dots=\mathbf{I}$, and thus $\mathbf{Y}_{n}=\sqrt{\mathbf{E}_{1}}\dots\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\dots\sqrt{\mathbf{E}_{1}}$.
Proof.
Note that
$\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})=\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\mathbb{E}({\mathbf{E}_{n}}|\mathcal{F}_{n-1})\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}$ (31)
$\preceq$ $\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\mathbf{F}_{n}^{-1}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}=\mathbf{Y}_{n-1},$ (32)
concluding the proof. ∎
We remark that by defining $\mathbf{E}_{n}^{\prime}=\sqrt{\mathbf{F}_{n}}\mathbf{E}_{n}\sqrt{\mathbf{F}_{n}}$ and $\mathbf{F}_{n}^{\prime}=\mathbf{I}$, we can obtain another $\mathcal{S}_{d}^{+}$-valued supermartingale $\sqrt{\mathbf{E}_{1}^{\prime}}\dots\sqrt{\mathbf{E}_{n}^{\prime}}\sqrt{\mathbf{E}_{n}^{\prime}}\dots\sqrt{\mathbf{E}_{1}^{\prime}}$. This option, however, involves more computation than (30) which can be computed incrementally by computing the two matrices $\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}$ and $\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}$ on the fly.
With Lemma 4.5, we can recover matrix versions of several recent “test supermartingales” (nonnegative supermartingales with initial value one) in the literature. In the following examples, we will state some conditions on a sequence of $\mathcal{S}_{d}$-valued random matrices $\{\mathbf{X}_{n}\}$ that will lead to a sequence of $\{\mathbf{E}_{n}\}$ and $\{\mathbf{F}_{n}\}$ that satisfy Lemma 4.5, so that a $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ can be constructed accordingly via (30).
First, we have the MGF-type supermartingales whose well-known scalar case is mentioned by e.g. Howard et al. [2020, Fact 1(a)].
Example 4.6 (Matrix MGF Supermartingale).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with common conditional mean matrix $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}$ and matrix-valued conditional MGF upper bound $\mathbf{G}(\gamma)\succeq\mathbb{E}(\mathrm{e}^{\gamma(\mathbf{X}_{n}-\mathbf{M})}|\mathcal{F}_{n-1})$. Then, for each $n$, we can take $\mathbf{E}_{n}=\mathrm{e}^{\gamma_{n}(\mathbf{X}_{n}-\mathbf{M})}$, $\mathbf{F}_{n}=\mathbf{G}(\gamma_{n})^{-1}$ where $\{\gamma_{n}\}$ are predictable scalars, in Lemma 4.5.
The MGF upper bounds to use may be found in Table 1. Second, the “betting” style martingales by Waudby-Smith and Ramdas [2023].
Example 4.7 (Betting).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}^{+}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and upper bounds $\mathbf{B}_{n}$ that are predictable (i.e., $\mathbf{X}_{n}\preceq\mathbf{B}_{n}$ a.s., and $\mathbf{B}_{n}$ is $\mathcal{F}_{n-1}$-measurable). We can take, in Lemma 4.5,
$\mathbf{E}_{n}=\mathbf{I}+\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n}),\quad\mathbf{F}_{n}=\mathbf{I},$ (33)
where
$-\frac{1}{\lambda_{\max}(\mathbf{B}_{n}-\mathbf{M}_{n})}<\gamma_{n}<\frac{1}{\lambda_{\max}(\mathbf{M}_{n})}$ (34)
is predictable.
To see that this is a valid $(\mathbf{E}_{n},\mathbf{F}_{n})$ pair for Lemma 4.5, it is clear that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})=\mathbf{I}$ and (34) ensures $\mathbf{E}_{n}\succeq 0$.
Besides these two examples where we extend prior scalar results to matrices, we also note that Howard et al. [2020, Section 6.5] prove a total of eight inequalities for matrix martingale differences $\{\mathbf{Z}_{n}\}$ in the style of
$\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\exp(\mathbf{C}_{n}^{\prime}),$ (35)
for some respectively $\mathcal{F}_{n}$-, $\mathcal{F}_{n-1}$-measurable matrices $\mathbf{C}_{n},\mathbf{C}_{n}^{\prime}$ under eight different conditions. These directly lead to $\mathcal{S}_{d}^{+}$-valued supermartingales via Lemma 4.5. However, Howard et al. [2020] do not consider combining these conditional expectation bounds into matrix-valued supermartingales via (30). For reasons we will discuss later in Section 5; they pursue scalar supermartingales instead. Of their eight inequalities, we will introduce two interesting cases now, omitting the other six due to their similar construction.
First, the self-normalized supermartingale that only assumes the second moment exists [Howard et al., 2020, Part (f) of Section 6.5], which we (re)prove in Section B.3.
Example 4.8 (Self-Normalized Matrix Supermartingale).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional variance upper bounds $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ that are predictable. We can take, in Lemma 4.5,
$\mathbf{E}_{n}=\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6),\quad\mathbf{F}_{n}=\exp\left(-\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right)$ (36)
where $\{\gamma_{n}\}$ are predictable positive scalars.
Second, a bound for symmetric random matrices that have a symmetric (around $0\in\mathcal{S}_{d}$), potentially heavy-tailed distribution [Howard et al., 2020, Part (d) of Section 6.5] which we prove as well in Section B.3.
Example 4.9.
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and satisfy $\mathbf{X}_{n}-\mathbf{M}_{n}|\mathcal{F}_{n-1}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}_{n}-\mathbf{X}_{n}|\mathcal{F}_{n-1}$. Then, we can take, in Lemma 4.5,
$\mathbf{E}_{n}=\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right),\quad\mathbf{F}_{n}=\mathbf{I},$ (37)
where $\{\gamma_{n}\}$ are predictable positive scalars.
Finally, let us present a novel matrix extension of the empirical Bernstein bound, in the style of Howard et al. [2021, Theorem 4], following whom we define the function
$\psi_{E}(x)=-\log(1-x)-x=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$ (38)
for $x\in[0,1)$. The notation comes from the fact that $\psi_{E}$ is the log-MGF of a centered standard exponential random variable. We distinguish the following result for its novelty by labeling it a “proposition” as opposed to an “example”.
Proposition 4.10 (Empirical Bernstein Matrix Supermartingale).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$. Further, suppose there is a predictable and integrable sequence of $\mathcal{S}_{d}$-valued random matrices $\{\widehat{\mathbf{X}}_{n}\}$ such that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$. Then, we can take, in Lemma 4.5,
$\mathbf{E}_{n}=\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}),\quad\mathbf{F}_{n}=\exp(-\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (39)
where $\{\gamma_{n}\}$ are predictable $(0,1)$-valued scalars.
We delay the proof that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ to Section B.4. The reason why it is called “empirical” Bernstein is the “plug-in empirical variance” component term $(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}$, as $\widehat{\mathbf{X}}_{n}$ can usually be taken as an empirical average (possibly weighted) from $\mathbf{X}_{1}$ to $\mathbf{X}_{n-1}$, in which case the condition $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ is satisfied when $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$. We shall come to this empirical Bernstein result again with some scalarization tricks in Section 5.3, where we obtain a strikingly sharp spectral concentration inequality in closed form, further justify its name, discuss the context of empirical Bernstein bounds, and compare it with other existing results.
These supermartingales lead to implicit concentration bounds when applied Theorem 4.4, and are thus unfavorable when one needs a confidence interval or confidence sequence. However, these supermartingales can be used to test whether $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ or not and are thus “test supermartingales” only.
4.3 Matrix e-Processes
Scalar e-processes are an important class of processes generalizing scalar nonnegative supermartingales [Ramdas et al., 2022]. We recall the definition of these processes below, noting that we will always consider a sequence of observations $X_{1},X_{2},\dots$ (while allowing their distribution to vary) and denote their canonical filtration by $\{\mathcal{F}_{n}\}$. We use notations $\mathbb{P}_{P}$, $\mathbb{E}_{P}$ and the general terminology “for $P$” (e.g. we may say a stochastic process adapted to $\{\mathcal{F}_{n}\}$ is a martingale for $P$) to refer to probabilistic statements when the observations $X_{1},X_{2},\dots$ are distributed jointly according to $P$.
Let $\mathcal{P}$ be a set of distributions over $\mathbb{R}\times\mathbb{R}\times\dots$. A nonnegative stochastic process $\{Y_{n}\}$ adapted to $\{\mathcal{F}_{n}\}$ is called an e-process for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, there exists a supermartingale $\{M_{n}^{P}\}$ adapted to $\{\mathcal{F}_{n}\}$ with $M_{0}^{P}\leqslant 1$ that upper bounds the process $\{Y_{n}\}$. Equivalently, $\{Y_{n}\}$ is an e-process if $\mathbb{E}_{P}[Y_{\tau}]\leqslant 1$ for any stopping time $\tau$ on $\{\mathcal{F}_{n}\}$ [Ramdas et al., 2020, Lemma 6]. Clearly, if $\{Y_{n}\}$ is a nonnegative supermartingale for $\mathcal{P}$ (meaning for every $P\in\mathcal{P}$), the optional stopping theorem implies that it is also an e-process for $\mathcal{P}$, but e-processes are a much more general class [Ramdas et al., 2022, Ruf et al., 2023].
A scalar e-process satisfies the following form of (VI) or (UVI):
for $\text{for any }P\in\mathcal{P},\mathbb{P}_{P}(Y_{\tau}\geqslant Ua)\leqslant a^{-1}$
for any stopping time $\tau$ and independent super-uniform random variable $U$ (and thus we also have that $\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$ by setting $U=1$). This property enables scalar e-processes to be used for testing a hypothesis that the data is drawn from some $P\in\mathcal{P}$: we may reject the null as soon as the e-process exceeds $1/\alpha$ (for some prespecified constant $\alpha\in(0,1)$) and be guaranteed that the type-I error is at most $\alpha$. E-processes are particularly useful for composite sequential testing problems for which nonnegative supermartingales are powerless [Ramdas et al., 2022, Ruf et al., 2023].
We can generalize the scalar e-processes to our matrix setting as well, with the sequence of observations that generates the filtration $\{\mathcal{F}_{n}\}$ now denoted by $\mathbf{X}_{1},\mathbf{X}_{2},\dots$.
Definition 4.11 (Matrix e-Process).
Let $\mathfrak{P}$ be a set of distributions over $\mathcal{S}_{d}\times\mathcal{S}_{d}\times\dots$. An $\mathcal{S}_{d}^{+}$-valued stochastic process $\{\mathbf{Y}_{n}\}$ adapted to $\{\mathcal{F}_{n}\}$ is called a matrix e-process for $\mathfrak{P}$ if, for any $\mathfrak{p}\in\mathfrak{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{\mathfrak{p}}\}$ adapted to $\{\mathcal{F}_{n}\}$ with $\mathbf{M}_{0}^{\mathfrak{p}}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{\mathfrak{p}}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $\mathfrak{p}$.
Similar to the scalar case, matrix e-processes satisfy their (UMVI).
Lemma 4.12 (Uniformly Randomized Matrix Ville’s Inequality for Matrix e-Processes).
Let $\{\mathbf{Y}_{n}\}$ be a matrix e-process for $\mathfrak{P}$. Then, for any $\mathfrak{p}\in\mathfrak{P}$, for any stopping time $\tau$ on $\{\mathcal{F}_{n}\}$, any trace super-uniform random matrix $\mathbf{U}$ taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ and any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}_{\mathfrak{p}}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1}).$ (40)
Proof.
Consider the upper bounding matrix supermartingale $\{\mathbf{M}_{n}^{\mathfrak{p}}\}$. By (UMVI), $\mathbb{P}(\mathbf{M}_{\tau}^{\mathfrak{p}}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1})$. Since $\mathbf{M}_{n}^{\mathfrak{p}}\succeq\mathbf{Y}_{n}$ almost surely for any $n$, $\mathbf{M}_{\tau}^{\mathfrak{p}}\succeq\mathbf{Y}_{\tau}$ almost surely as well. Therefore, $\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathbb{P}(\mathbf{M}_{\tau}^{\mathfrak{p}}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1}).$
Simple examples of matrix e-processes following the previous subsection include lower bounds of $\mathcal{S}_{d}^{+}$-valued supermartingales. First, we define the following “matrix minimum”: for $\mathbf{A},\mathbf{B}\in\mathcal{S}_{d}$, we define $\mathbf{A}\curlyvee\mathbf{B}$ to be $\mathbf{A}$ if $\mathbf{A}\preceq\mathbf{B}$, and $\mathbf{A}-\lambda_{\max}(\mathbf{A}-\mathbf{B})\cdot\mathbf{I}$ if otherwise. Note that $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{A}$ and $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{B}$, and we remark that in general there is no largest symmetric matrix lower bound of $\mathbf{A}$ and $\mathbf{B}$ in the $\preceq$ order [Kadison, 1951, Theorem 6].
Let $\mathbf{M},\mathbf{M}^{\prime}\preceq\mathbf{B}$ be three matrices in $\mathcal{S}_{d}^{++}$, and we define $\mathcal{B}(\mathbf{B},\mathbf{M})$ to be the set of distributions on $\mathcal{S}_{d}^{+}$ upper bounded by $\mathbf{B}$ with mean $\mathbf{M}$. Further, we define the distribution families
$\mathfrak{P}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M})\text{ for all }n\}$ (41)
and
$\mathfrak{P}^{\prime}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M}^{\prime})\text{ for all }n\}.$ (42)
According to Example 4.7, we can construct a $\mathcal{S}_{d}^{+}$-valued betting process $\{\mathbf{Y}_{n}\}$ which is a supermartingale under any of $\mathfrak{p}\in\mathfrak{P}$, and another $\{\mathbf{Y}_{n}^{\prime}\}$ which is a supermartingale under any of $\mathfrak{p}\in\mathfrak{P}^{\prime}$. Then the minimum process $\{\mathbf{Y}_{n}\curlyvee\mathbf{Y}_{n}^{\prime}\}$ is a matrix $\mathfrak{P}\cup\mathfrak{P}^{\prime}$-e-process, useful for testing sequentially the null that the true mean is always either $\mathbf{M}$ or $\mathbf{M}^{\prime}$.
4.4 Matrix-Valued Submartingales and Backward Supermartingales
Next, we note that all the above are still valid if we replace “supermartingale”s with “submartingale”s. To wit, we say that $\{\mathbf{Y}_{n}\}$ is a submartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq\mathbf{Y}_{n-1}$. It is clear from the proof of Lemma 4.3 that,
Lemma 4.13.
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued submartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Then, $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a submartingale on $\{\mathcal{F}_{n}\}$. Consequently, if $\tau\leqslant N$ almost surely for some $N\in\mathbb{N}$, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{N}$.
Doob’s submartingale inequality, which is “dual” to Ville’s supermartingale inequality, is also matrix generalizable.
Theorem 4.14 (Matrix Doob’s inequality).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale. For every $N\in\mathbb{N}$ and every $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}\left(\exists n\leqslant N,\mathbf{Y}_{n}\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbb{E}(\mathbf{Y}_{N}\mathbbmss{1}_{\{\exists n\leqslant N,\mathbf{Y}_{n}\npreceq\mathbf{A}\}})\mathbf{A}^{-1})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{N})\mathbf{A}^{-1}).$ (43)
Proof.
Let $E$ be the event $\{\exists n\leqslant N,\mathbf{Y}_{n}\npreceq\mathbf{A}\}$. Define the stopping time
$\tau=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}\wedge N,$ (44)
so it satisfies $\tau\leqslant N$ and
$\{\mathbf{Y}_{\tau}\mathbbmss{1}_{E}\npreceq\mathbf{A}\}=E,\quad\mathbf{Y}_{\tau}\mathbbmss{1}_{\overline{E}}=\mathbf{Y}_{N}\mathbbmss{1}_{\overline{E}}.$ (45)
Thus by Lemmas 4.13 and 3.3,
$\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{N})\mathbf{A}^{-1/2})\geqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})$ (46)
$=$ $\mathrm{tr}(\mathbf{A}^{-1/2}\mathbb{E}(\mathbf{Y}_{\tau}\mathbbmss{1}_{E})\mathbf{A}^{-1/2})+\mathrm{tr}(\mathbf{A}^{-1/2}\mathbb{E}(\mathbf{Y}_{\tau}\mathbbmss{1}_{\overline{E}})\mathbf{A}^{-1/2})$ (47)
$\geqslant$ $\mathbb{P}(\mathbf{Y}_{\tau}\mathbbmss{1}_{E}\npreceq\mathbf{A})+\mathrm{tr}(\mathbf{A}^{-1/2}\mathbb{E}(\mathbf{Y}_{N}\mathbbmss{1}_{\overline{E}})\mathbf{A}^{-1/2}),$ (48)
implying that $\mathbb{P}(E)\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}\mathbb{E}(\mathbf{Y}_{N}\mathbbmss{1}_{E})\mathbf{A}^{-1/2})$. ∎
Next, we illustrate how Theorem 4.14 remains valid even without assuming that the process $\{\mathbf{Y}_{n}\}$ starts at $n=0$ — or starts at all; it is applicable even if $\{\mathbf{Y}_{n}\}$ is indexed on $n\in\{\dots,-2,-1,0,1,\dots,N\}$. We formalize this by considering a backward filtration $\{\mathcal{G}_{n}\}$, i.e., $\mathcal{G}_{0}\supseteq\mathcal{G}_{1}\supseteq\dots$, and an adapted backward submartingale $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{Y}_{n}$ is $\mathcal{G}_{n}$-measurable and $\mathbb{E}[\mathbf{Y}_{n}|\mathcal{G}_{n+1}]\succeq\mathbf{Y}_{n+1}$, while keeping the time indices $n\geqslant 0$. This, in the scalar regime, sets the stage for the backward Ville’s inequality which has been shown equivalent to the (forward) Doob’s inequality (hence the duality between Ville’s and Doob’s inequality; see the second proof of Theorem 2 in Manole and Ramdas [2023]), for which we have the following matrix version.
Theorem 4.15 (Matrix Backward Ville’s Inequality).
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued backward submartingale adapted to the backward filtration $\{\mathcal{G}_{n}\}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (49)
Proof.
Take an arbitrary $N\in\mathbb{N}$ and consider, indexed by $0\leqslant n\leqslant N$, the process $\mathbf{P}_{n}=\mathbf{Y}_{N-n}$ on the (forward) filtration $\mathcal{F}_{n}=\mathcal{G}_{N-n}$. It is clear that $\{\mathbf{P}_{n}\}$ is a (forward) submartingale on $\{\mathcal{F}_{n}\}$, to which we can apply Theorem 4.14,
$\mathbb{P}\left(\exists n\leqslant N,\mathbf{P}_{n}\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{P}_{N})\mathbf{A}^{-1}).$ (50)
Substituting $M_{n}$ back to the inequality above, we have,
$\mathbb{P}\left(\exists n\leqslant N,\mathbf{Y}_{n}\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (51)
The desired inequality is immediately proved by appealing to the fact that
$\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}=\liminf_{N\to\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}=\bigcup_{k=1}^{\infty}\bigcap_{N=k}^{\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}.$ (52)
4.5 Exchangeable Matrices
We now have the tools to prove a matrix version of the exchangeable Chebyshev inequality (XCI) we mentioned earlier. Recall from Fact 2 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $21\leqslant p<$ 2) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
$\frac{1}{n}(\mathbf{X}_{1}^{p}+\dots+\mathbf{X}_{n}^{p})\succeq\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}\right)^{p}.$ (53)
This gives rise to the following backward submartingale for exchangeable matrices.
Lemma 4.16.
Let $p\in[1,2]$, and $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued, $L^{p}$ random matrices. Suppose $\mathbf{C}\in\mathcal{S}_{d}$ is deterministic, and either of the following holds:
  • $p=2$; or
  • $\mathbf{X}_{n}\succeq\mathbf{C}$ almost surely for all $n$.
Then, the process
$\mathbf{R}_{n}=(\overline{\mathbf{X}}_{n}-\mathbf{C})^{p}$ (54)
is a backward submartingale adapted to the exchangeable backward filtration
$\mathcal{E}_{n}=\sigma(f(\mathbf{X}_{1},\mathbf{X}_{2},\dots):f\text{ symmetric w.r.t.\ first $n$ arguments}).$ (55)
Proof.
Recalling that the conditional expectation w.r.t. the exchangeable $\sigma$-algebra $\mathcal{E}_{n+1}$ equals the average over all $(n+1)!$ permutations of the indices $11,2,\dots,n+$ 1 (see e.g. Klenke [2013, Theorem 12.10]), we have
$\mathbb{E}[\mathbf{R}_{n}|\mathcal{E}_{n+1}]=\frac{1}{(n+1)!}\sum_{\sigma\in\mathrm{S}_{n+1}}\left(\frac{\mathbf{X}_{\sigma(1)}+\dots+\mathbf{X}_{\sigma(n)}}{n}-\mathbf{C}\right)^{p}$ (56)
(${{\succeq}}$ $\left(\frac{\sum_{\sigma\in\mathrm{S}_{n+1}}\left(\frac{\mathbf{X}_{\sigma(1)}+\dots+\mathbf{X}_{\sigma(n)}}{n}-\mathbf{C}\right)}{(n+1)!}\right)^{p}=\mathbf{R}_{n+1},$ (57)
concluding the proof that $\{\mathbf{R}_{n}\}$ is a backward submartingale. ∎
Applying the matrix backward Ville’s inequality (Theorem 4.15) to this process $\{\mathbf{R}_{n}\}$, we immediately have the following two matrix extensions of (XCI).
Theorem 4.17 (Exchangeable Matrix Chebyshev Inequality).
Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with mean $\mathbf{M}$ and variance upper bound $\mathbb{E}(\mathbf{X}_{1}-\mathbf{M})^{2}\preceq\mathbf{V}$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\exists n,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (XMCI)
Further, if $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$, we have for any fixed $N$,
$\mathbb{P}(\exists n\geqslant N,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant N^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (58)
Proof.
Applying Theorem 4.15 to $\mathbf{R}_{n}=(\overline{\mathbf{X}_{n}}-\mathbf{M})^{2}$, which is a backward submartingale according to Lemma 4.16,
$\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{R}_{1})\mathbf{A}^{-2})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (59)
Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/2}$,
$\mathbb{P}(\exists n,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2}),$ (60)
concluding the proof of (XMCI). To show (58), simply apply Theorem 4.15 to the submartingale $\mathbf{R}_{n}=(\overline{\mathbf{X}_{n}}-\mathbf{M})^{2}$ with $n\geqslant N$, obtaining
$\mathbb{P}(\exists n\geqslant N,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{R}_{n})\mathbf{A}^{-2})\leqslant N^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ (61)
due to Theorem 3.5. ∎
Theorem 4.18 (Exchangeable Matrix $p$-Chebyshev Inequality).
Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}^{+}$-valued random matrices with raw $p$th moment $\mathbf{V}_{p}=\mathbb{E}\mathbf{X}_{1}^{p}$, where $p\in[1,2]$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
$\mathbb{P}(\exists n,\;\overline{\mathbf{X}}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (62)
Proof.
Applying Theorem 4.15 to $\{\mathbf{R}_{n}\}$, with $\mathbf{C}$ being 00, in Lemma 4.16,
$\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n})^{p}\npreceq\mathbf{A}^{p})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (63)
Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/p}$,
$\mathbb{P}(\exists n,\;\overline{\mathbf{X}}_{n}\npreceq\mathbf{A})\leqslant\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n})^{p}\npreceq\mathbf{A}^{p}),$ (64)
concluding the proof. ∎
Indeed, we can weaken the second-moment assumption to a $p$th moment, leading to, however, a concentration not around the mean but a Markov-type upper bound. This is largely because the scalar map $x\mapsto|x|^{p}$ is convex but not its matrix extension, and so the raw $p$th moment is used above. Nonetheless, due to Fact 4, the following trace inequality holds.
$\mathrm{tr}\left(\frac{1}{n}((\operatorname{abs}\mathbf{X}_{1})^{p}+\dots+(\operatorname{abs}\mathbf{X}_{n})^{p})\right)\geqslant\mathrm{tr}\left(\left(\operatorname{abs}\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}\right)^{p}\right).$ (65)
Therefore,
$r_{n}=\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{C})^{p}\right)$ (66)
is a real-valued backward submartingale. With the scalar backward Ville’s inequality, this leads to the following anytime-valid extension of Lemma 3.7.
Theorem 4.19 (Exchangeable Matrix $p$-Chebyshev Trace Inequality).
Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with mean $\mathbf{M}$ and central $p$th moment $\mathbf{V}_{p}=\mathbb{E}\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M})^{p}$, where $p\geqslant 1$. Then, for any $a>0$,
$\mathbb{P}(\exists n,\;\|\overline{\mathbf{X}}_{n}-\mathbf{M}\|_{\mathrm{op}}\geqslant a)\leqslant\mathbb{P}\left(\exists n,\;\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})^{p}\right)\geqslant a^{p}\right)\leqslant a^{-p}\mathrm{tr}\mathbf{V}_{p}.$ (67)
The bound above differs from Theorem 4.18 in that the concentration is now around the mean $\mathbf{M}$.
5 Randomized Matrix Bounds via Scalar e-Processes
5.1 A Conversion Lemma from Matrix to Scalar Supermartingales
We have studied the concentration of random matrices via matrix e-statistics, i.e. random matrices with bounded expectations, matrix supermartingales, matrix e-values, and matrix backward submartingales, since these quantities yield concentration inequalities via our Markov and Ville’s inequalities for matrices, oftentimes in randomized forms. Some further randomized matrix concentration inequalities can be obtained by revisiting some bounds obtained via scalar e-quantities that already exist in the literature.
As we mentioned earlier in Section 4.2, Howard et al. [2020, Section 6.5] prove, under different conditions, numerous inequalities for symmetric matrix martingale difference $\{\mathbf{Z}_{n}\}$ in the style of
$\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\exp(\mathbf{C}_{n}^{\prime}),$ (68)
for some respectively $\mathcal{F}_{n}$-, $\mathcal{F}_{n-1}$-measurable symmetric matrices $\mathbf{C}_{n},\mathbf{C}_{n}^{\prime}$.
At this step, our previous methods and that of Howard et al. [2020] had diverged. We have been able to directly construct matrix-valued supermartingales in Section 4.2, leading potentially to computable sequential tests by virtue of our new (UMVI). Howard et al. [2020], on the other hand, construct scalar e-processes from the condition (68). Let us rephrase their “master lemma” behind this construction.
Lemma 5.1 (Lemma 4 in Howard et al. [2020], rephrased and generalized).
Let $\{\mathbf{Z}_{n}\}$ be an $\mathcal{S}_{d}$-valued martingale difference sequence adapted to $\{\mathcal{F}_{n}\}$. Let $\{\mathbf{C}_{n}\}$ be an $\mathcal{S}_{d}$-valued adapted process, $\{\mathbf{C}_{n}^{\prime}\}$ be an $\mathcal{S}_{d}$-valued predictable process, w.r.t. the same filtration. If (68) holds for all $n$, then, the process
$L_{n}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)$ (69)
is a (scalar) supermartingale. Further,
$L_{n}\geqslant\exp\left(\lambda_{\max}\left(\sum_{i=1}^{n}\mathbf{Z}_{i}\right)-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\right).$ (70)
The proof of this lemma can be found in Section B.5. Note that the upper bounded expression (70) is an instance of e-processes, introduced earlier in Section 4.3. We shall see soon that both (69) and (70) can be useful.
5.2 Examples: Self-Normalization, Bernstein, and Hoeffding
Recall from Example 4.8 the self-normalized bound, that if $\{\mathbf{X}_{n}\}$ are symmetric matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\{\mathbf{M}_{n}\}$ and predictable conditional variance upper bounds $\{\mathbf{V}_{n}\}$,
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\mid\mathcal{F}_{n-1}\right)\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (71)
an instance of (68). By Lemma 5.1, this implies the following statement whose weaker version is stated as Lemma 3(f) by Howard et al. [2020].
Example 5.2 (Matrix Self-Normalized Scalar Supermartingale).
Under the same assumptions as Example 4.8, the process
$L_{n}^{\textsf{SN}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)$ (72)
is a supermartingale.
The proof is simply taking $\mathbf{Z}_{n}=\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})$, $\mathbf{C}_{n}=\frac{\gamma_{n}^{2}}{6}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}$, and $\mathbf{C}_{n}^{\prime}=\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}$ in Lemma 5.1. Some remarks on this result are in order. First, the expression of the scalar supermartingale $\{L_{n}^{\textsf{SN}}\}$ explains the name “self-normalized”, as an empirical variance term exists within the bound. Second, unlike the scalar case where a closed-form confidence sequence can be solved from the supermartingale combining (VI) if the means are a constant [Wang and Ramdas, 2023a, Section 5], in our matrix case, we only have an implicit bound when applying (UVI).
Corollary 5.3 (Uniformly Randomized Stopped Matrix Self-Normalized Bound).
Under the same assumptions as Example 4.8, for any stopping time $\tau$, $\alpha\in(0,1)$, and a super-uniform $U$ independent from $\mathcal{F}_{\infty}$,
$\mathbb{P}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{\tau}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{\tau}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)\geqslant dU/\alpha\right)\leqslant\alpha.$ (73)
The bound above, as those matrix supermartingales in Section 4.2, does not bear closed-form confidence sets but still implies a computable randomized sequential test for the null, e.g., that the means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})$ all equal $\mathbf{M}$ with marginal type I error rate within $\alpha$.
However, a closed-form concentration can be computed when further assumption is introduced on top of the finite variance in Example 4.8. This is the Hoeffding-type e-process by Howard et al. [2020, Lemma 3], which we present as a Bernstein-type result akin to Mackey et al. [2014, Corollary 4.2] for generality.
Example 5.4 (Matrix Bernstein Scalar e-Process).
Under the same assumptions as Example 4.8, suppose further that $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}\preceq\mathbf{B}_{n}$ almost surely for some predictable sequence $\{\mathbf{B}_{n}\}$. Then, the self-normalized supermartingale $\{L_{n}^{\textsf{SN}}\}$ upper bounds the following process:
$L_{n}^{\textsf{SN}}\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)\right\}.$ (74)
Proof.
Taking $\psi(x)=x^{2}/2$, $\mathbf{C}_{n}=\frac{1}{3}(\mathbf{X}_{n}-\mathbf{M})^{2}$, and $\mathbf{C}_{n}^{\prime}=\frac{2}{3}\mathbf{V}_{n}$ in Lemma 5.1, from (70),
$L_{n}^{\textsf{SN}}$ $\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)\right\}$ (75)
$\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)\right\},$ (76)
concluding the proof. ∎
A closed-form, randomized, stopping time-valid matrix Bernstein or Hoeffding inequality on the concentration of weighted sample average is immediately in force by applying (UVI) to $\{L_{n}^{\textsf{SN}}\}$. Below, we denote by $\overline{\mathbf{X}}_{n}^{\gamma}$ the weighted average $\frac{\gamma_{1}\mathbf{X}_{1}+\dots+\gamma_{n}\mathbf{X}_{n}}{\gamma_{1}+\dots+\gamma_{n}}$ w.r.t. some positive weight sequence $\{\gamma_{n}\}$.
Theorem 5.5 (Uniformly Randomized Stopped Matrix Bernstein/Hoeffding Inequality).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$, $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}\preceq\mathbf{B}_{n}$ and $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ almost surely for predictable $\{\mathbf{B}_{n}\}$ and $\{\mathbf{V}_{n}\}$. Then, for any stopping time $\tau$, predictable positive sequence $\{\gamma_{n}\}$, $\alpha\in(0,1)$, and super-uniform random scalar $U$ independent of $\mathcal{F}_{\infty}$,
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{6}\lambda_{\max}\left(\sum{\gamma_{i}^{2}}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)$$\leqslant$ (77)
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{2}\lambda_{\max}\left(\sum_{i=1}^{\tau}\gamma_{i}^{2}\mathbf{B}_{i}\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)$ $\leqslant\alpha.$ (78)
We now remark at length on this theorem. First, the Hoeffding term $\frac{1}{2}\lambda_{\max}\left(\sum\gamma_{i}^{2}\mathbf{B}_{i}\right)$ comes from applying $\mathbf{V}_{i}\preceq\mathbf{B}_{i}$ to the Bernstein term $\frac{1}{6}\lambda_{\max}\left(\sum\gamma_{i}^{2}(\mathbf{B}_{i}^{2}+2\mathbf{V}_{i})\right)$, among which one may choose, as is with scalar Bernstein versus Hoeffding, depending on whether variance bounds $\{\mathbf{V}_{n}\}$ better than the total square bounds $\{\mathbf{B}_{n}\}$ are known. Second, the strength of this bound can be seen from the following fact: taking $\{\mathbf{M}_{n}\}$ to be a constant $\mathbf{M}$, $\{\mathbf{B}_{n}\}$ a constant $\mathbf{B}$, $\{\mathbf{V}_{n}\}$ a constant $\mathbf{V}$, $\tau$ a deterministic sample size $n$ and letting $U=1$, $\gamma_{1}=\dots=\gamma_{n}=\sqrt{\frac{6\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B}+2\mathbf{V})}}$, (77) reads
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\right)\geqslant\sqrt{\frac{2\log(d/\alpha)\lambda_{\max}(\mathbf{B}+2\mathbf{V})}{3n}}\right)\leqslant\alpha,$ (79)
where $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$. This even improves the optimal matrix Hoeffding inequality by Mackey et al. [2014, Corollary 4.2] whose dependence on $\mathbf{B}$ and $\mathbf{V}$ is $\frac{1}{2}\lambda_{\max}(\mathbf{B}+\mathbf{V})$ compared to our smaller $\frac{1}{3}\lambda_{\max}(\mathbf{B}+2\mathbf{V})$, [^4] The optimality of Mackey et al. [2014, Corollary 4.2] is in terms of the overall multiplicative constant which (79) achieves as well, not the distribution of the constant between $\mathbf{B}$ and $\mathbf{V}$. Even tighter dependence in this regard shall be presented soon in Section 5.3. and recovers (MHI), both via $\mathbf{V}\preceq\mathbf{B}$. The uncompromised inequality (77) further allows for stopping time-validity, time-varying and adaptive weighting, and randomized tightening simultaneously. Third, in the i.i.d. (or homoscedastic more generally) sequential (sample size $\tau$ not fixed) case, following Waudby-Smith and Ramdas [2020], Wang and Ramdas [2023a], the weights $\{\gamma_{n}\}$ can be set as a vanishing sequence such that each $\gamma_{n}$ matches the fixed-time optimal choice of $\gamma_{n}$ with sample size $n$, so we recommend $\gamma_{n}=\sqrt{\frac{6\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B}+2\mathbf{V})}}$. By doing so, the concentration tightens itself as time goes on; or, the corresponding sequential test has asymptotic power 1 [Waudby-Smith and Ramdas, 2023, Table 1]. Moreover, it can be seen that the closed-form concentration of $\overline{\mathbf{X}}_{\tau}$ here is achieved via multiple steps of scalarization, each step loosening the bound by a little. Therefore, if one does not want to quantify uncertainty but is only interested in sequentially testing if the means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}$, the more powerful test in the current subsection is still with the largest process, the self-normalized supermartingale $\{L^{\textsf{SN}}_{n}\}$ (i.e., rejecting whenever $L^{\textsf{SN}}_{n}$ surpasses $Ud/\alpha$), with variance upper bounds $\mathbf{V}_{n}$ supplied by the upper bounds $\mathbf{B}_{n}$.
5.3 Matrix Empirical Bernstein Inequality
We now present a closed-form empirical Bernstein inequality that enjoys the same stopping time-validity, adaptive weighting, and randomization as well, foreshadowed by a matrix expectation bound in Proposition 4.10. Empirical Bernstein inequalities for scalars differ from “classical” Bernstein inequalities, e.g. our (77) when $d=1$, in that the population variance is replaced by an empirical variance in the deviation bound. They therefore achieve the advantage of Bernstein’s inequalities over Hoeffding’s inequalities without requiring knowing the variance beforehand.
Scalar empirical Bernstein inequalities are often derived in two different ways. First, a union bound between a classical Bernstein inequality and a concentration inequality on the empirical variance, a method that early empirical Bernstein results [Audibert et al., 2009, Maurer and Pontil, 2009] all share. Second, a self-normalization technique pioneered by Howard et al. [2021, Theorem 4] and later followed by Waudby-Smith and Ramdas [2023], Chugg et al. [2023], among others, which enables supermartingale-based, stopping time-valid concentration, and variance proxy by predictable estimates other than the empirical variance. These two methods are inherently different and as argued convincingly by Howard et al. [2021, Appendix A.8]: the latter’s avoidance of the union bound produces a sharper concentration. No matrix empirical Bernstein results, especially those in the second fashion, are previously known.
Our matrix extension of Howard et al. [2021, Theorem 4] starts from the bound we realier prove as Proposition 4.10. Recall that $\psi_{E}(x)=-\log(1-x)-x=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$ for $x\in[0,1)$, and that
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\middle|\mathcal{F}_{n-1}\right)\preceq\exp(\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n}))$ (80)
under, among other, a key assumption that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$. We can use Lemma 5.1 again to produce a scalar supermartingale and a closed-form spectral concentration inequality.
Theorem 5.6 (Uniformly Randomized Stopped Matrix Empirical Bernstein Inequality).
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$. Let $\{\widehat{\mathbf{X}}_{n}\}$ be a sequence of predictable and integrable $\mathcal{S}_{d}$-valued random matrices such that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ almost surely. Then, for any predictable $(0,1)$-valued sequence $\{\gamma_{n}\}$,
$L_{n}^{\mathsf{EB}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}\right)$ (81)
is a supermartingale. Further, for any stopping time $\tau$, $\alpha\in(0,1)$, and super-uniform scalar random variable $U$ independent of $\mathcal{F}_{\infty}$, we have
$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{\tau}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)\leqslant\alpha.$ (82)
Before we remark on the uncompromised Theorem 5.6, let us first write down its fixed-time, non-randomized, fine-tuned special case with $\tau=n$ and $U=1$ which shall justify the “empirical Bernstein” name it bears.
Corollary 5.7 (Matrix Empirical Bernstein Inequality).
Suppose $\alpha\in(0,1)$. Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be $\mathcal{S}_{d}^{[0,1]}$-valued i.i.d. random matrices with mean $\mathbf{M}=\mathbb{E}\mathbf{X}_{1}$. Let $\overline{\mathbf{X}}_{i}=\frac{1}{i}(\mathbf{X}_{1}+\dots+\mathbf{X}_{i})$ and $\overline{\mathbf{X}}_{0}=0$. Then, for any predictable $(0,1)$-valued sequence $\gamma_{1},\dots,\gamma_{n}$,
[…]
Further, if we define the following variance proxies
$\overline{\mathbf{V}}_{0}=0,\quad\overline{\mathbf{V}}_{k}=\frac{1}{k}\sum_{i=1}^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{k})^{2},\quad\overline{v}_{k}=\lambda_{\max}(\overline{\mathbf{V}}_{k})\vee\frac{5\log(d/\alpha)}{n},$ (84)
$\widetilde{s}_{n}=\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}^{2})}{\overline{v}_{i-1}}\right),$ (85)
and set $\gamma_{i}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}$ for $i=1,\dots,n$, then,
$D_{n}^{\mathsf{EB}}\leqslant\sqrt{\frac{\log(d/\alpha)}{2n}}\frac{1+2\widetilde{s}_{n}}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}\quad\approx\sqrt{\frac{9\log(d/\alpha)\lambda_{\max}(\mathbf{V})}{2n}}$ (86)
where $\mathbf{V}=\mathbf{Var}(\mathbf{X}_{1})$; further,
$\lim_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}=\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}\quad\text{almost surely}.$ (87)
We prove both Theorem 5.6 and Corollary 5.7 in Section B.4. The asymptotic behavior (87) of deviation bound $D_{n}^{\mathsf{EB}}$ is striking as it adapts fully to, without knowing, the true variance $\mathbf{V}$, a strict improvement even compared to the deviation bound $D_{n}^{\mathsf{B}}$ of the our previous matrix Bernstein inequality (79),
$\sqrt{n}D_{n}^{\mathsf{B}}=\sqrt{2\log(d/\alpha){\frac{1+2\lambda_{\max}(\mathbf{V})}{3}}},$ (88)
which requires knowing beforehand, but only adapts partially to, the variance $\mathbf{V}$. Indeed, when $\mathbf{V}$ is small, (88) can be arbitrarily larger than (87). We can also compare Corollary 5.7 to another matrix Bernstein inequality in the literature by Tropp [2012, Theorem 1.4], who proves that $\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\right)\geqslant D_{n}^{\mathsf{TB}}\right)\leqslant\alpha$ under the same assumptions (plus knowing the variance $\mathbf{V}$), where its deviation bound
$D_{n}^{\mathsf{TB}}=\frac{\log(d/\alpha)+\sqrt{\log^{2}(d/\alpha)+18\log(d/\alpha)n\lambda_{\max}(\mathbf{V})}}{3n}$ (89)
satisfies the same asymptotic $\lim_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{TB}}=\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}$ as (87), which we achieve without requiring knowing $\mathbf{V}$. Other matrix Bernstein bounds in the literature, like the ones by Mackey et al. [2014, Corollary 5.2] and by Oliveira [2009], are looser than the Tropp [2012, Theorem 1.4] bound we discuss here as acknowledged by Mackey et al. [2014] themselves.
Our Theorem 5.6 and Corollary 5.7 owe much to the techniques developed by Waudby-Smith and Ramdas [2023, Theorem 2 and Remark 1] in the scalar case (who in turn build on the earlier result by Howard et al. [2021, Theorem 4] via the “predictable mixing” sequence $\{\gamma_{n}\}$). We recover exactly as special cases those two known results when $d=1$.
We reiterate our remark following Proposition 4.10 that $\widehat{\mathbf{X}}_{n}$ in Theorem 5.6 can usually be taken as the empirical average from $\mathbf{X}_{1}$ to $\mathbf{X}_{n-1}$, which is exactly what is done in Corollary 5.7. In the i.i.d. (or homoscedastic more generally) sequential case without the sample size $n$ fixed in advance, we can again (recalling a similar discussion following Theorem 5.5) set the weights $\{\gamma_{n}\}$ to be a vanishing sequence such that each $\gamma_{n}$ matches the fixed-time optimal choice of $\gamma_{n}$ with sample size $n$, so we recommend $\gamma_{n}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{n-1}}}$. Under this weight sequence, we see the choice of a weighted average $\widehat{\mathbf{X}}_{n}=\overline{\mathbf{X}}_{n-1}^{\psi_{E}(\gamma)}$ is more reasonable as it roughly minimizes the weighted sum of squares $\sum_{i=1}^{\tau}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}$ in (82) into a weighted empirical variance — a heuristic which should work for Proposition 4.10 as well. Obviously, as long as $\widehat{\mathbf{X}}_{n}$ is any average, weighted or not, of $\mathbf{X}_{1},\dots,\mathbf{X}_{n-1}$, the condition $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ is met when $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$.
6 Experiments: To Scalarize or not to Scalarize
We have developed thus far two categories of matrix concentration bounds: those obtained via the pure matrix approach, typically expressed in the $\preceq$ order, in Sections 3 and 4; and those obtained via scalarization, i.e. taking the trace, in Section 5. It is therefore a natural question to ask if either of these is tighter than the other.
A fair comparison can be raised between the self-normalized matrix and scalar test supermartingales, Example 4.8 and Example 5.2 — for both are under the assumption of finite variance and based on the key one-step self-normalized bound (71). We place them in the setting of the following sequential testing problem: let $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ be an i.i.d. sequence of matrices in $\mathcal{S}_{d}$, with mean $\mathbf{M}$ and variance upper bound $\mathbf{I}$. We would like to sequentially test the null hypothesis $H_{0}:\mathbf{M}=0$. Let us call the tests derived from Example 4.8 and Example 5.2 for this null Matrix and Scalar respectively. Ignoring the randomizer for simplicity (plug in $\mathbf{U}=\mathbf{I}$), Matrix computes the matrix-valued process
$\mathbf{Y}_{n}^{\textsf{SN}}=\exp\left(-\frac{\gamma_{1}^{2}+\dots+\gamma_{n}^{2}}{3}\right)\cdot\exp\left(\frac{\gamma_{1}\mathbf{X}_{1}}{2}-\frac{\gamma_{1}^{2}\mathbf{X}_{n}^{2}}{12}\right)\dots\exp\left(\frac{\gamma_{n}\mathbf{X}_{n}}{2}-\frac{\gamma_{n}^{2}\mathbf{X}_{n}^{2}}{12}\right)\\ \exp\left(\frac{\gamma_{n}\mathbf{X}_{n}}{2}-\frac{\gamma_{n}^{2}\mathbf{X}_{n}^{2}}{12}\right)\dots\exp\left(\frac{\gamma_{1}\mathbf{X}_{1}}{2}-\frac{\gamma_{1}^{2}\mathbf{X}_{1}^{2}}{12}\right),$ (90)
rejecting when $\mathbf{Y}_{n}^{\textsf{SN}}\npreceq\mathbf{A}$ for some threshold matrix $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$; Scalar computes the scalar-valued process
$L_{n}^{\textsf{SN}}=\exp\left(-\frac{\gamma_{1}^{2}+\dots+\gamma_{n}^{2}}{3}\right)\cdot\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}\mathbf{X}_{i}-\sum_{i=1}^{n}\frac{\gamma_{i}^{2}\mathbf{X}_{i}^{2}}{6}\right),$ (91)
rejecting when $L_{n}^{\textsf{SN}}\geqslant d/\alpha$. Note that $L_{n}^{\textsf{SN}}$ does not equal $\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ in general, an important point that we shall come to later.
Under the null, both Matrix and Scalar safely control the type 1 error (the probability of falsely rejecting the null) under the predefined level $\alpha$ due to (MVI) and (VI). For power comparison, we shall first discuss the case when only $n=1$ observation is revealed, and then move on to the case with an indefinite sequence of observations, with examples where Matrix outpowers Scalar under decreasingly restrictive conditions and finally no restriction at all.
6.1 Comparison with a Single Observation
When only $n=1$ observation is revealed, $L_{1}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{1}^{\textsf{SN}})$. We shall see that the comparison depends heavily on the rejection threshold matrix $\mathbf{A}$ that Matrix picks beforehand.
Proposition 6.1.
Suppose Matrix picks $\mathbf{A}=(d/\alpha)\mathbf{I}$. Then, if Matrix rejects at $n=1$, Scalar rejects as well. In this case Scalar is at least as powerful as Matrix.
Proof.
If $\mathbf{A}=(d/\alpha)\mathbf{I}$ and Matrix rejects, then $\mathbf{Y}_{1}^{\textsf{SN}}\npreceq(d/\alpha)\mathbf{I}$, meaning that $\lambda_{\max}(\mathbf{Y}_{1}^{\textsf{SN}})>d/\alpha$. Therefore, $L_{1}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{1}^{\textsf{SN}})\geqslant\lambda_{\max}(\mathbf{Y}_{1}^{\textsf{SN}})>d/\alpha$. ∎
Proposition 6.2.
Suppose that $d\geqslant 2$ and that the random matrix $\mathbf{Y}_{1}^{\mathsf{SN}}$ has full support, meaning it can equal any matrix in $\mathcal{S}_{d}^{++}$. Then,
  • $(d/\alpha)\mathbf{I}$ is the only matrix that Matrix can pick as $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that if Matrix rejects at $n=1$, Scalar always rejects as well. Therefore, as long as $\mathbf{A}\neq(d/\alpha)\mathbf{I}$, Scalar is not at least as powerful as Matrix;
  • It is impossible that if Scalar rejects at $n=1$, Matrix always rejects as well. So Matrix is never at least as powerful as Scalar.
In summary, for general $\mathbf{A}$ and $n=1$, neither Scalar nor Matrix dominates the other.
Proof.
For the first statement, consider the spectral decomposition of the matrix $\mathbf{A}\neq(d/\alpha)\mathbf{I}$,
$\mathbf{A}=\lambda_{1}\mathbf{P}_{1}+\dots+\lambda_{d}\mathbf{P}_{d},\quad\lambda_{1}\leqslant\dots\leqslant\lambda_{d}.$ (92)
The matrix $\lambda_{1}\mathbf{P}_{1}$ satisfies $\mathrm{tr}(\lambda_{1}\mathbf{P}_{1})=\lambda_{1}<d/\alpha$. Applying a small pertubation, $\lambda_{1}\mathbf{P}_{1}+\varepsilon\mathbf{I}\npreceq\mathbf{A}$, $\mathrm{tr}(\lambda_{1}\mathbf{P}_{1}+\varepsilon\mathbf{I})<d/\alpha$. So, if $\mathbf{Y}_{1}^{\mathsf{SN}}$ takes a value that is close to $\lambda_{1}\mathbf{P}_{1}+\varepsilon\mathbf{I}$, Matrix rejects at $n=1$ but Scalar does not.
For the second statement, the matrix $0.9\mathbf{A}$ satisfies $\mathrm{tr}(0.9\mathbf{A})\geqslant 0.9d^{2}/\alpha>d/\alpha$ by Cauchy-Schwarz inequality and $0.9\mathbf{A}\preceq\mathbf{A}$. Therefore, if $\mathbf{Y}_{1}^{\mathsf{SN}}$ takes a value that is close to $0.9\mathbf{A}$, Scalar rejects at $n=1$ but Matrix does not. ∎
Therefore, one may understand the difference between Matrix and Scalar as that Matrix has the freedom to choose the threshold matrix $\mathbf{A}$ in anticipation of the potential alternative; whereas Scalar represents some kind of Pareto optimal procedure if completely agnostic about the reality. For Matrix to be more powerful it has to choose $\mathbf{A}$ wisely according to some prior knowledge. The following example further showcases this point.
Example 6.3.
Suppose Matrix knows beforehand what $\mathbf{X}_{1}$ is (and hence what $\mathbf{Y}_{1}^{\mathsf{SN}}$ is), and $\lambda_{d}>1/\alpha$ where
$\mathbf{Y}_{1}^{\mathsf{SN}}=\lambda_{1}\mathbf{P}_{1}+\dots+\lambda_{d}\mathbf{P}_{d},\quad\lambda_{1}\leqslant\dots\leqslant\lambda_{d}$ (93)
is its spectral decomposition. Then, Matrix can always make sure to reject by picking
$\mathbf{A}=\varepsilon^{-1}(d-1)\mathbf{P}_{1}+\dots\varepsilon^{-1}(d-1)\mathbf{P}_{d-1}+(\alpha-\varepsilon)^{-1}\mathbf{P}_{d}$ (94)
with $\varepsilon\in(0,\alpha-\lambda_{d}^{-1})$. The condition $\lambda_{d}>1/\alpha$ here is strictly weaker than the condition under which Scalar rejects: $\lambda_{1}+\dots+\lambda_{d}=d/\alpha$.
Indeed, as we can see from these examples with $n=1$ (thus we have only compared the matrix Markov inequality with the scalar Markov inequality on the trace), for Matrix to outpower Scalar, apart from prior knowledge, it is favorable that $\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})$ is an outlier in the spectrum of $\mathbf{Y}_{1}^{\mathsf{SN}}$, i.e., ${\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{1}^{\mathsf{SN}})}$ is close to 1. This is exactly what we would predict from the discussion on (UMMI) early on, where we mention that (UMMI) is tighter when the random matrix has (approximately) a lower rank. In the very next subsection, we shall see that whereas the condition ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\approx 1$ seems artificial and contrived here with $n=1$, it becomes an almost certain phenomenon with large $n$.
6.2 Comparison with a Sequence of Observations
Let us consider first an example of a sequence of commuting observations, where Matrix’s excess power compared to Scalar increases as time goes on.
Example 6.4.
Suppose each observation has the spectral decomposition
$\mathbf{X}_{n}=\lambda_{1n}\mathbf{P}_{1}+\dots+\lambda_{dn}\mathbf{P}_{d}$ (95)
with common deterministic projection matrices $\mathbf{P}_{1},\dots,\mathbf{P}_{d}$. For each $k$, $\lambda_{k1},\lambda_{k2},\dots\overset{\mathrm{iid}}{\sim}\pi_{k}$ with mean $\mathbb{E}\lambda_{k1}=\lambda_{k}$ and variance $\mathbf{Var}\lambda_{k1}=\sigma_{k}^{2}\leqslant 1$. Then, the variance $\mathbf{Var}(\mathbf{X}_{1})=\sigma_{1}^{2}\mathbf{P}_{1}+\dots+\sigma_{d}^{2}\mathbf{P}_{d}\preceq\mathbf{I}$. The matrix-valued process that Matrix computes is
$\mathbf{Y}_{n}^{\textsf{SN}}=\sum_{k=1}^{d}Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})\mathbf{P}_{k},$ (96)
and the scalar-valued process that Scalar computes is
$L_{n}^{\textsf{SN}}=\sum_{k=1}^{d}Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}}).$ (97)
In both expressions the function $Y_{n}^{\textsf{SN}}(\cdot)$ denotes the self-normalized test process for scalar observations,
$Y_{n}^{\textsf{SN}}(x_{1},\dots,x_{n})=\exp\left(-\frac{1}{3}\sum_{i=1}^{n}\gamma_{i}^{2}+\sum_{i=1}^{n}\gamma_{i}x_{i}-\frac{1}{6}\sum_{i=1}^{n}\gamma_{i}^{2}x_{i}^{2}\right).$ (98)
The scalar component process $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows exponentially fast if $\pi_{k}$ has mean $\lambda_{k}\neq 0$, and shrinks to 0 if $\lambda_{k}=0$; moreover, the rate in which $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows depends exponentially on the size $|\lambda_{k}|$. Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will be approximately low-rank with ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})\mathbf{P}_{d}$ will exponentially outgrow the rest. In this case if Matrix correctly anticipates the largest component $\mathbf{P}_{d}$, by picking again the threshold matrix
$\mathbf{A}=\varepsilon^{-1}(d-1)\mathbf{P}_{1}+\dots\varepsilon^{-1}(d-1)\mathbf{P}_{d-1}+(\alpha-\varepsilon)^{-1}\mathbf{P}_{d}$ (99)
with small $\varepsilon$ it will outperform Scalar: it only takes $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})$ to surpass $(\alpha-\varepsilon)^{-1}\approx 1/\alpha$ for Matrix to reject, whereas it takes $L_{n}^{\textsf{SN}}\approx Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})$ to surpass $d/\alpha$ for Scalar to reject. Similar reasoning applies as long as Matrix knows a subset of the $d$ components with larger means.
We experiment with the following setup. We pick an orthonormal basis in $d=3$ dimension: $\mathbf{u}_{1}=(1/\sqrt{2},1/\sqrt{2},0)^{\mathsf{T}}$, $\mathbf{u}_{2}=(0,0,1)^{\mathsf{T}}$, $\mathbf{u}_{3}=(1/\sqrt{2},-1/\sqrt{2},0)^{\mathsf{T}}$, and let $\mathbf{P}_{1}=\mathbf{u}_{1}\mathbf{u}_{1}^{\mathsf{T}}$, $\mathbf{P}_{2}=\mathbf{u}_{2}\mathbf{u}_{2}^{\mathsf{T}}$, $\mathbf{P}_{3}=\mathbf{u}_{3}\mathbf{u}_{3}^{\mathsf{T}}$. The spectral distributions are set as $\pi_{1}=\mathcal{N}(0.1,1)$, $\pi_{2}=\mathcal{N}(-0.2,1)$, and $\pi_{3}=\mathcal{N}(0.3,1)$. We consider the following three sequential tests: smart Matrix, one that follows the recipe (99) with $\varepsilon=\alpha/20$; naive Matrix, one that is agnostic about the real distribution and picks $\mathbf{A}=(d/\alpha)\mathbf{I}$; and Scalar, all with the parameters $\gamma_{n}=n^{-1/2}$ and $\alpha=0.05$. We record the times at which these tests correctly reject the null, as well as the evolution of ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$. The findings are displayed in Figure 1: smart Matrix indeed outpowers the other two, whereas naive Matrix and Scalar are similar in power, explained by the quick convergence of ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$ to 1.
Figure 1: Empirical findings with commuting i.i.d. random matrices. Left. Comparison of the cumulative distribution functions of rejection times of the sequential tests corresponding to the matrix- and scalar-valued test processes under the same alternative distribution. Right. Evolution of the spectral distribution of the matrix-valued test process under the alternative.
While this indicates that
smart Matrix $>$ naive Matrix $\approx$ Scalar
in power, we now come to the striking (in light of Proposition 6.1) phenomenon that with the commutativity assumption removed, it is actually possible that
smart Matrix $>$ naive Matrix $>$ Scalar.
In the following example that involves non-commuting sequential observations, we consider the application scenario of covariance estimation [Ke et al., 2019].
Example 6.5.
Suppose $\mathbf{X}_{n}=\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}}-(2d)^{-1}\mathbf{I}$ where $\mathbf{x}_{1},\mathbf{x}_{2},\dots\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,\mathbf{\Sigma})$. We are to test if the null $\mathbf{\Sigma}=(2d)^{-1}\mathbf{I}$ is true, [^5] The null covariance $(2d)^{-1}\mathbf{I}$ is picked in the mere interest of simplicity so that the variance upper bound $\mathbf{Var}(\mathbf{X}_{n})\preceq\mathbf{I}$ is easily realizable in our example; it is easy to see that any $\mathbf{\Sigma}\in\mathcal{S}_{d}^{++}$ works as well by adjusting the variance upper bound accordingly. i.e., $\mathbb{E}\mathbf{X}_{n}=0$, under the 4th moment assumption $\mathbb{E}\|\mathbf{x}_{n}\|^{4}=(\mathrm{tr}\mathbf{\Sigma})^{2}+2\mathrm{tr}(\mathbf{\Sigma}^{2})\leqslant 1$, which implies
$\mathbf{Var}(\mathbf{X}_{n})=\mathbf{Var}(\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}})\preceq\mathbb{E}(\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}}\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}})=\mathbb{E}(\|\mathbf{x}_{n}\|^{2}\cdot\mathbf{x}_{n}\mathbf{x}_{n}^{\mathsf{T}})\preceq\mathbb{E}(\|\mathbf{x}_{n}\|^{2}\cdot\|\mathbf{x}_{n}\|^{2}\mathbf{I})\preceq\mathbf{I}.$ (100)
Suppose in reality Matrix knows that $\mathbf{\Sigma}\neq(2d)^{-1}\mathbf{I}$ and $\mathbf{u}_{d}$ is an eigenvector of $\mathbf{\Sigma}-(2d)^{-1}\mathbf{I}$ with large eigenvalue, so that the data $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ will manifest non-zero random signal approximately around this spectral component, it can first complete the orthonormal basis $\mathbf{u}_{1},\dots,\mathbf{u}_{d-1}$ and picks
$\mathbf{A}=\varepsilon^{-1}(d-1)\mathbf{u}_{1}\mathbf{u}_{1}^{\mathsf{T}}+\dots\varepsilon^{-1}(d-1)\mathbf{u}_{d-1}\mathbf{u}_{d-1}^{\mathsf{T}}+(\alpha-\varepsilon)^{-1}\mathbf{u}_{d}\mathbf{u}_{d}^{\mathsf{T}}$ (101)
to outpower Scalar.
Our experiment is set again in the $d=3$ case with $\mathbf{u}_{1}=(1/\sqrt{2},1/\sqrt{2},0)^{\mathsf{T}}$, $\mathbf{u}_{2}=(0,0,1)^{\mathsf{T}}$, $\mathbf{u}_{3}=(1/\sqrt{2},-1/\sqrt{2},0)^{\mathsf{T}}$, and the covariance matrix
$\mathbf{\Sigma}=\left(\begin{matrix}1/6&1/24&0\\ 1/24&1/6&0\\ 0&0&1/6\end{matrix}\right)=\text{null hypothesis}+\left(\begin{matrix}0&1/24&0\\ 1/24&0&0\\ 0&0&0\end{matrix}\right),$ (102)
which indeed satisfies $(\mathrm{tr}\mathbf{\Sigma})^{2}+2\mathrm{tr}(\mathbf{\Sigma}^{2})=\frac{61}{144}<1$. The three sequential tests as before are compared in Figure 2: the smart Matrix that follows (101) with $\varepsilon=\alpha/20$; the naive Matrix with $\mathbf{A}=(d/\alpha)\mathbf{I}$; Scalar, all with $\gamma_{n}=n^{-1/2}$ and $\alpha=0.05$. Besides the unsurprising fact that the smart Matrix is again the most powerful, rejecting the null earlier than the other two, we also observe the phenomenon that, while in the $n=1$ case, the matrix-valued test process with “naive” $\mathbf{A}=(d/\alpha)\mathbf{I}$ is always less powerful than the scalar-valued test process, this is reversed under non-commutativity with large $n$ as naive Matrix rejects slightly earlier than Scalar in our simulation. As also shown in Figure 2, the combination of two auxiliary findings contributes to this phenomenon: ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$ even when observations do not commute; and, while $L_{n}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ in all our previous examples, now $\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ grows slightly faster than $L_{n}^{\textsf{SN}}$. It takes further theoretical investigations to understand why this happens (either in this particular example or in general), but to some extent it hints nonetheless at the possibility that in practice where the effect size is small (requiring large $n$) and matrix observations are non-commuting, there is an inherent advantage using the matrix-valued test process we propose.
Figure 2: Empirical findings with non-commuting i.i.d. random matrices for covariance testing. Left. Comparison of the cumulative distribution functions of rejection times of the sequential tests corresponding to the matrix- and scalar-valued test processes under the same alternative distribution. Middle. Evolution of the spectral distribution of the matrix-valued test process under the alternative. Right. Comparison between the trace of the matrix-valued test process, and the scalar-valued test process under the same alternative distribution.
7 Conclusion
We prove in this paper several novel concentration inequalities for random symmetric matrices. Some of these bounds are tighter than previous ones in the literature via randomization or Loewner order, others enjoy the virtues of validity at stopping times and/or the more general applicability to non-i.i.d. exchangeable matrices. These results shall find applications in, for example, covariance estimation and testing.
We also leave open some questions for future studies to address. These include a deeper understanding of the matrix randomizer $\mathbf{U}$ in our paper, which belongs to the “trace super-uniform” distribution family; tighter $p$-Chebyshev inequalities; further investigation into matrix-valued e-processes; as well as the theoretical explanation of our comparison between matrix-valued test supermartingales and scalar-valued test supermartingales when testing sequentially the same null hypotheses. Our experiments show that matrix-valued test processes are often more powerful in these sequential tests.
Acknowledgement
The authors thank Joel A. Tropp, Lester Mackey, and Tudor Manole for suggestions and discussions, and acknowledge support from NSF grant DMS-2310718.
Appendix A $p$-Chebyshev Inequalities
A.1 The Scalar (and Vector) $p$-Chebyshev Inequality
The following statement shows that for i.i.d. random scalars, the central $p$th moment of the sum grows linearly in the number of scalars just like the variance.
Proposition A.1 (Contraction of the scalar $p$th moment, Lemma 4.2 of Cherapanamjeri et al. [2022]).
Suppose $1\leqslant p\leqslant 2$. Let $X_{1},\dots,X_{n}$ be i.i.d. scalar with mean $\mu=\mathbb{E}X_{1}$ and central $p$th moment $v_{p}=\mathbb{E}|X_{1}-\mu|^{p}$. Then, $\mathbb{E}|X_{1}+\dots+X_{n}-n\mu|^{p}\leqslant 2^{2-p}nv_{p}$.
Wang et al. [2021] generalize the statement to vectors under the martingale difference assumption.
Proposition A.2 (Concentration of the vector $p$th moment, Lemma 7 of Wang et al. [2021]).
Suppose $1\leqslant p\leqslant 2$. Let $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ be $d$-dimensional random vectors adapted to a filtration $\mathcal{F}_{0},\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ with conditional mean $\mu=\mathbb{E}(\mathbf{x}_{i}|\mathcal{F}_{i-1})$ and conditional central $p$th moment upper bound $v_{p}\geqslant\mathbb{E}(\|\mathbf{x}_{i}-\mu\|^{p}|\mathcal{F}_{i-1})$. Then, $\mathbb{E}\|\mathbf{x}_{1}+\dots+\mathbf{x}_{n}-n\mu\|^{p}\leqslant 2^{2-p}d^{1-p/2}nv_{p}$.
[…]
Corollary A.3 (Vector $p$-Chebyshev Inequality).
Suppose $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ satisfy the same assumptions as in Proposition A.2. Then, for any $a>0$,
$\mathbb{P}\left(\left\|\overline{\mathbf{x}}_{n}-\mu\right\|\geqslant a\right)\leqslant\frac{2^{2-p}d^{1-p}v_{p}}{n^{p-1}a^{p}}.$ (103)
A.2 Matrix $p$-Chebyshev Inequalities
The following conjecture is an ideal extension of Propositions A.2 and A.3.
Conjecture A.4 (Matrix $p$-Chebyshev Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and $p$th central moment matrix $\mathbf{V}_{p}:=\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M}))^{p}$. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$, there exists a function $f:[1,2]\times\mathbb{N}\to(0,\infty)$ that grows sublinearly in its second argument, and bounded in its first argument, such that the $p$th central moment contracts,
$\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n}-n\mathbf{M})^{p})\preceq n\cdot f(p,d)\cdot\mathbf{V}_{p};$ (104)
and consequently, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$,
$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{1-p}\cdot f(p,d)\cdot\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (105)
If it holds, the $\mathbf{A}$ above can be randomized into $(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p}$ as well.
We do have the following heavily scalarized and dimension-dependent bound that scales correctly w.r.t. the sample size $n$. We find this bound trivial as it regards the matrix as a mere collection of $d^{2}$ scalars and is strictly looser than the “standard” vector bound (Proposition A.2) applied to the flattened matrix.
Proposition A.5 (Matrix $p$-Chebyshev Spectral Inequality).
Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d. random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and $p$th spectral central moment $v_{p}:=\mathbb{E}\|\mathbf{X}_{1}-\mathbf{M}\|^{p}$. Then,
$\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}-n\mathbf{M}\|^{p}\leqslant 2^{2-p}d^{2-p/2}\cdot n\cdot v_{p}.$ (106)
Proof.
Without loss of generality suppose $\mathbf{M}=0$. Denote by $\|\cdot\|_{L_{p}}$ the entrywise matrix $p$-norm and $\|\cdot\|_{\mathrm{F}}=\|\cdot\|_{L_{2}}$ the Frobenius norm. Recall that $\|\mathbf{A}\|\leqslant\|\mathbf{A}\|_{\mathrm{F}}\leqslant\sqrt{d}\|\mathbf{A}\|$ and $\|\mathbf{A}\|_{\mathrm{F}}\leqslant\|\mathbf{A}\|_{L_{p}}\leqslant d^{\frac{2}{p}-1}\|\mathbf{A}\|_{\mathrm{F}}$. Then, $\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}\|^{p}$ is upper bounded by
$\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}\|_{L_{p}}^{p}=\mathbb{E}\left[\sum_{1\leqslant j,k\leqslant d}\left|\sum_{i=1}^{n}X_{i(jk)}\right|^{p}\right]=\sum_{1\leqslant j,k\leqslant d}\mathbb{E}\left|\sum_{i=1}^{n}X_{i(jk)}\right|^{p}$ (107)
$\stackrel{{\scriptstyle(*)}}{{\leqslant}}\,$ $\sum_{1\leqslant j,k\leqslant d}2^{2-p}n\mathbb{E}\left|X_{1(jk)}\right|^{p}=2^{2-p}n\mathbb{E}\|\mathbf{X}_{1}\|_{L_{p}}^{p}\leqslant 2^{2-p}d^{2-p}\cdot n\cdot\mathbb{E}\|\mathbf{X}_{1}\|^{p}_{\mathrm{F}}$ (108)
[…]
where the inequality (*)(*)( * ) comes from the contraction of scalar $p$th moment (Proposition A.1). ∎
Appendix B Additional Lemmas and Omitted Proofs
B.1 Proofs of Statements in Table 1
Proof.
For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a direct Taylor expansion argument. Since the matrix log is monotone (Fact 2), $\log\mathbf{G}(\gamma/n)\preceq\gamma^{2}\mathbf{C}^{2}/2n^{2}$. Finally, invoking the monotonicity of $\mathbf{A}\mapsto\mathrm{tr}\exp\mathbf{A}$ (Fact 3) we see that $\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$. Bennett I follows from a similar reasoning applied to the Bennett MGF bound [Tropp, 2012, Lemma 6.7]. For Bennett II, we can apply the variance bound at last: $\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{Var}(\mathbf{X}_{1})})\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$. The uni-Gaussian MGF bound is also proved by Tropp [2015, Lemma 4.6.2] via a direct Taylor expansion argument. For symmetric Hoeffding, introducing an independent $R\sim\operatorname{rad}_{1/2}$,
$\log\mathbf{G}(\gamma)=\log\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}=\log\mathbb{E}\mathrm{e}^{\gamma R(\mathbf{X}_{1}-\mathbf{M})}\preceq\gamma^{2}\mathbf{B}/2$ (110)
due to [Tropp, 2012, Lemma 7.7], leading to the corresponding upper bound on $\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$ as before. ∎
B.2 Dominated Integrability Lemma
It is well-known that if $X$ and $Y$ are scalar random variables such that $c\leqslant X\leqslant Y$ almost surely for some constant $c$ and that $\mathbb{E}|Y|<\infty$, it follows that $\mathbb{E}|X|<\infty$ as well, and $\mathbb{E}X\leqslant\mathbb{E}Y$. This type of “implied integrability” appears frequently in scalar concentration bounds. Let us prove its symmetric matrix extension for the sake of self-containedness.
Lemma B.1 (Dominated Integrability).
Let $\mathbf{X}$ and $\mathbf{Y}$ be $\mathcal{S}_{d}^{[c,\infty)}$-valued random matrices for some $c\in\mathbb{R}$ such that $\mathbf{X}\preceq\mathbf{Y}$ almost surely. Further, suppose $\mathbb{E}\mathbf{Y}$ exists. Then, so does $\mathbb{E}\mathbf{X}$ and $\mathbb{E}\mathbf{X}\preceq\mathbb{E}\mathbf{Y}$.
Proof.
Let us prove that each element $X_{ij}$ of the random matrix $\mathbf{X}$ is integrable. Note that for any deterministic $\mathbf{v}\in\mathbb{R}^{d}$, $\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v}\leqslant\mathbf{v}^{\mathsf{T}}\mathbf{Y}\mathbf{v}$ almost surely. First, taking $\mathbf{v}=(0,\dots,0,1,0,\dots 0)^{\mathsf{T}}$, we have
$c\leqslant X_{jj}\leqslant Y_{jj}\quad\text{almost surely},$ (111)
concluding that the diagonal element $X_{jj}$ must be integrable (since $Y_{jj}$ is). Next, taking $\mathbf{v}=(0,\dots,0,1,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$, we have
$2c\leqslant 2X_{ij}+X_{ii}+X_{jj}\leqslant 2Y_{ij}+Y_{ii}+Y_{jj}\quad\text{almost surely},$ (112)
concluding that $2X_{ij}+X_{ii}+X_{jj}$ must be integrable (since $2Y_{ij}+Y_{ii}+Y_{jj}$ is). Therefore, the off-diagonal element $X_{ij}$ is integrable since $X_{ii}$ and $X_{jj}$ are.
Now that we have established the existence of $\mathbb{E}\mathbf{X}$, it is clear that $\mathbb{E}\mathbf{X}\preceq\mathbb{E}\mathbf{Y}$ since for any $\mathbf{v}\in\mathbb{R}^{d}$, $\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{X})\mathbf{v}=\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y})\mathbf{v}$. ∎
B.3 Proofs of Self-Normalization Statements
The following transfer rule [Tropp, 2012, Equation 2.2] shall commonly be used: suppose $I\subseteq\mathbb{R}$ and $f,g:I\to\mathbb{R}$ satisfies $f(x)\leqslant g(x)$, then, since the function $g-f:I\to\mathbb{R}_{\geqslant 0}$ identifies canonically with a matrix-to-matrix function $g-f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{+}$, it follows that $f(\mathbf{X})\preceq g(\mathbf{X})$ for any $\mathbf{X}\in\mathcal{S}_{d}^{I}$.
Proof of Example 4.8.
Let us replicate the key steps of the proof of Lemma 3(f) in Howard et al. [2020]. Recalling from Delyon [2009, Proposition 12] that $\exp(x-x^{2}/6)\leqslant 1+x+x^{2}/3$ for all $x\in\mathbb{R}$, we have, due to the transfer rule
$\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\preceq 1+\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})+\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/3.$ (113)
This allows us to use Lemma B.1,
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\mid\mathcal{F}_{n-1}\right)$ (114)
[…]
=\ = $1+\mathbb{E}\left(\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/3\mid\mathcal{F}_{n-1}\right)\preceq 1+\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (116)
where in the last step we use the transfer rule again with $1+x\leqslant\exp(x)$ for all $x\in\mathbb{R}$. ∎
Proof of Example 4.9.
Let us replicate the key steps of the proof of Lemma 3(d) in Howard et al. [2020], which is also a self-normalization argument. Due to the symmetry of $\mathbf{X}_{n}-\mathbf{M}_{n}$, we introduce a Rademacher random variable $\epsilon$ independent from $\mathcal{F}_{n}$,
$\mathbb{E}\left(\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)$ (117)
$=$ $\mathbb{E}\left(\exp\left(\gamma_{n}\epsilon(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)$ (118)
$=$ $\mathbb{E}\left\{\mathbb{E}\left(\exp\left(\gamma_{n}\epsilon(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n}\right)\middle|\mathcal{F}_{n-1}\right\}$ (119)
$=$ $\mathbb{E}\left\{\frac{\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)+\exp\left(-\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)}{2}\middle|\mathcal{F}_{n-1}\right\}.$ (120)
Note that the function $f(x)=\frac{\exp(x-x^{2}/2)+\exp(-x-x^{2}/2)}{2}$ satisfies $f(x)\leqslant 1$ for all $x\in\mathbb{R}$, we see from the transfer rule and Lemma B.1 that
$\mathbb{E}\left(\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)=\mathbb{E}\left(f(\mathbf{X}_{n}-\mathbf{M}_{n})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbf{I}.$ (121)
B.4 Proofs of Empirical Bernstein Statements
Proof of Proposition 4.10.
Recall that $\psi_{E}(\gamma)=-\log(1-\gamma)-\gamma$. An inequality by Fan et al. [2015] quoted by Howard et al. [2021, Appendix A.8] states that, for all $10\leqslant\gamma<$ 1 and $\xi\geqslant-1$,
$\exp(\gamma\xi-\psi_{E}(\gamma)\xi^{2})\leqslant 1+\gamma\xi.$ (122)
Since $\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}\in\mathcal{S}_{d}^{[-1,\infty)}$, we can apply the transfer rule mentioned in Section B.3, replacing the scalar $\xi$ above by the matrix $\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}$, and plugging in $\gamma=\gamma_{n}\in(0,1)$,
$\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\preceq 1+\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}).$ (123)
Lemma B.1 then guarantees the integrability of the left hand side, and that
$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbb{E}\left(1+\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\middle|\mathcal{F}_{n-1}\right)$ (124)
$=$ $1+\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})\preceq\exp(\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (125)
where in the final step we use the transfer rule again with $1+x\leqslant\exp(x)$ for all $x\in\mathbb{R}$. This concludes the proof. ∎
Proof of Theorem 5.6.
We apply Lemma 5.1 with $\mathbf{Z}_{n}=\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})$, $\mathbf{C}_{n}=\gamma_{n}(\widehat{\mathbf{X}}_{n}-\mathbf{M}_{n})+\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}$, and $\mathbf{C}_{n}^{\prime}=\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})$ to see that
$L_{n}^{\mathsf{EB}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\psi_{E}(\gamma_{i})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)$ (126)
is a supermartingale, which upper bounds
$\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)\right\}.$ (127)
Applying (UVI),
$\mathbb{P}\left(\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)\geqslant\log(dU/\alpha)\right)\leqslant\alpha,$ (128)
and the desired result follows from rearranging. ∎
Lemma B.2.
Let $\psi_{E}(x)=-\log(1-x)-x$. When $0\leqslant x\leqslant\sqrt{2/5}$, $\psi_{E}(x)\leqslant x^{2}$.
Proof.
Let $g(x)=\psi_{E}(x)-x^{2}$. The claim follows from $g^{\prime\prime}(x)=(1-x)^{-2}-2\geqslant 0$ for $x\in[0,\sqrt{2/5}]$, and $g(0)=0$, $g(\sqrt{2/5})<0$. ∎
Proof of Corollary 5.7.
First, it is straightforward that $\lambda_{\max}({\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}})\geqslant-1$ for every $i=1,\dots,n$ since both $\mathbf{X}_{i}$ and $\overline{\mathbf{X}}_{i-1}$ take values in $\mathcal{S}_{d}^{[0,1]}$, so Theorem 5.6 is applicable. Let us prove the two claims about the deviation bound $D_{n}^{\mathsf{EB}}$ under $\gamma_{i}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}$. Recall that
$\overline{\mathbf{V}}_{0}=0,\quad\overline{\mathbf{V}}_{k}=\frac{1}{k}\sum_{i=1}^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{k})^{2},\quad\overline{v}_{k}=\lambda_{\max}(\overline{\mathbf{V}}_{k})\vee\frac{5\log(d/\alpha)}{n},$ (129)
$\widetilde{s}_{n}=\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}^{2})}{\overline{v}_{i-1}}\right).$ (130)
First, via our definition, $\overline{v}_{i}\geqslant\frac{5\log(d/\alpha)}{n}$, therefore $\gamma_{i}\leqslant\sqrt{2/5}$. So Lemma B.2 implies that $\psi_{E}(\gamma_{i})\leqslant\gamma_{i}^{2}$. Therefore,
$D_{n}^{\mathsf{EB}}$ $=\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}}$ (131)
$\leqslant\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}}$ (132)
$=\sqrt{\frac{\log(d/\alpha)}{2n}}\frac{1+2\widetilde{s}_{n}}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}.$ (133)
Then, let us show that $\lim_{n\to\infty}D_{n}^{\mathsf{EB}}/D_{n}^{\mathsf{B}}$ almost surely. Via the boundedness of $\mathbf{X}_{1}$ and the strong law of large numbers, we can see the following limits hold almost surely (cf. Waudby-Smith and Ramdas [2023, Lemmas 4-6]):
$\lim_{k\to\infty}\overline{\mathbf{X}}_{k}=\mathbf{M},\quad\lim_{k\to\infty}\overline{\mathbf{V}}_{k}=\mathbf{V}=\mathbf{Var}(\mathbf{X}_{1}),$ (134)
$\lim_{k\to\infty}\overline{v}_{k}=\lambda_{\max}(\mathbf{V}),\quad\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}=\mathbf{V},$ (135)
$\lim_{n\to\infty}\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}\right)=\lambda_{\max}(\mathbf{V}),$ (136)
$\lim_{n\to\infty}\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}{\overline{v}_{i-1}}\right)=1.$ (137)
Therefore, we have, via the expansion $\psi_{E}(x)=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$,
$\limsup_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}$ (138)
$=$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (139)
$\leqslant$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{2}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{2}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (140)
$+$ $\sum_{k=3}^{\infty}\limsup_{n\to\infty}\frac{\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{k}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (141)
$=$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{2}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{2}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (142)
$=$ $\limsup_{n\to\infty}\sqrt{\frac{\log(d/\alpha)}{2}}\frac{\left(1+\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}{\overline{v}_{i-1}}\right)\right)}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}$ (143)
$=$ $\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}.$ (144)
Similarly, one can show that $\liminf_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}\geqslant\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}$, concluding the proof. ∎
B.5 Proof of Lemma 5.1
Proof.
Due to the monotonicity of $\log$ (Fact 2), the condition (68) implies
$\log\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\mathbf{C}_{n}^{\prime}.$ (145)
Now recall Lieb’s concavity theorem [Lieb, 1973]: for any $\mathbf{H}\in\mathcal{S}_{d}$, the map $\mathbf{X}\mapsto\mathrm{tr}\exp(\mathbf{H}+\log\mathbf{X})$ ($\mathcal{S}_{d}^{++}\to(0,\infty)$) is concave. Therefore,
$\mathbb{E}(L_{n}|\mathcal{F}_{n-1})=$ $\mathbb{E}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{n-1}\mathbf{Z}_{i}-\sum_{i=1}^{n-1}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})-\mathbf{C}_{n}^{\prime}+\log\mathrm{e}^{\mathbf{Z}_{n}-\mathbf{C}_{n}}\right)\middle|\mathcal{F}_{n-1}\right)$ (146)
(Jensen’s inequality)
[…]
(by (145) and Fact 3)
[…]
concluding the proof that $\{L_{n}\}$ is a supermartingale. Finally, observe that
$L_{n}$ $=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)$ (149)
$\geqslant\mathrm{tr}\exp\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\mathbf{I}\right)$ (150)
$\geqslant\lambda_{\max}\exp\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\mathbf{I}\right)$ (151)
$=\exp\lambda_{\max}\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\mathbf{I}\right)$ (152)
$=\exp\left(\lambda_{\max}\left(\sum_{i=1}^{n}\mathbf{Z}_{i}\right)-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\right),$ (153)
concluding the proof. ∎
Appendix C Further Notes on Matrix-Valued Martingales
C.1 Doob’s $L^{p}$ Inequality on Trace
Applying an operator-monotone and concave function to a matrix supermartingale, we get another matrix supermartingale. In particular,
Lemma C.1.
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale on $\{\mathcal{F}_{n}\}$ and $p\in[1,\infty)$. Then, $\{\mathbf{M}_{n}^{1/p}\}$ and $\{\log\mathbf{M}_{n}\}$ are also supermartingales.
Proof.
The function $\mathbf{X}\mapsto\mathbf{X}^{1/p}$ is operator-monotone and concave. So
$\mathbb{E}(\mathbf{M}_{n}^{1/p}|\mathcal{F}_{n-1})\preceq\mathbb{E}(\mathbf{M}_{n}|\mathcal{F}_{n-1})^{1/p}\preceq\mathbf{M}_{n-1}^{1/p}.$ (154)
The same proof applies for the logarithm as well. ∎
Fact 3 and Fact 4, on the other hand, convert a matrix-valued (super/sub)martingale to a scalar-valued one. For example,
Lemma C.2.
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then, $\{\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})^{p}\}$ is a $\mathbb{R}_{\geqslant 0}$-valued submartingale.
Proof.
Using Fact 4,
$\mathbb{E}(\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})^{p}|\mathcal{F}_{n-1})=\mathrm{tr}\mathbb{E}((\operatorname{abs}\mathbf{M}_{n})^{p}|\mathcal{F}_{n-1})\geqslant\mathrm{tr}(\operatorname{abs}\mathbb{E}(\mathbf{M}_{n}|\mathcal{F}_{n-1}))^{p}.$ (155)
If $\{\mathbf{M}_{n}\}$ is a martingale, the right hand side above is $\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n-1})^{p}$. If $\mathbf{M}_{n}\succeq 0$ is a submartingale, then we have by Fact 3,
$\mathrm{tr}(\operatorname{abs}\mathbb{E}(\mathbf{M}_{n}|\mathcal{F}_{n-1}))^{p}\geqslant\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n-1})^{p},$ (156)
since the function $x\mapsto|x|^{p}$ is monotone on $x\geqslant 0$. ∎
Proposition C.3 (Trace Doob’s $L^{p}$ Inequality).
Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then,
$\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{\mathrm{op}}^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}(\mathrm{tr}\operatorname{abs}\mathbf{M}_{n})^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\mathrm{tr}\operatorname{abs}\mathbf{M}_{N})^{p}\leqslant\left(\frac{pd}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|_{\mathrm{op}}^{p}.$ (157)
Proof.
The second inequality is directly from applying Doob’s $L^{p}$ inequality on the nonnegative submartingale $\{\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})\}$. The first and the third $\leqslant$’s are both trivial. ∎
Appendix D Non-symmetric, Non-square Matrices
We finally remark on the extension of our methods to non-symmetric and possibly non-square matrices. Denote by $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$ the space of all $d_{1}\times d_{2}$ real-valued matrices, and we consider the random observations $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ now as taking values in $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$. We can employ the classical idea of dilations (see e.g. [Tropp, 2015, Section 2.1.16]). Consider the linear map
$D:\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})\to\mathcal{S}_{d_{1}+d_{2}},\quad D(\mathbf{X})=\left(\begin{matrix}0&\mathbf{X}\\ \mathbf{X}^{\mathsf{T}}&0\end{matrix}\right),$ (158)
and we can apply our previously developed bounds or tests on $\{D(\mathbf{X}_{n})\}$, which shall in turn provide concentration inequalities for $\{\mathbf{X}_{n}\}$ themselves: either via the relation $\|D(\mathbf{X})\|=\|\mathbf{X}\|$ with the linearity of $D$ to obtain spectral bounds for $\mathbf{X}_{1},\mathbf{X}_{2},\dots$, or one may directly use the resulting (scalar or matrix) e-process as the test process.
In Table 2, we see that quadratic assumptions on the possibly non-square random matrix $\mathbf{X}_{1}$ translate directly to those on the symmetric random matrix $D(\mathbf{X}_{1})$.
Assumption on $\mathbf{X}_{1}$ Assumption on $D(\mathbf{X}_{1})$
$\mathbf{X}_{1}^{\mathsf{T}}\mathbf{X}_{1}\preceq\mathbf{B}_{1}$ and $\mathbf{X}_{1}\mathbf{X}_{1}^{\mathsf{T}}\preceq\mathbf{B}_{2}$ $D^{2}(\mathbf{X}_{1})\preceq\left(\begin{matrix}\mathbf{B}_{1}&\\ &\mathbf{B}_{2}\end{matrix}\right)$
$\mathbf{Cov}(\mathbf{X}_{1}^{\mathsf{T}})\preceq\mathbf{V}_{1}$ and $\mathbf{Cov}(\mathbf{X}_{1})\preceq\mathbf{V}_{2}$ $\mathbf{Var}(D(\mathbf{X}_{1}))\preceq\left(\begin{matrix}\mathbf{V}_{1}&\\ &\mathbf{V}_{2}\end{matrix}\right)$
Table 2: Distribution assumptions on $\{\mathbf{X}_{n}\}$ imply those on $\{D(\mathbf{X}_{n})\}$. Here, the covariance of a random matrix is defined as $\mathbf{Cov}(\mathbf{X})=\mathbb{E}(\mathbf{X}-\mathbb{E}\mathbf{X})(\mathbf{X}-\mathbb{E}\mathbf{X})^{\mathsf{T}}$.
For example, the self-normalized matrix supermartingale (Example 4.8) can be generalized as follows.
Proposition D.1.
Let $\{\mathbf{X}_{n}\}$ be $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional covariance upper bounds $\mathbf{Cov}(\mathbf{X}_{n}^{\mathsf{T}}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{1n}$, $\mathbf{Cov}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{2n}$ that are predictable. For any predictable positive scalars $\{\gamma_{n}\}$, let
$\scriptsize\mathbf{E}_{n}=\left(\begin{matrix}-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})(\mathbf{X}_{n}-\mathbf{M}_{n})^{\mathsf{T}}/6&\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})\\ \gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})^{\mathsf{T}}&-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{\mathsf{T}}(\mathbf{X}_{n}-\mathbf{M}_{n})/6\end{matrix}\right),\;\mathbf{F}_{n}=-\frac{\gamma_{n}^{2}}{3}\left(\begin{matrix}\mathbf{V}_{1n}&\\ &\mathbf{V}_{2n}\end{matrix}\right),$ (159)
the process
$\mathbf{Y}_{n}=\mathrm{e}^{\mathbf{F}_{1}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{2}/2}\dots\mathrm{e}^{\mathbf{F}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{F}_{n}/2}\dots\mathrm{e}^{\mathbf{E}_{2}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{1}/2}.$ (160)
is a $\mathcal{S}_{d_{1}+d_{2}}^{+}$-valued supermartingale.

### v4

@@
-Positive Semidefinite Supermartingales and Randomized Matrix Concentration Inequalities
+Positive Semidefinite Matrix Supermartingales
 Abstract
-We present new concentration inequalities for either martingale dependent or exchangeable random symmetric matrices under a variety of tail conditions, encompassing now-standard Chernoff bounds to self-normalized heavy-tailed settings.
-These inequalities are often randomized in a way that renders them strictly tighter than existing deterministic results in the literature, are typically expressed in the Loewner order, and are sometimes valid at arbitrary data-dependent stopping times.
-Along the way, we explore the theory of positive semidefinite supermartingales and maximal inequalities, a natural matrix analog of scalar nonnegative supermartingales that is potentially of independent interest.
+We explore the asymptotic convergence and nonasymptotic maximal inequalities of supermartingales and backward submartingales in the space of positive semidefinite matrices.
+These are natural matrix analogs of scalar nonnegative supermartingales and backward nonnegative submartingales, whose convergence and maximal inequalities are the theoretical foundations for a wide and ever-growing body of results in statistics, econometrics, and theoretical computer science.
+Our results lead to new concentration inequalities for either martingale dependent or exchangeable random symmetric matrices under a variety of tail conditions, encompassing now-standard Chernoff bounds to self-normalized heavy-tailed settings.
+Further, these inequalities are usually expressed in the Loewner order, are sometimes valid simultaneously for all sample sizes or at an arbitrary data-dependent stopping time, and can often be tightened via an external randomization factor.
 1 Introduction
-Bounding the tail probability of the sum of independent and identically distributed (i.i.d.) real-valued random variables is one of the most crucial tasks in the nonasymptotic theory of statistics, for which a copious collection of methods has flourished [Boucheron et al., 2013].
-This paper concerns concentration inequalities for sums of random symmetric matrices under martingale dependence or exchangeability (both generalizing the i.i.d.
-assumption).
-Notable previous results include the study of Rademacher and random graph adjacency matrices by Oliveira [2009, 2010]; those in a series of seminal papers by Tropp [2012, 2015], who famously ports the techniques of cumulant generating functions (CGFs) and Chernoff-type bounds, which are central in scalar concentration, into the matrix regime; some new bounds by Mackey et al.
-[2014] using exchangeability; and a recent study on “anytime-valid” concentration inequalities by Howard et al.
-[2020, 2021] which improves and unifies several matrix Chernoff-type bounds.
-We return to several of these works later in the current paper.
-To set the stage for our exposition, we quote a celebrated result by Mackey et al.
-[2014, Corollary 4.2] (and also Tropp [2012, Theorem 1.3]) which states that, if $d\times d$ random i.i.d.
-symmetric matrices $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ satisfy $(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})^{2}\preceq\mathbf{B}$ (meaning that $\mathbf{B}-(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})^{2}$ is positive semidefinite) almost surely, then
-$\mathbb{P}\left(\lambda_{\max}\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}-\mathbb{E}\mathbf{X}_{1}\right)\geqslant\sqrt{\frac{2\lambda_{\max}(\mathbf{B})\log(d/\alpha)}{n}}\right)\leqslant\alpha.$ (MHI)
-Above, $\lambda_{\max}(\,\cdot\,)$ refers to the largest eigenvalue of a symmetric matrix, a typical scalarization of matrix-valued tails.
-When $d=1$, it recovers, with the same optimal constant, the scalar Hoeffding’s inequality as a special case; hence its name “matrix Hoeffding’s inequality”.
-In this paper, we present numerous new techniques to strengthen this type of matrix concentration inequalities. Each of our new bounds enjoys a combination of some of the following virtues:
+Let $\mathbf{Y}_{0},\mathbf{Y}_{1},\dots$ be a sequence of random symmetric $d\times d$ matrices adapted to a filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
+The process $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ is called a (matrix-valued) supermartingale if
+for $\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}.$ (1)
+Here, crucially, the order $\preceq$ is the Loewner order, where $\mathbf{A}\preceq\mathbf{B}$ if and only if $\mathbf{B}-\mathbf{A}$ is a positive semidefinite matrix.
+This introduces the rich spectral theory of matrices to the familiar $d=1$ scalar special case as well as the matrix martingale special case where the $\preceq$” in (1) takes the entry-wise “$=$”.
+Similarly, if the matrices $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ are adapted instead to a backward filtration $\{\mathcal{G}_{n}\}_{n\geqslant 0}$, that is, $\mathcal{G}_{0}\supseteq\mathcal{G}_{1}\supseteq\dots$, it is called a (matrix-valued) backward submartingale if
+for $\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n-1}|\mathcal{G}_{n})\succeq\mathbf{Y}_{n}.$ (2)
+In this paper, we primarily study matrix-valued supermartingales and backward submartingales that are almost surely positive semidefinite (PSD).
+These matrix-valued stochastic processes are natural and non-trivial generalizations of scalar nonnegative supermartingales and nonnegative backward submartingales, studied extensively by the probability and statistics literature.
+In particular, as we shall review soon in Section 2.2, a scalar nonnegative supermartingale or backward submartingale $\{Y_{n}\}$ satisfies Ville’s inequality
+$\mathbb{P}(\exists n,\ Y_{n}\geqslant a)\leqslant(\mathbb{E}Y_{0})a^{-1},$ (3)
+which is interconnected with various results on the asymptotic and non-asymptotic behavior of these classes of processes, and has led to numerous methods in sequential statistics.
+We extend all these results and applications from scalars to matrices.
+In particular, we present the matrix Ville’s inequality
+$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1})$ (4)
+for both classes of processes, their numerous equivalent or strengthened forms, and a variety of statistical applications.
+A list of major contributions of this paper is as follows.
   1. 1.
-    Anytime validity. We derive new inequalities that hold not only for a fixed number ($n$) of matrices but also a stopping time number of them [Howard et al., 2020].
+    The convergence and optional stopping theorems, and the Ville’s inequality for PSD matrix supermartingales (Theorems 3.2, 3.3 and 3.4) that generalize the corresponding scalar results.
   2. 2.
-    Randomization.
-We losslessly tighten many of the previously known inequalities by randomly decreasing the tail threshold (decreasing the $a$ in (MHI) on the left hand side but not the right hand side), extending a technique that was recently explored by Ramdas and Manole [2023] for scalars.
+    Sequential testing procedure with a matrix test supermartingale (Algorithm 1) that can be applied to the meaning testing of bounded (Examples 3.7 and 3.10), symmetric (Example 3.9), or heavy-tailed (Examples 3.8 and 3.12) martingale-dependent matrices.
   3. 3.
-    Exchangeability. We go beyond the i.i.d. or martingale difference assumptions on random matrices and derive bounds that allow them to just be exchangeable instead.
+    Matrix e-values and e-processes (Definitions 3.13 and 3.14), generalizing the corresponding scalar concepts that have recently been extensively studied.
   4. 4.
-    Tighter (or no) scalarization.
-In some bounds we use the trace functional $\mathrm{tr}(\,\cdot\,)$ which is larger than the usual $\lambda_{\max}(\,\cdot\,)$ (e.g.
-in (MHI)), and introduce some bounds completely in the matrix order $\preceq$ without any scalarization, which is partly achieved by studying the maximal inequalities of matrix-valued sub- and super-martingales.
+    The matrix Ville’s inequality for backward PSD matrix submartingales (Theorem 4.3), leading to the exchangeable matrix Chebyshev’s inequality (Theorem 4.5).
   5. 5.
-    Heavy tails. Instead of assuming the matrices to be bounded or (sub)Gaussian, we sometimes allow them to only have a finite variance or $p$th moment, or be symmetric.
-To showcase these virtues, let us pick two of the bounds we shall prove in the exposition that follows. Under the assumptions of the (MHI) above, our Theorem 5.5 states that
-$\mathbb{P}\left(\lambda_{\max}\left(\frac{\gamma_{1}\mathbf{X}_{1}+\dots+\gamma_{\tau}\mathbf{X}_{\tau}}{\gamma_{1}+\dots+\gamma_{\tau}}-\mathbb{E}\mathbf{X}_{1}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{2}\lambda_{\max}(\mathbf{B})\sum_{i=1}^{\tau}\gamma_{i}^{2}}{\gamma_{1}+\dots+\gamma_{\tau}}\right)\leqslant\alpha$ (1)
-for any predictable positive tuning parameters $\{\gamma_{n}\}$, stopping time $\tau$, and an independent random variable $U\sim\text{Unif}_{(0,1)}$.
-The above equips (MHI) with virtues 1 and 2 without compromises, since replacing $U$ with 1 and taking $\tau=n$, $\gamma_{1}=\dots=\gamma_{n}=\sqrt{\frac{2\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B})}}$ in above, one obtains the exact deviation bound $\sqrt{\frac{2\log(d/\alpha)\lambda_{\max}(\mathbf{B})}{n}}$ as in (MHI).
-A bound of very different flavor is that of our Theorem 4.17, which states that for exchangeable matrices $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ with variance $\mathbf{V}$,
-$\mathbb{P}\left(\exists n\geqslant 1:\operatorname{abs}\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}-\mathbb{E}\mathbf{X}_{1}\right)\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ (2)
-for any positive definite $\mathbf{A}$.
-This bound enjoys virtues 3, 4, and 5, sequentializing the matrix Chebyshev inequality $\mathbb{P}\left(\operatorname{abs}\left(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1}\right)\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ by Ahlswede and Winter [2002, Appendix C] to an infinite time horizon.
-Outline.
+    The randomized matrix Markov’s and Ville’s inequalities (Theorems 5.3 and 5.5).
+  6. 6.
+    A sequential mean testing example with the matrix supermartingale outperforming the scalar supermartingale (Example 6.5).
 The rest of the paper is organized as follows.
-In Section 2, we briefly review the technical background (and challenges) of matrix concentration bounds as well as a collection of classical and recent scalar bounds that anticipate our matrix extensions.
-We present in Section 3 our first kind of bounds, derived via randomized matrix Markov, Chebyshev, and Chernoff inequalities.
-We develop and apply the theory of matrix-valued supermartingales and backward submartingales in Section 4, further obtaining bounds for i.i.d.
-and exchangeable matrices.
-Finally, in Section 5, we revisit scalar supermartingales (and the scalar processes they dominate) constructed via matrices which yield some other randomized bounds as well, concluding with a discussion on the respective advantages of matrix- and scalar- based sequential tests in both theoretical and empirical dimensions.
-2 Preliminaries
+We review in Section 2 the technical background of matrix analysis, sequential statistics with scalar processes in the abovementioned classes, as well as matrix concentration inequalities in the literature.
+We study forward PSD matrix supermartingales, their Ville’s inequality and applications in Section 3.
+We then study backward PSD matrix submartingales, their Ville’s inequality and applications in Section 4.
+In Section 5 we provide several perspectives comparing our inherent matrix approach to some existing scalar-born or scalarized results, and discuss a randomized improvement that is further developed in Appendix A.
+An interesting empirical study is presented in Section 6, including a problem where our matrix method yields higher out-of-the-box power compared to the previous scalar method.
+Omitted proofs are found in Appendix B; some further references and independent theoretical results in Appendix C.
+2 Preliminaries and Related Work
 2.1 Matrix Analysis
 Let $\mathcal{S}_{d}$ denote the set of all $d\times d$ real-valued symmetric matrices, which is the only class of matrices considered in this paper.
-[^1] We only briefly discuss the case with non-symmetric and non-square matrices in Appendix D.
+[^1] We only briefly discuss the case with non-symmetric and non-square matrices in Section C.3.
 These matrices are denoted by bold upper-case letters $\mathbf{A},\mathbf{B}$, etc; whereas $d$ dimensional (column) vectors by bold lower-case letters $\mathbf{u},\mathbf{v}$, etc.
@@
 $\mathcal{S}_{d}^{[0,\infty)}$, the set of positive semidefinite and $\mathcal{S}_{d}^{(0,\infty)}$, the set of positive definite matrices are simply denoted by $\mathcal{S}_{d}^{+}$ and $\mathcal{S}_{d}^{++}$ respectively.
-The Loewner partial order is denoted $\preceq$, where $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite, and $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite.
+We consider random matrices taking values in $\mathcal{S}_{d}$.
+It is easy to see that if $\mathbf{X}$ is an integrable random matrix taking values in $\mathcal{S}_{d}^{I}$ where $I$ is an interval, any conditional expectation $\mathbb{E}(\mathbf{X}|\mathcal{A})$ also takes values in $\mathcal{S}_{d}^{I}$.
+The Loewner partial order $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite, and $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite.
 Any function $f:I\to J$ (where $I,J\subseteq\mathbb{R}$) is identified canonically with a function $f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$ in the following way: for a diagonal matrix $\mathbf{\Sigma}\in\mathcal{S}_{d}^{I}$, $f(\mathbf{\Sigma})$ is obtained by applying $f$ entrywise; otherwise it is obtained by applying $f$ after orthogonal diagonalization, $f(\mathbf{U}^{\mathsf{T}}\mathbf{\Sigma}\mathbf{U})=\mathbf{U}^{\mathsf{T}}f(\mathbf{\Sigma})\mathbf{U}$.
@@
 $\operatorname{abs}:\mathcal{S}_{d}\to\mathcal{S}^{+}_{d}$, and powers $\mathbf{X}\mapsto\mathbf{X}^{k}$ for any $k\in\mathbb{R}$.
-Take heed that usually $\exp(\mathbf{A}+\mathbf{B})\neq\exp\mathbf{A}+\exp\mathbf{B}$ and $\log\mathbf{A}\mathbf{B}\neq\log\mathbf{A}+\log\mathbf{B}$, but a trace equality holds on the latter:
-Fact 1 (Trace Logarithm Identity).
-$\mathrm{tr}\log\mathbf{A}\mathbf{B}=\mathrm{tr}\log\mathbf{A}+\mathrm{tr}\log\mathbf{B}$.
 […]
-It is rare to find matrix-to-matrix functions that are both monotone (i.e.
+It is rare to find matrix-to-matrix functions that are monotone (i.e.
 preserving the $\preceq$ order) and convex (i.e.
 $f$ such that $f(\sum\lambda_{i}\mathbf{X}_{i})\preceq\sum{\lambda_{i}}f(\mathbf{X}_{i})$ if $\sum\lambda_{i}=1$ and each $\lambda_{i}\geqslant 0$; concavity is defined similarly).
-Scalar-to-scalar monotonicity ($I\to J$) does not imply matrix-to-matrix monotonicity ($\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$), and the same holds for convexity.
+Scalar-to-scalar monotonicity ($I\to J$) does not imply matrix-to-matrix monotonicity ($\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$), and the same with convexity.
 For example, the matrix square $\mathbf{X}\mapsto\mathbf{X}^{2}$ is not monotone on the domain $\mathcal{S}_{d}^{+}$ and the matrix 4th power $\mathbf{X}\mapsto\mathbf{X}^{4}$ is neither monotone nor convex.
-The following are some of the few well-known nontrivial “operator-monotone” and “operator-convex” functions in the matrix analysis literature (cf.
+The following are some of the few well-known nontrivial “operator monotone” and “operator convex” functions in the matrix analysis literature (cf.
 Bhatia [2013, Chapter V]).
-Fact 2 (Operator-monotone and -concave Functions).
+Fact 1 (Operator Monotone and Convex Functions).
 The matrix logarithm $\mathbf{X}\mapsto\log\mathbf{X}$ ($\mathcal{S}_{d}^{++}\to\mathcal{S}_{d}$) is monotone and concave.
@@
 For $k\in(1,2]$, the matrix power $\mathbf{X}\mapsto\mathbf{X}^{k}$ ($\mathcal{S}_{d}^{+}\to\mathcal{S}_{d}^{+}$ if $k\neq 2$; $\mathcal{S}_{d}\to\mathcal{S}_{d}^{+}$ if $k=2$) is convex.
-[…]
-Fact 3 (Trace Monotonicity).
+With trace, however, real-to-real functions translate into matrix-to-real functions with monotonicity and convexity intact. That is,
+Fact 2 (Trace Monotonicity).
 Let $f:I\to\mathbb{R}$ be monotone increasing ($I\subseteq\mathbb{R}$). Then so is $\mathrm{tr}\circ f:\mathcal{S}_{d}^{I}\to\mathbb{R}$, i.e.,
 $\mathbf{A}\preceq\mathbf{B}\implies\mathrm{tr}f(\mathbf{A})\leqslant\mathrm{tr}f(\mathbf{B}).$ (5)
-Fact 4 (Trace Jensen).
+Fact 3 (Trace Jensen).
 For an interval $I\subseteq\mathbb{R}$, assume $f:I\to\mathbb{R}$ is convex. Then so is $\mathrm{tr}\circ f:\mathcal{S}_{d}^{I}\to\mathbb{R}$, i.e.,
@@
 if $\sum\lambda_{i}=1$ and each $\lambda_{i}\geqslant 0$.
-2.2 Concentration of Scalars
-We briefly review some classical and recent scalar tail bounds in this subsection.
-Markov inequality, the foundation of a large family of scalar concentration inequalities, states that if $X$ is a nonnegative random variable,
+2.2 Scalar Nonnegative Supermartingales and Statistical Applications
+A scalar-valued nonnegative supermartingale $\{Y_{n}\}_{n\geqslant 0}$ on the filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ satisfies Ville’s inequality [Ville, 1939],
+$\mathbb{P}(\exists n,\ Y_{n}\geqslant a)\leqslant(\mathbb{E}Y_{0})a^{-1},$ (VI)
+or equivalently,
+$\mathbb{P}(Y_{\tau}\geqslant a)\leqslant(\mathbb{E}Y_{0})a^{-1}\quad\text{for any stopping time }\tau.$ (VI)
+The equivalence is demonstrated by, for example, Howard et al. [2021, Lemma 3]. Ville’s inequality is often seen as a time-uniform, or a “stopped” generalization of Markov’s inequality
 $\mathbb{P}(X\geqslant a)\leqslant(\mathbb{E}X)a^{-1}$ (MI)
-for any $a>0$. There have been many generalizations of (MI). Ville’s inequality [Ville, 1939] states that a similar tail bound holds uniformly over time for a nonnegative supermartingale $\{M_{n}\}$,
-$\mathbb{P}(\exists n,\ M_{n}\geqslant a)\leqslant(\mathbb{E}M_{0})a^{-1},$ (VI)
-or equivalently, $\mathbb{P}(M_{\tau}\geqslant a)\leqslant(\mathbb{E}M_{0})a^{-1}$ for any stopping time $\tau$ [Howard et al., 2020].
-Ville’s inequality (VI) has been the workhorse for many recent time-uniform concentration inequalities (“confidence sequences”) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b] that have been favored by, for example, the multi-armed bandit literature.
-Applying (MI) to the square of the sum of i.i.d. random variables $X_{1},\dots,X_{n}$ of mean $\mu$ and variance $\sigma^{2}$, we have the famous Chebyshev inequality
-$\mathbb{P}\left(\left|\overline{X}_{n}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{na^{2}},$ (CI)
-where $\overline{X}_{n}$ is the arithmetic mean of $X_{1},\dots,X_{n}$.
-A perhaps less-known extension of (CI) is that if $X_{1},\dots,X_{n}$ have central $p$th moment $v_{p}$ ($1\leqslant p\leqslant 2$) instead of variance,
-$\mathbb{P}\left(\left|\overline{X}_{n}-\mu\right|\geqslant a\right)\leqslant\frac{2^{2-p}v_{p}}{n^{p-1}a^{p}},$ (5)
-which we prove in Section A.1 for completeness.
-In the other direction, if $X_{i}$’s are light-tailed and have moment generating function (MGF) $G(\gamma)=\mathbb{E}\mathrm{e}^{\gamma(X_{1}-\mu)}$, (MI) applied to the exponential of their mean yields the Chernoff-Hoeffding inequality
-$\mathbb{P}(\overline{X}_{n}-\mu\geqslant a)\leqslant G(\gamma/n)^{n}\mathrm{e}^{-\gamma a}.$ (6)
-Recently, Ramdas and Manole [2023] discovered some interesting strengthened forms of the bounds above.
-First, uniform randomization.
-It is found that if $U$ is a $\mathrm{Unif}_{(0,1)}$-distributed random variable independent from $X$ or $\{M_{n}\}$, (MI) and (VI) and be losslessly tightened into
-$\mathbb{P}(X\geqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (UMI)
-and
-$\mathbb{P}(M_{\tau}\geqslant Ua)\leqslant(\mathbb{E}M_{0})a^{-1}$ (UVI)
-respectively.
-The downstream bounds including (CI), (5) and (6) can thus also be randomized into strictly tighter forms.
-Second, Ramdas and Manole [2023] introduce exchangeable generalizations of (CI) which partly shares the anytime virtue of (VI)-type bounds, stating that as long as the infinite sequence $X_{1},X_{2},\dots$ is exchangeable,
-$\mathbb{P}\left(\exists n,\ \left|\overline{X}_{m}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{a^{2}};$ (XCI)
-and further, if i.i.d., (XCI) implies that for any fixed $N$,
-$\mathbb{P}\left(\exists n\geqslant N,\ \left|\overline{X}_{m}-\mu\right|\geqslant a\right)\leqslant\frac{\sigma^{2}}{Na^{2}}.$ (7)
-Extensions of (MI) and (VI) to random symmetric matrices are proved by Ahlswede and Winter [2002, Appendix C]; of (6) by Tropp [2012, 2015].
-We will bring up these previous extensions again soon in the following.
-Our current paper is partly animated by the question if these novel bounds (UMI), (UVI), (XCI), and (7) by Ramdas and Manole [2023] have matrix siblings as well; and if yes, to what extent of generality.
-We shall show that there is indeed a plethora of matrix tail bounds that recover as $d=1$ special cases the results mentioned above, using classical matrix analysis tools plus some new probability techniques.
-2.3 Templates of Our Bounds
-We now sketch out the skeletons that our bounds may have, and discuss their relation.
-Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d.
-random symmetric matrices whose common distribution belongs to a family $\mathcal{P}$.
-Denote by $\mathbf{M}$ their common mean and $\overline{\mathbf{X}}_{n}$ their empirical average $\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$.
-The following are the concentration events that we may prove to be of high probability.
-$\mathbf{A}\in\mathcal{S}^{+}_{d}$ denotes a constant matrix of deviation, and $r(\cdot)$ is a “randomization” map producing an $r(\mathbf{A})$ that is random but always smaller than $\mathbf{A}$ in Loewner order (the $Ua$ term in (UMI) is a one-dimensional example of $r(a)$).
-An arrow $A\to B$ denote the inclusion of events $A\subseteq B$, and hence also that $A$ is a stronger claim than $B$.
-{tikzcd}
-Most results in the past, for example (MHI), have been spectral bounds, which are weaker than many of the bounds we shall present in this paper.
-3 Randomized Fixed-Time Matrix Tail Bounds
-3.1 Randomized Matrix Markov Inequality
-We begin our exposition with a matrix version of uniformly randomized Markov inequality (UMI) by Ramdas and Manole [2023]; a randomized version of the matrix Markov inequality by Ahlswede and Winter [2002, Theorem 12],
+for a nonnegative random variable $X$.
+As Markov’s inequality lays the foundation for a large body of concentration inequalities for a fixed number of random variables leading up to non-asymptotic statistical methods (e.g.
+confidence intervals) valid at fixed finite sample sizes, Ville’s inequality (VI) has been the workhorse for many recent time-uniform concentration inequalities and accordingly sequential statistical methods which are valid at data-dependent sample sizes.
+For example, a $(1-\alpha)$-confidence sequence, which is derived almost exclusively via (VI) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b], is a sequence of confidence sets $\{C_{n}\}_{n\geqslant 1}$ for the parameter $\theta$ such that $\mathbb{P}(\theta\in C_{n}\text{ for all }n)\geqslant 1-\alpha$, or equivalently $\mathbb{P}(\theta\in C_{\tau})\geqslant 1-\alpha$ for any stopping time $\tau$ that can possibly depend on the data.
+More generally, the stopped version of (VI) is used in sequential hypothesis testing, as it implies that to construct a conservative sequential test of $\alpha$ level it suffices to construct a process $\{Y_{n}\}$ with $Y_{0}=1$ that is a nonnegative supermartingale under the null hypothesis, and reject it when $Y_{n}$ first surpasses $1/\alpha$.
+Various examples of sequential tests constructed via nonnegative supermartingales can be found in the recent review article by Ramdas et al.
+[2023, Sections 4 and 5].
+We return to this topic of test processes with a more involved discussion in Section 3.4.
+Numerous variations and extensions on (VI) exist.
+First, the backward Ville’s inequality (see e.g.
+Manole and Ramdas [2023, Theorem 2]).
+A process $\{Y_{n}\}$ adapted to a backward filtration $\{\mathcal{G}_{n}\}$ is called a backward submartingale if $\mathbb{E}(Y_{n-1}|\mathcal{G}_{n})\geqslant Y_{n}$ which, if nonnegative, satisfies (VI) as well,
+$\mathbb{P}(\exists n,\ Y_{n}\geqslant a)\leqslant(\mathbb{E}Y_{0})a^{-1}.$ (7)
+Second, the uniformly randomized Ville’s inequality [Ramdas and Manole, 2024, Theorem 4.1], which states that (VI) can be tightened into
+$\mathbb{P}(Y_{\tau}\geqslant Ua)\leqslant(\mathbb{E}Y_{0})a^{-1}\quad\text{forany stopping time }\tau,$ (8)
+where $U\sim\operatorname{Unif}_{(0,1)}$ is independent from the filtration $\{\mathcal{F}_{n}\}$.
+Ramdas and Manole [2024, Section 4.3] outline its implication for sequential statistics, which we discuss in Section 5.3 in the context of matrices.
+2.3 Matrix Concentration Inequalities
+2.3.1 Spectral Bounds
+There has been a wealth of results bounding the tails of the largest eigenvalue (or the spectral norm) of the sum of independent or martingale dependent random symmetric matrices.
+Most notably, Oliveira [2009, 2010], Tropp [2011, 2012, 2015], and their collaborators achieve the technically involved extension of the Cramér-Chernoff method, the central tool in deriving exponentially decaying scalar concentration inequalities, to random matrices satisfying similar tail conditions.
+The most prominent step along this route is the application of Lieb’s concavity theorem [Lieb, 1973] that ends up bounding the expected value of the scalar
+$\mathrm{tr}\exp\left(\sum\mathbf{X}_{i}\right)$ (9)
+where $\mathbf{X}_{i}$’s are independent [Tropp, 2012, Section 3.4] or martingale dependent [Tropp, 2011, Section 2.2] matrices.
+Fixed-time tail bounds follow from Markov’s inequality (MI).
+Already noticed by Tropp [2011] and further exploited by Howard et al.
+[2020], the trace process (9) is a (scalar) nonnegative supermartingale, and Ville’s inequality (VI) thus implies time-uniform spectral tail bounds for the same classes of random matrices.
+Our new results will take a different approach as we shall construct matrix-valued supermartingales and Loewner-ordered ($\preceq$) concentration inequalities.
+Their fixed-time counterparts are reviewed next in Section 2.3.2.
+Our results can often recover these “scalarized” results of Oliveira, Tropp, and Howard et al.
+mentioned above.
+2.3.2 Ahlswede-Winter Bounds
+Ahlswede and Winter [2002, Theorem 12] prove the following Loewner-ordered matrix Markov’s inequality
 $\mathbb{P}(\mathbf{X}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}),$ (MMI)
 where $\mathbf{X}$ is an $\mathcal{S}_{d}^{+}$-valued random matrix and $\mathbf{A}\in\mathcal{S}_{d}^{++}$ is deterministic.
-To achieve this, let us first define what it means for a random matrix to be uniformly and super-uniformly distributed.
-Note that it is easy to see that the randomizer $U$ in (UMI) and (UVI) can be replaced by any random variable that is stochastically larger; that is, one satisfying the following definition:
-Definition 3.1.
-We call a positive random scalar $U$ super-uniform if $\mathbb{P}(U\leqslant x)\leqslant x$ for all $x\geqslant 0$; that is, $U$ is stochastically greater than or equal to the uniform distribution $\operatorname{Unif}_{(0,1)}$.
-This has two viable extensions to matrices, one in terms of the trace and the other of the spectral norm.
-Definition 3.2.
-We call an $\mathcal{S}_{d}^{+}$-valued random matrix $\mathbf{U}$ spectrally super-uniform if $\mathbb{P}(\mathbf{U}\nsucceq\mathbf{Y})\leqslant\lambda_{\max}(\mathbf{Y})$ for all $\mathbf{Y}\in\mathcal{S}_{d}^{+}$, and $\mathcal{Y}$-spectrally uniform if, further, equality holds for all $\mathbf{Y}\in\mathcal{Y}\subseteq\mathcal{S}_{d}^{+}$.
-We say $\mathbf{U}$ is trace super-uniform if $\mathbb{P}(\mathbf{U}\nsucceq\mathbf{Y})\leqslant\mathrm{tr}\mathbf{Y}$ for all $\mathbf{Y}\in\mathcal{S}_{d}^{+}$, and $\mathcal{Y}$-trace uniform if, further, equality holds for all $\mathbf{Y}\in\mathcal{Y}\subseteq\mathcal{S}_{d}^{+}$.
-Note that every spectrally super-uniform matrix is a trace super-uniform matrix since $\lambda_{\max}(\mathbf{Y})\leqslant\mathrm{tr}\mathbf{Y}$.
-If $U$ is a super-uniform random scalar, then $U\mathbf{I}$ is a spectrally (hence trace) super-uniform random matrix, because $\mathbb{P}(U\mathbf{I}\nsucceq\mathbf{Y})=\mathbb{P}(U\leqslant\lambda_{\max}(\mathbf{Y}))\leqslant\lambda_{\max}(\mathbf{Y}).$
-Straightforwardly the identity matrix $\mathbf{I}$ is spectrally (hence trace) super-uniform.
-If $U\sim\operatorname{Unif}_{(0,1)}$, $U\mathbf{I}$ is $\mathcal{S}_{d}^{[0,1]}$-spectrally uniform, and $\left\{\mathbf{v}\mathbf{v}^{\mathsf{T}}:\|\mathbf{v}\|\leqslant 1\right\}$-trace uniform.
-Having defined trace super-uniformity, we can now state a matrix extension of the uniformly randomized Markov’s inequality (UMI).
-Theorem 3.3 (Uniformly Randomized Matrix Markov Inequality).
-Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{X}\subseteq\mathcal{S}_{d}^{+}$, and $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$.
-For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
-$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}).$ (UMMI)
-The above holds with equality if and only if $\mathbf{U}$ is $\mathbf{A}^{-1/2}\mathcal{X}\mathbf{A}^{-1/2}$-trace uniform.
-Proof.
-By our assumptions,
-$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})=\mathbb{P}(\mathbf{A}^{-1/2}\mathbf{X}\mathbf{A}^{-1/2}\npreceq\mathbf{U})=\mathbb{E}(\mathbb{P}(\mathbf{A}^{-1/2}\mathbf{X}\mathbf{A}^{-1/2}\npreceq\mathbf{U}|\mathbf{X}))$ (8)
-[…]
-concluding the proof. ∎
-[…]
-$\mathbb{P}(\lambda_{\max}(\mathbf{X})>\mathrm{tr}(\mathbf{A}\mathbf{U}))\leqslant\mathbb{P}(\mathrm{tr}\mathbf{X}>\mathrm{tr}(\mathbf{A}\mathbf{U}))\leqslant\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2}),$ (10)
-subject to the same upper bound $\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1})$.
-The following is an example where the inequality of (UMMI) holds with equality.
-We take $\mathbf{U}=U\mathbf{I}$ where $U\sim\operatorname{Unif}_{(0,1)}$, and $\mathbf{A}\in\mathcal{S}_{d}^{++}$.
-Let $\mathbf{x}$ be a random vector in the ellipsoid $\{\mathbf{x}:\mathbf{x}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{x}\leqslant 1\}\subseteq\mathbb{R}^{d}$.
-Then,
-$\mathbb{P}(\|\mathbf{A}^{-1/2}\mathbf{x}\|^{2}\leqslant U)=\mathbb{P}(\mathbf{x}\mathbf{x}^{\mathsf{T}}\npreceq U\mathbf{A})=\mathrm{tr}(\mathbf{A}^{-1}\mathbb{E}(\mathbf{x}\mathbf{x}^{\mathsf{T}}))=\mathbb{E}\|\mathbf{A}^{-1/2}\mathbf{x}\|^{2},$ (11)
-where the second equality above is (UMMI) applied on the random matrix $\mathbf{x}\mathbf{x}^{\mathsf{T}}$.
-It is not hard to see here the condition for equality is met.
-Intuitively, this hints that (UMMI) is tighter when the random matrix $\mathbf{X}$ is approximately of lower rank, a point that we shall return to later in this paper with statistical examples.
-Concluding this section, we remark on trace super-uniformity that clearly $U\mathbf{I}+\mathbf{Y}$ is trace super-uniform when $U\sim\operatorname{Unif}_{(0,1)}$ and $\mathbf{Y}\succeq 0$, and this provides an interesting class for our purposes.
-It may be interesting to fully characterize the set of all trace super-uniform matrices, but we leave it as future work.
-3.2 Randomized Matrix Chebyshev Inequalities
-Recall that the variance of a random symmetric matrix $\mathbf{X}$ is defined simply as $\mathbf{Var}(\mathbf{X})=\mathbb{E}(\mathbf{X}-\mathbb{E}\mathbf{X})^{2}$.
-Applying Theorem 3.3 to the square of one square-integrable matrix and the square of the sum of several i.i.d.
-square-integrable matrices, we have the following randomization of the one-matrix Chebyshev inequality by Ahlswede and Winter [2002, Theorem 14],
-$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2})$ (12)
-and the implied law of large numbers for the average $\overline{\mathbf{X}}_{n}$ of $n$ i.i.d.
-symmetric matrices with mean $\mathbf{M}$ and variance $\mathbf{V}$ by Ahlswede and Winter [2002, Theorem 15],
-$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (MCI)
-Lemma 3.4 (Uniformly Randomized One-Matrix Chebyshev Inequality).
-Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$, $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$.
-For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
-$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2}).$ (13)
-Proof.
-By Fact 2, $\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant\mathbb{P}((\mathbf{X}-\mathbb{E}\mathbf{X})^{2}\npreceq\mathbf{A}\mathbf{U}\mathbf{A})$.
-By Theorem 3.3, we have $\mathbb{P}((\mathbf{X}-\mathbb{E}\mathbf{X})^{2}\npreceq\mathbf{A}\mathbf{U}\mathbf{A})\leqslant\mathrm{tr}(\mathbf{Var}(\mathbf{X})\mathbf{A}^{-2})$.
-∎
-We can immediately apply Lemma 3.4 to multiple matrices, obtaining the following weak-law-of-large-numbers style inequality that shrinks when $n$ gets larger.
-Recall that the anticommutator between two sqaure matrices $\mathbf{A},\mathbf{B}$ is defined as $\{\mathbf{A},\mathbf{B}\}=\mathbf{A}\mathbf{B}+\mathbf{B}\mathbf{A}$.
-Theorem 3.5 (Uniformly Randomized $n$-Matrix Chebyshev Inequality).
-Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{i}=\mathbf{M}$ and variance matrix upper bound $\mathbf{Var}(\mathbf{X}_{i})\preceq\mathbf{V}$, such that $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$; $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from them all.
-Then, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$, we have
-$\mathbf{Var}(\overline{\mathbf{X}}_{n})\preceq n^{-1}\mathbf{V}.$ (14)
-Consequently, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$.
-$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq(\mathbf{A}\mathbf{U}\mathbf{A})^{1/2})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}),$ (UMCI)
-Proof.
-Note that
-$\small\mathbf{Var}(\overline{\mathbf{X}}_{n})=n^{-2}\mathbb{E}\left(\sum_{i=1}^{n}(\mathbf{X}_{i}-\mathbf{M})\right)^{2}=n^{-2}\left(\sum_{i=1}^{n}\mathbf{Var}(\mathbf{X}_{i})+\sum_{i<j}\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}\right)\preceq n^{-1}\mathbf{V},$ (15)
-and apply Lemma 3.4 to conclude the proof. ∎
-It is easy to see that (UMCI) recovers as special cases (MCI) and (CI).
-Remark 3.6.
-The assumption $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for $i\neq j$ is satisfied when $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ are pairwise uncorrelated, which is in turn satisfied under martingale difference assumption (i.e., $\{\mathbf{X}_{i}\}$ is adapted to some filtration $\{\mathcal{F}_{i}\}$ with $\mathbb{E}(\mathbf{X}_{i}|\mathcal{F}_{i-1})=\mathbf{M}$ and $\mathbf{Var}(\mathbf{X}_{i}|\mathcal{F}_{i-1})\preceq\mathbf{V}$), which is in turn implied if they are i.i.d.
-For $p\in[1,2)$, a $p$th moment bound can lead to an infinite variance version of (13), whose non-randomized ($\mathbf{U}=\mathbf{I}$) version is mentioned by Tu [2017].
-Lemma 3.7 (Uniformly Randomized One-Matrix $p$-Chebyshev Inequality).
-Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$ that has a finite $p$th central moment $\mathbf{V}_{p}:=\mathbb{E}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}$ and $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$.
-For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
-$\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (16)
-Proof.
-By Fact 2, $\mathbb{P}(\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X})\npreceq(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p})\leqslant\mathbb{P}((\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}\npreceq\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})$.
-By Theorem 3.3, we have $\mathbb{P}((\operatorname{abs}(\mathbf{X}-\mathbb{E}\mathbf{X}))^{p}\npreceq\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p})$.
-∎
-However, an $L^{p}$ analogy of (UMCI) (i.e.
-a matrix version of the $p$-Chebyshev inequality (5)) is not straightforward.
-This is largely because while the contraction of variance rule (14) exists for matrices, we do not know of any matrix version of $p$th moment contraction rule.
-In fact, the scalar $p$th moment contraction rule was only recently established (by Cherapanamjeri et al.
-[2022, Lemma 4.2] and Wang et al.
-[2021, Lemma 7], see the proof of (5) in Section A.1).
-Section A.2 contains a conjectured result along these lines, as well as a looser result that is the best we can currently achieve.
-3.3 Randomized Matrix Chernoff Bounds
-We now establish the classical Chernoff-type bounds with randomization for light-tailed random matrices that have finite moment generating functions (MGFs).
-Lemma 3.8 (Uniformly Randomized One-Matrix Chernoff Bound).
-Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{S}_{d}$, $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$.
-For any $\mathbf{A}\in\mathcal{S}_{d}$ and real number $\gamma$,
-$\mathbb{P}\left(\mathbf{X}\npreceq\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right)\leqslant\mathrm{tr}(\mathrm{e}^{-2\gamma\mathbf{A}}\,\mathbb{E}\mathrm{e}^{2\gamma\mathbf{X}}).$ (17)
-Proof.
-By Fact 2, we have $\mathbb{P}\left(2\gamma\mathbf{X}\npreceq\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right)\leqslant\mathbb{P}(\mathrm{e}^{2\gamma\mathbf{X}}\npreceq\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})$.
-By Theorem 3.3, we have $\mathbb{P}(\mathrm{e}^{2\gamma\mathbf{X}}\npreceq\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\leqslant\mathrm{tr}(\mathrm{e}^{-2\gamma\mathbf{A}}\,\mathbb{E}\mathrm{e}^{2\gamma\mathbf{X}})$, concluding the proof.
-∎
-The seemingly indigestible tail threshold $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})$ in (17) can be largely simplified in any of the following scenarios.
-First, one can simply take the trace on the tail event above, obtaining due to Fact 1 and the monotonicity of the trace,
-$\mathbb{P}\left(\mathrm{tr}\mathbf{X}>\mathrm{tr}\mathbf{A}+\frac{1}{2\gamma}\mathrm{tr}\log\mathbf{U}\right)\leqslant\mathbb{P}\left(\mathbf{X}\npreceq\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})\right),$ (18)
-without any additional assumptions; Second, if $\mathbf{U}=U\mathbf{I}$ where $U$ is scalar super-uniform, $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})=\mathbf{A}+\frac{1}{2\gamma}\mathbf{I}\log U$; Third, if $\mathbf{A}=a\mathbf{I}$, $\frac{1}{2\gamma}\log(\mathrm{e}^{\gamma\mathbf{A}}\mathbf{U}\mathrm{e}^{\gamma\mathbf{A}})=a\mathbf{I}+\frac{1}{2\gamma}\log\mathbf{U}$.
-It is known that, unlike the scalar case, the one-matrix MGF is not linearly additive. Once again, the complication is partially resolved when the trace is applied.
-Fact 5 (Lemma 3.4 of Tropp [2012]).
-Let $\mathbf{Z}_{1},\dots,\mathbf{Z}_{n}$ be independent random matrices taking values in $\mathcal{S}_{d}$. Then,
-$\mathrm{tr}\left\{\mathbb{E}\exp\left(\sum\mathbf{Z}_{i}\right)\right\}\leqslant\mathrm{tr}\exp\left(\sum\log\mathbb{E}\mathrm{e}^{\mathbf{Z}_{i}}\right).$ (19)
-This implies the following bound that recovers as special cases both the scalar Chernoff-Hoeffding inequality (6) when taking $d=\mathbf{U}=1$, and the “master tail bound” on $\lambda_{\max}(\overline{\mathbf{X}}_{n}-\mathbf{M})$ by Tropp [2012, Theorem 3.6] when taking $\mathbf{U}=\mathbf{I}$.
-Theorem 3.9 (Uniformly Randomized Matrix Chernoff-Hoeffding Inequality).
-Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d.
-random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and matrix-valued MGF $\mathbf{G}(\gamma)=\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}$; $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from them all.
-Let $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$.
-For any $a,t>0$,
-$\mathbb{P}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\npreceq a\mathbf{I}+\frac{\log\mathbf{U}}{\gamma}\right)\leqslant\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})\mathrm{e}^{-\gamma a}.$ (20)
-Proof.
-By Fact 5, $\mathrm{tr}(\mathbb{E}\mathrm{e}^{\gamma(\overline{\mathbf{X}}_{n}-\mathbf{M})})\leqslant\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})$.
-Using Lemma 3.8 on $\gamma(\overline{\mathbf{X}}_{n}-\mathbf{M})$ with $\mathbf{A}=a\mathbf{I}$ concludes the proof.
-∎
-Again, the bound remains valid if only an upper bound of the matrix MGF $\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}\preceq\mathbf{G}(\gamma)$ or (more commonly) the weaker log-MGF bound $\log\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}\preceq\log\mathbf{G}(\gamma)$ is known.
-Table 1 contains some important known matrix (log-)MGFs (see Tropp [2012, 2015]).
-An asymmetric Hoeffding bound is missing here but shall be discussed in Section 5 later since it requires some further techniques beyond the matrix log-MGF.
-Name Condition $\mathbf{G}(\gamma)$ $\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})$
-Rademacher $\mathbf{X}_{1}-\mathbf{M}=R_{1}\mathbf{C},R_{1}\sim\operatorname{rad}_{1/2}$ $\preceq\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$ $\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$
-Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $=\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$ $=\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$
-Bennett I $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})=\mathbf{V}$ $\preceq\mathrm{e}^{(\mathrm{e}^{\gamma}-\gamma-1)\mathbf{V}}$ $\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$
-Bennett II $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})\preceq\mathbf{V}$ $\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$
-Symmetric Hoeffding $(\mathbf{X}_{1}-\mathbf{M})^{2}\preceq\mathbf{B},\ \mathbf{X}_{1}-\mathbf{M}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}-\mathbf{X}_{1}$ $\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{B}/2n})$
-Table 1: Some important matrix MGFs and log-MGFs.
-Here $\mathbf{B},\mathbf{C},\mathbf{V}$ are deterministic matrices, and =$\stackrel{{\scriptstyle d}}{{=}}$ means two random matrices having the same distribution.
-The proofs are in Section B.1.
-4 Maximal Inequalities of Matrix (Super-/Sub-)Martingales
-Some interesting tail bounds are obtained if we consider matrix-valued supermartingales and submartingales and their maximal inequalities.
-These processes can be defined on forward or backward filtrations and in both these cases, well-known scalar-valued maximal inequalities can be extended to matrices.
-4.1 Matrix-Valued Supermartingales
-Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued stochastic process adapted to the filtration $\{\mathcal{F}_{n}\}$.
-[^3] We shall often omit the filtration $\{\mathcal{F}_{n}\}$ when unambiguous from the context.
-We say that $\{\mathbf{Y}_{n}\}$ is a supermartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}$ holds for every $n$.
+This leads to the direct corollary of a matrix Chebyshev inequality that serves as a weak law of large numbers for the average $\overline{\mathbf{X}}_{n}$ of $n$ i.i.d.
+symmetric matrices with mean $\mathbf{M}$ and (matrix) variance $\mathbf{V}=\mathbb{E}(\mathbf{X}-\mathbf{M})^{2}$ [Ahlswede and Winter, 2002, Corollary 16],
+$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (10)
+Ahlswede and Winter [2002] also attempt to use this “un-scalarized” method to generalize the Cramér-Chernoff method for scalars.
+However, as summarized by [Tropp, 2012, Section 3.7], this leads to suboptimal bounds that involve “the sum of eigenvalues” as opposed to “the eigenvalue of a sum” which the sharper bounds (obtained by “scalarizing” with Lieb’s concavity theorem as mentioned in Section 2.3.1) involve.
+Our time-uniform, martingale-based main results are often obtained via (MMI).
+In particular, we show that this Loewner-ordered approach pioneered by Ahlswede and Winter [2002] does lead to sharp matrix concentration inequalities and tests, fixed-time and time-uniform alike, comparable to and sometimes better than the results cited in Section 2.3.1.
+We provide a self-contained proof of (MMI) in Appendix A that also comes with the extension of randomization.
+3 Forward PSD Supermartingales
+We consider in this section a probability space with a usual forward filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
+That is, $\mathcal{F}_{0}\subseteq\mathcal{F}_{1}\subseteq\dots\subseteq\mathcal{F}_{\infty}$.
+Within the scope of this section, we shall often omit the filtration $\{\mathcal{F}_{n}\}$ when unambiguous from the context, calling a scalar- or matrix-valued process “adapted” if the $n$-indexed element is $\mathcal{F}_{n}$-measurable; “predictable” if $\mathcal{F}_{n-1}$-measurable.
+Stopping times are all with respect to $\{\mathcal{F}_{n}\}$.
+3.1 Convergence, Optional Stopping, and Maximal Inequalities
+Recall our definition of a matrix-valued supermartingale from the beginning of the paper that an adapted $\mathcal{S}_{d}$-valued stochastic process $\{\mathbf{Y}_{n}\}$ is a supermartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}$ holds for every $n$.
 Before we proceed, note that the order $\preceq$ is preserved by addition, limit, and Lebesgue integration (hence including convergent infinite sums, expectations, conditional expectations, etc., essentially because $\mathcal{S}^{+}=\{\mathbf{A}:\mathbf{A}\succeq 0\}$ is closed both topologically and under $+$).
@@
 Lemma 3.1.
-An $\mathcal{S}_{d}$-valued process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every deterministic vector $\mathbf{v}\in\mathbb{R}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale.
-Proof.
-$\{\mathbf{Y}_{n}\}$ is a supermartingale (21)
-$\iff$ For all $n$, $\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq 0$ (22)
-$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}(\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1}))\mathbf{v}\geqslant 0$ (23)
-$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n-1}\mathbf{v}-\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}|\mathcal{F}_{n-1})\geqslant 0$ (24)
-$\iff$ For all $\mathbf{v}$, $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale. (25)
-This completes the proof. ∎
-This leads to an analog of Doob’s martingale convergence theorem for positive semidefinite matrix supermartingales.
+An $\mathcal{S}_{d}$-valued adapted process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every non-random vector $\mathbf{v}\in\mathbb{R}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale.
+The very straightforward proof of Lemma 3.1 can be found in Section B.2. It leads to an analog of Doob’s martingale convergence theorem for PSD matrix supermartingales.
 Theorem 3.2 (Matrix Supermartingale Convergence Theorem).
@@
 First, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that $\mathbf{Y}_{n}$’s diagonal converges to integrable scalars.
-Then, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that the rest of $\mathbf{Y}_{n}$’s entries also converge to integrable scalars.
+Then, setting $\mathbf{v}$ to $(0,\dots,0,1,0,\dots,0,1,0,\dots,0)^{\mathsf{T}}$ we see that $\mathbf{Y}_{n}$’s off-diagonal entries also converge to integrable scalars.
 The limit $\mathbf{Y}_{\infty}=\lim_{n\to\infty}\mathbf{Y}_{n}$ is in $\mathcal{S}_{d}^{+}$ because $\mathcal{S}_{d}^{+}$ is a closed set.
@@
 so $\mathbb{E}\mathbf{Y}_{\infty}\preceq\mathbb{E}\mathbf{Y}_{0}$. ∎
-Next, we see that using a similar trick, any scalar version of the optional stopping (sampling) theorems extends to symmetric matrices as well. For example:
-Lemma 4.3 (Matrix Optional Stopping).
-Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$. Then,
+The convergence of PSD supermartingales ensures that we can speak of the stopped matrix $\mathbf{Y}_{\tau}$ for any stopping time $\tau$, even when $\mathbb{P}(\tau=\infty)>0$.
+We can use a similar method to extend some scalar versions of the optional stopping (sampling) theorems to matrices.
+For example:
+Theorem 3.3 (Matrix Optional Stopping).
+Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\tau$ be a stopping time. Then,
   1. 1.
-    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale on $\{\mathcal{F}_{n}\}$.
+    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale.
   2. 2.
@@
 3.
-Apply Theorem 4.2 to the $\mathcal{S}_{d}^{+}$-valued stopped supermartingale $\{\mathbf{Y}_{\tau\wedge n}\}$, we have $\mathbb{E}\mathbf{Y}_{\tau\wedge\infty}\preceq\mathbb{E}\mathbf{Y}_{\tau\wedge 0}$, i.e., $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
+Applying Theorem 3.2 to the $\mathcal{S}_{d}^{+}$-valued stopped supermartingale $\{\mathbf{Y}_{\tau\wedge n}\}$, we have $\mathbb{E}\mathbf{Y}_{\tau\wedge\infty}\preceq\mathbb{E}\mathbf{Y}_{\tau\wedge 0}$, i.e., $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
 ∎
-This gives rise to the matrix versions of (VI) and (UVI) via some manipulation of crossing events.
-Theorem 4.4 ((Uniformly Randomized) Matrix Ville’s Inequality).
-Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale on $\{\mathcal{F}_{n}\}$, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$.
-Let $\mathbf{U}$ be a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathcal{F}_{\infty}$.
-Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
-$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (UMVI)
-Consequently,
-$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
+This gives rise to the matrix versions of the two mutually equivalent statements of Ville’s inequality (VI) via some manipulation of crossing events.
+Theorem 3.4 (Matrix Ville’s Inequality).
+Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale. Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
+$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1});$ (MVI)
+or equivalently,
+for any stopping time $\text{for any stopping time }\tau,\quad\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
 Proof.
-First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Lemma 4.3. By Theorem 3.3 and the monotinicity of the trace,
-$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1/2}),$ (27)
-which is (UMVI). Next, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$. Setting $\mathbf{U}$ to $\mathbf{I}$ and $\tau$ to $\nu\wedge N$ in (UMVI), we have
+We prove the stopped inequality first. Due to Part 3 of Theorem 3.3, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. Now by (MMI) and the monotinicity of the trace,
+$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1/2}),$ (12)
+which is the stopped (MVI). Next, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$. Applying the result above with $\tau=\nu\wedge N$,
 $\mathbb{P}(\nu\leqslant N)=\mathbb{P}(\mathbf{Y}_{\nu\wedge N}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1})$ (13)
@@
 $\mathbb{P}(\nu<\infty)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1})$ (14)
-which concludes the proof of (MVI). ∎
-We remark that while there are two equivalent ways of stating the Ville’s inequality, the “any-time” statement (VI) and the stopping time statement (stated after (VI)), it appears that randomization only applies to the latter [Ramdas and Manole, 2023].
-The same applies to our matrix versions of Ville’s inequality.
-Finally, we remark that some other maximal inequalities for scalar sub- and super-martingales can be extended to matrices as well, of which we derive a few in Appendix C.
-4.2 Matrix Generalizations of Classical Test Supermartingales
-Recall that scalar-valued nonnegative supermartingales are often constructed incrementally by multiplying each time a nonnegative variable with conditional expectation $\leqslant 1$.
-Analogously, we may use the following straightforward fact to construct $\mathcal{S}_{d}^{+}$-valued supermartingales.
-Lemma 4.5.
-Let $\mathbf{E}_{n}$ be $\mathcal{S}_{d}^{+}$-valued and $\mathcal{F}_{n}$-measurable, $\mathbf{F}_{n}$ be $\mathcal{S}_{d}^{++}$-valued and $\mathcal{F}_{n-1}$-measurable, such that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ almost surely.
-Then, the following $\mathcal{S}_{d}^{+}$-valued process is a supermartingale on $\{\mathcal{F}_{n}\}$:
-$\mathbf{Y}_{n}=\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}.$ (30)
-In some cases, we can take $\mathbf{F}_{1}=\mathbf{F}_{2}=\dots=\mathbf{I}$, and thus $\mathbf{Y}_{n}=\sqrt{\mathbf{E}_{1}}\dots\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\dots\sqrt{\mathbf{E}_{1}}$.
+which concludes the proof of both forms of (MVI). ∎
+A natural question arises: in comparison to the scalar (VI), does (MVI) provide tighter, looser, or incomparable concentration results?
+This question is essential as from every PSD supermartingale $\{\mathbf{Y}_{n}\}$ one can construct scalar nonnegative supermartingales $\{\mathrm{tr}\mathbf{Y}_{n}\}$ and $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$.
+We shall answer the question in detail, theoretically in Sections 5.1 and 5.2 and empirically in Section 6.
+As our exposition unfolds, the following three conclusions shall become clear:
+  1. 1.
+    (MVI) is in general not reducible to the scalar (VI).
+  2. 2.
+    To conceptualize the difference between (MVI) and (VI), it suffices to look at the one-time case, comparing (MMI) and (MI).
+The spectra of both the “threshold” matrix $\mathbf{A}$ and the random matrix $\mathbf{X}$ matter.
+  3. 3.
+    While the one-time Markov case reveals the potential tightness benefit of (MVI), in real data the conditions for such benefit often arise naturally in the sequential, anytime-valid testing case that necessitates Ville-type inequalities.
+As an aside, we present the additional result of a matrix analog of Doob’s $L^{p}$ inequality for forward matrix martingales in Section C.2.
+It is not necessarily relevant to our upcoming discussion in statistical contexts but might be of independent interest.
+3.2 Matrix Test Supermartingales
+Scalar-valued nonnegative supermartingales are often constructed incrementally: if one can find an adapted nonnegative process $\{E_{n}\}$ that has the conditional expectation bound $\mathbb{E}(E_{n}|\mathcal{F}_{n})\leqslant 1$, then their product $Y_{n}=\prod_{i=1}^{n}E_{i}$ is a nonnegative supermartingale.
+In particular, the supermartingale satisfies $Y_{0}=1$ and thus satisfies (VI) with $\mathbb{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$.
+It is called a test supermartingale.
+We propose the following analogous way to construct $\mathcal{S}_{d}^{+}$-valued supermartingales that involves a bit more flexibility.
+Lemma 3.5 (Construction of Matrix Test Supermartingales).
+If an adapted $\mathcal{S}_{d}^{+}$-valued process $\{\mathbf{E}_{n}\}_{n\geqslant 1}$ and a predictable $\mathcal{S}_{d}^{++}$-valued process $\{\mathbf{F}_{n}\}_{n\geqslant 1}$ satisfy $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ almost surely for all $n\geqslant 1$.
+Then, the following is a PSD supermartingale:
+$\mathbf{Y}_{n}=\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}},$ (15)
+with $\mathbf{Y}_{0}=\mathbf{I}$. It is called a matrix test supermartingale.
+In some cases, we can take $\mathbf{F}_{1}=\mathbf{F}_{2}=\dots=\mathbf{I}$, where the process has a simpler expression $\mathbf{Y}_{n}=\sqrt{\mathbf{E}_{1}}\dots\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\dots\sqrt{\mathbf{E}_{1}}$.
 Proof.
@@
 concluding the proof. ∎
-We remark that by defining $\mathbf{E}_{n}^{\prime}=\sqrt{\mathbf{F}_{n}}\mathbf{E}_{n}\sqrt{\mathbf{F}_{n}}$ and $\mathbf{F}_{n}^{\prime}=\mathbf{I}$, we can obtain another $\mathcal{S}_{d}^{+}$-valued supermartingale $\sqrt{\mathbf{E}_{1}^{\prime}}\dots\sqrt{\mathbf{E}_{n}^{\prime}}\sqrt{\mathbf{E}_{n}^{\prime}}\dots\sqrt{\mathbf{E}_{1}^{\prime}}$.
-This option, however, involves more computation than (30) which can be computed incrementally by computing the two matrices $\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}$ and $\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}}$ on the fly.
-With Lemma 4.5, we can recover matrix versions of several recent “test supermartingales” (nonnegative supermartingales with initial value one) in the literature.
-In the following examples, we will state some conditions on a sequence of $\mathcal{S}_{d}$-valued random matrices $\{\mathbf{X}_{n}\}$ that will lead to a sequence of $\{\mathbf{E}_{n}\}$ and $\{\mathbf{F}_{n}\}$ that satisfy Lemma 4.5, so that a $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ can be constructed accordingly via (30).
-First, we have the MGF-type supermartingales whose well-known scalar case is mentioned by e.g. Howard et al. [2020, Fact 1(a)].
+The expression of $\mathbf{Y}_{n}$ involves splitting the stepwise matrices into square roots and multiplying bidirectionally from the outside, which, we note, ensures the product is still in $\mathcal{S}_{d}$ and has the conditional expectation property we want; in particular, $\mathbf{F}_{n}\mathbf{E}_{n}$ and $\mathbf{E}_{n}\mathbf{F}_{n}$ may not be symmetric.
+The process $\{\mathbf{Y}_{n}\}$ can be computed incrementally on the fly by noting that
+$\mathbf{Y}_{n}=\mathbf{L}_{n}\mathbf{R}_{n},\quad\mathbf{L}_{n}=\mathbf{L}_{n-1}\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}},\quad\mathbf{R}_{n}=\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\mathbf{R}_{n-1}.$ (18)
+We remark that by defining $\mathbf{E}_{n}^{\prime}=\sqrt{\mathbf{F}_{n}}\mathbf{E}_{n}\sqrt{\mathbf{F}_{n}}$ and $\mathbf{F}_{n}^{\prime}=\mathbf{I}$, we can obtain another PSD supermartingale $\sqrt{\mathbf{E}_{1}^{\prime}}\dots\sqrt{\mathbf{E}_{n}^{\prime}}\sqrt{\mathbf{E}_{n}^{\prime}}\dots\sqrt{\mathbf{E}_{1}^{\prime}}$.
+This option, however, involves slightly more matrix product computation than (15).
+Algorithm 1 Sequential Test via a Matrix Test Supermartingale
+Null hypothesis $\mathcal{H}_{0}$ on the data stream $X_{1},X_{2},\dots$ taking values in $\mathcal{X}$;
+Functions $\mathbf{E}:\mathcal{X}\to\mathcal{S}_{d}^{+}$ and $\mathbf{F}:\mathcal{X}\to\mathcal{S}_{d}^{++}$ such that $\mathbb{E}(\mathbf{E}(X_{n})|X_{1},\dots,X_{n-1})\preceq\mathbf{F}^{-1}(X_{n})$ whenever $\mathcal{H}_{0}$ is true;
+Type-I error level $\alpha\in(0,1)$; $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$
+$\mathcal{H}_{0}$ is falsely rejected with probability at most $\alpha$
+$\mathbf{L}\leftarrow\mathbf{I}$, $\mathbf{R}\leftarrow\mathbf{I}$
+for $n=1,2,\dots$ do
+     observe $X_{n}$
+     $\mathbf{L}\leftarrow\mathbf{L}\sqrt{\mathbf{F}(X_{n})}\sqrt{\mathbf{E}(X_{n})}$
+     $\mathbf{R}\leftarrow\sqrt{\mathbf{E}(X_{n})}\sqrt{\mathbf{F}(X_{n})}\mathbf{R}$
+     if  $\mathbf{L}\mathbf{R}\npreceq\mathbf{A}$  then
+         reject $\mathcal{H}_{0}$
+     end if
+end for
+With Lemma 3.5, we can recover matrix versions of several scalar test supermartingales in the recent sequential statistics literature.
+In the next subsection, we will state some conditions on a sequence of $\mathcal{S}_{d}$-valued random matrices $\{\mathbf{X}_{n}\}$ that will lead to two sequences of $\{\mathbf{E}_{n}\}$ and $\{\mathbf{F}_{n}\}$ that satisfy Lemma 3.5, so that a $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{Y}_{n}\}$ can be constructed accordingly via (15).
+Combining with (MVI), we form the following generic scheme for sequentially testing a null hypothesis $\mathcal{H}_{0}$ using a PSD supermartingale outlined in Algorithm 1, where if $\mathcal{H}_{0}$ is true the procedure has at most $\alpha$ probability ever rejecting it.
+We further remark that, while with scalar nonnegative supermartingales formed multiplicatively, one often works on the log scale to ease the computation.
+That is, if $M_{n}=\prod_{i=1}^{n}E_{i}$ is a scalar nonnegative supermartingale, one often adds up $\log E_{1}+\dots+\log E_{n}$ and rejects the null when this sum exceeds $\log(1/\alpha)$.
+With the PSD supermartingale formed in Lemma 3.5, while it is indeed true that
+[…]
+due to the monotonicity of the matrix logarithm (Fact 1), one can not compute $\log\mathbf{Y}_{n}$ via summing $\log\mathbf{E}_{i}$ and $\log\mathbf{F}_{i}$.
+An additive alternative to (MVI) that avoids matrix multiplication might be of interest for future work.
+3.3 Examples: Conditional Mean Testing Problems
+We have numerous examples for Lemma 3.5 where one can conduct sequential hypothesis testing with Algorithm 1.
+All examples below, we note, involve the null hypothesis in the form of “the conditional mean of $\mathbf{X}_{n}$ equals $\mathbf{M}_{n}$” where $\{\mathbf{M}_{n}\}$ is a pre-specified predictable sequence.
+A common special case is when the matrices $\{\mathbf{X}_{n}\}$ are i.i.d.
+and one tests if their common mean equals a specific value.
+First, we have the moment generating function (MGF)-based supermartingales whose well-known scalar case is mentioned by e.g. Howard et al. [2020, Fact 1(a)].
 Example 3.6 (Matrix MGF Supermartingale).
-Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with common conditional mean matrix $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}$ and matrix-valued conditional MGF upper bound $\mathbf{G}(\gamma)\succeq\mathbb{E}(\mathrm{e}^{\gamma(\mathbf{X}_{n}-\mathbf{M})}|\mathcal{F}_{n-1})$.
-Then, for each $n$, we can take $\mathbf{E}_{n}=\mathrm{e}^{\gamma_{n}(\mathbf{X}_{n}-\mathbf{M})}$, $\mathbf{F}_{n}=\mathbf{G}(\gamma_{n})^{-1}$ where $\{\gamma_{n}\}$ are predictable scalars, in Lemma 4.5.
-The MGF upper bounds to use may be found in Table 1. Second, the “betting” style martingales by Waudby-Smith and Ramdas [2023].
-Example 4.7 (Betting).
-Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}^{+}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and upper bounds $\mathbf{B}_{n}$ that are predictable (i.e., $\mathbf{X}_{n}\preceq\mathbf{B}_{n}$ a.s., and $\mathbf{B}_{n}$ is $\mathcal{F}_{n-1}$-measurable).
+Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and matrix-valued conditional MGF upper bound $\mathbf{G}_{n}(\gamma)\succeq\mathbb{E}(\mathrm{e}^{\gamma(\mathbf{X}_{n}-\mathbf{M}_{n})}|\mathcal{F}_{n-1})$.
+Then, for each $n$, we can take $\mathbf{E}_{n}=\mathrm{e}^{\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})}$, $\mathbf{F}_{n}=\mathbf{G}_{n}(\gamma_{n})^{-1}$ where $\{\gamma_{n}\}$ are predictable scalars, in Lemma 3.5.
+It is worth noting that in general it is hard to prove matrix MGF bounds and they are known only in very few cases.
+On the other hand, much more is known on bounding the matrix log-MGF $\log\left(\mathbf{G}_{n}(\cdot)\right)$ under various distribution assumptions, obtained via the Lieb-Tropp technique mentioned in Section 2.3.1.
+These log-MGF upper bounds only lead to scalar-valued supermartingales (9) instead of matrix-valued supermartingales.
+We review some existing results of matrix MGFs by Tropp [2012, 2015] in Section C.1.
+Our second example extends the wealth process of betting against the fairness of a coin.
+The scalar case was initially proposed for Bernoulli coin tosses, generalized to any bounded random scalars by Shafer and Vovk [2005, Chapters 3.1 and 3.2 resp.], and recently applied to sequential mean testing by Waudby-Smith and Ramdas [2023].
+We have the following corresponding result for bounded random matrices.
+Example 3.7 (Matrix Betting).
+Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}^{+}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and upper bounds $\mathbf{B}_{n}$ that are predictable (i.e., $\mathbf{X}_{n}\preceq\mathbf{B}_{n}$ a.s., and $\mathbf{B}_{n}$ is $\mathcal{F}_{n-1}$-measurable).
 We can take, in Lemma 3.5,
@@
 Besides these two examples where we extend prior scalar results to matrices, we also note that Howard et al.
-[2020, Section 6.5] prove a total of eight inequalities for matrix martingale differences $\{\mathbf{Z}_{n}\}$ in the style of
+[2020, Section 6.5] prove a total of eight inequalities for $\mathcal{S}_{d}$-valued martingale differences $\{\mathbf{Z}_{n}\}$ in the style of
 $\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\exp(\mathbf{C}_{n}^{\prime}),$ (22)
@@
 However, Howard et al.
-[2020] do not consider combining these conditional expectation bounds into matrix-valued supermartingales via (30).
-For reasons we will discuss later in Section 5; they pursue scalar supermartingales instead.
-Of their eight inequalities, we will introduce two interesting cases now, omitting the other six due to their similar construction.
-First, the self-normalized supermartingale that only assumes the second moment exists [Howard et al., 2020, Part (f) of Section 6.5], which we (re)prove in Section B.3.
+[2020] do not consider combining these conditional expectation bounds into matrix-valued supermartingales via (15); they pursue scalar supermartingales instead, with methods we discuss later in Section 5.2.
+Of their eight inequalities, we will introduce two interesting cases now and (re-)prove them in Section B.2, omitting the other six due to their similar construction.
+First, the self-normalized supermartingale that only assumes the second moment exists [Howard et al., 2020, Part (f) of Section 6.5].
 Example 3.8 (Self-Normalized Matrix Supermartingale).
-Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional variance upper bounds $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ that are predictable.
+Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional variance upper bounds $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ that are predictable.
 We can take, in Lemma 3.5,
@@
 where $\{\gamma_{n}\}$ are predictable positive scalars.
-Second, a bound for symmetric random matrices that have a symmetric (around $0\in\mathcal{S}_{d}$), potentially heavy-tailed distribution [Howard et al., 2020, Part (d) of Section 6.5] which we prove as well in Section B.3.
+Second, a bound for symmetric random matrices that have a symmetric (around $0\in\mathcal{S}_{d}$), potentially heavy-tailed distribution [Howard et al., 2020, Part (d) of Section 6.5].
 Example 3.9.
-Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and satisfy $\mathbf{X}_{n}-\mathbf{M}_{n}|\mathcal{F}_{n-1}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}_{n}-\mathbf{X}_{n}|\mathcal{F}_{n-1}$.
+Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and satisfy $\mathbf{X}_{n}-\mathbf{M}_{n}|\mathcal{F}_{n-1}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}_{n}-\mathbf{X}_{n}|\mathcal{F}_{n-1}$.
 Then, we can take, in Lemma 3.5,
@@
 where $\{\gamma_{n}\}$ are predictable positive scalars.
-Finally, let us present a novel matrix extension of the empirical Bernstein bound, in the style of Howard et al. [2021, Theorem 4], following whom we define the function
-$\psi_{E}(x)=-\log(1-x)-x=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$ (38)
+Further, let us present two examples in the light- and heavy-tail regimes respectively that do not appear among the ones due to Howard et al.
+[2020, Section 6.5].
+First, a matrix extension of empirical Bernstein inequalities, in the style of Howard et al.
+[2021, Theorem 4], following whom we define the function
+$\psi_{\mathrm{E}}(x)=-\log(1-x)-x$ (25)
 for $x\in[0,1)$.
-The notation comes from the fact that $\psi_{E}$ is the log-MGF of a centered standard exponential random variable.
-We distinguish the following result for its novelty by labeling it a “proposition” as opposed to an “example”.
-Proposition 4.10 (Empirical Bernstein Matrix Supermartingale).
+The notation comes from the fact that $\psi_{\mathrm{E}}$ is the log-MGF of a centered standard exponential random variable.
+The following is a rephrasing of a recent result by Wang and Ramdas [2024, Lemma 4.1] in the language of our Lemma 3.5.
+Example 3.10 (Empirical Bernstein-Type Matrix Supermartingale).
 Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$.
@@
 Then, we can take, in Lemma 3.5,
-$\mathbf{E}_{n}=\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}),\quad\mathbf{F}_{n}=\exp(-\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (39)
+$\mathbf{E}_{n}=\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{\mathrm{E}}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}),\quad\mathbf{F}_{n}=\exp(-\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (26)
 where $\{\gamma_{n}\}$ are predictable $(0,1)$-valued scalars.
-We delay the proof that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ to Section B.4.
-The reason why it is called “empirical” Bernstein is the “plug-in empirical variance” component term $(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}$, as $\widehat{\mathbf{X}}_{n}$ can usually be taken as an empirical average (possibly weighted) from $\mathbf{X}_{1}$ to $\mathbf{X}_{n-1}$, in which case the condition $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ is satisfied when $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$.
-We shall come to this empirical Bernstein result again with some scalarization tricks in Section 5.3, where we obtain a strikingly sharp spectral concentration inequality in closed form, further justify its name, discuss the context of empirical Bernstein bounds, and compare it with other existing results.
-These supermartingales lead to implicit concentration bounds when applied Theorem 4.4, and are thus unfavorable when one needs a confidence interval or confidence sequence.
-However, these supermartingales can be used to test whether $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ or not and are thus “test supermartingales” only.
-4.3 Matrix e-Processes
-Scalar e-processes are an important class of processes generalizing scalar nonnegative supermartingales [Ramdas et al., 2022].
-We recall the definition of these processes below, noting that we will always consider a sequence of observations $X_{1},X_{2},\dots$ (while allowing their distribution to vary) and denote their canonical filtration by $\{\mathcal{F}_{n}\}$.
+Let us briefly explain the name “empirical Bernstein”.
+The original scalar version of the inequality above (that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$) was proved by Howard et al.
+[2021, Theorem 4], generalized and studied closer by Waudby-Smith and Ramdas [2023, Theorem 2], producing a confidence sequence for the common conditional mean of martingale dependent, $[0,1]$-valued scalars that is asymptotically as sharp as the Bernstein inequality without assuming knowing the underlying variance while adapting to it, and beats numerous then-existing empirical Bernstein inequalities in the literature.
+The matrix inequality that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ above was recently proved by Wang and Ramdas [2024, Lemma 4.1], who applied the same scalarization technique as Howard et al.
+[2020] (see Section 5.2) to obtain a scalar test supermartingale and a closed-form spectral empirical Bernstein inequality for $\mathcal{S}_{d}^{[0,1]}$-valued random matrices, of similar sharpness.
+Here, on the other hand, we demonstrate that if a closed-form spectral confidence set is not of interest, one can sequentially test the null $\mathbb{E}[\mathbf{X}_{n}|\mathcal{F}_{n-1}]=\mathbf{M}$, under the assumption that $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$, by constructing a matrix-valued test supermartingale using Lemma 3.5 of potential larger power compared to the scalarization approach of Wang and Ramdas [2024].
+The hyperparameters (predictable scalars $\{\gamma_{n}\}$ and matrices $\{\widehat{\mathbf{X}}_{n}\}$) in Example 3.10 can be set according to the discussion of Wang and Ramdas [2024, Section 4].
+Our last example generalizes the scalar “Catoni supermartingales” due to Catoni [2012] in the fixed-time case and Wang and Ramdas [2023a] in the sequential case.
+With scalars, this influence function-based technique produces subGaussian confidence intervals and sequences for heavy-tailed random variables with a $p$th moment upper bound where $p\in(1,2]$.
+In the $p=2$ special case Wang and Ramdas [2023a] demonstrate that the “Catoni-style confidence sequence” is tighter than the self-normalization technique which we already generalized to matrices in Example 3.8, under a variance upper bound assumption.
+Our matrix extension of these results starts with the standard definition of Catoni’s influence function (see e.g.
+Chen et al.
+[2021, Section 1], Wang and Ramdas [2023a, Section 9]).
+Definition 3.11.
+An increasing function $\phi:\mathbb{R}\to\mathbb{R}$ is upper $p$-logarithmically contractive if $\phi(x)\leqslant\log(1+x+|x|^{p}/p)$; lower $p$-logarithmically contractive if $-\log(1-x+|x|^{p}/p)\leqslant\phi(x)$.
+For example, the function
+$\phi_{p}(x)=\begin{cases}\log(1+x+x^{p}/p),&x\geqslant 0;\\ -\log(1-x+(-x)^{p}/p),&x<0.\end{cases}$ (27)
+is both upper and lower $p$-logarithmically contraction.
+Example 3.12 (Matrix Catoni Supermartingales).
+Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional $p$th moment upper bounds $\mathbb{E}((\operatorname{abs}(\mathbf{X}_{n}-\mathbf{M}_{n}))^{p}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ that are predictable.
+For any upper $p$-logarithmically contractive function, we can take in Lemma 3.5,
+$\mathbf{E}_{n}=\exp(\phi(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n}))),\quad\mathbf{F}_{n}=\left(\mathbf{I}+\frac{\gamma_{n}^{p}}{p}\mathbf{V}_{n}\right)^{-1};$ (28)
+and for any lower $p$-logarithmically contractive function, we can take in Lemma 3.5,
+$\mathbf{E}_{n}=\exp(-\phi(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n}))),\quad\mathbf{F}_{n}=\left(\mathbf{I}+\frac{\gamma_{n}^{p}}{p}\mathbf{V}_{n}\right)^{-1}.$ (29)
+Above, $\{\gamma_{n}\}$ are predictable positive scalars.
+Example 3.12 is proved in Section B.2.
+We remark that, in the scalar case, the two supermartingales constructed this way remain supermartingales under one-sided nulls, one for each side.
+This fails to hold with matrices (e.g., the null that $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{M}_{n}$) as there is no guarantee that $\exp(\pm\phi(\cdot))$ are operator monotone.
+Still, the two matrix test supermartingales play the role of detecting null-deviating behavior of opposite directions, and a powerful test for the point null $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ can be formed by a convex combination of the two.
+3.4 Matrix e-Values and e-Processes
+We have thus far generalized scalar nonnegative supermartingales to PSD matrix supermartingales, as well as their numerous applications in sequential statistics as “test supermartingales”.
+However, with scalars, it is known that nonnegative supermartingales fail to be powerful in some examples involving large composite null hypotheses, and a wider class of processes called e-processes is often used instead [Ramdas et al., 2022].
+Loosely speaking, an e-process is a nonnegative process upper bounded by supermartingales.
+We first recall the formal definitions of scalar e-values and e-processes, and generalize them to matrices as well.
+Throughout this subsection, we consider $X_{1},X_{2},\dots$ a sequence of random observations taking values in some space $\mathcal{X}$ and they generate the filtration $\{\mathcal{F}_{n}\}$.
 We use notations $\mathbb{P}_{P}$, $\mathbb{E}_{P}$ and the general terminology “for $P$” (e.g.
-we may say a stochastic process adapted to $\{\mathcal{F}_{n}\}$ is a martingale for $P$) to refer to probabilistic statements when the observations $X_{1},X_{2},\dots$ are distributed jointly according to $P$.
-Let $\mathcal{P}$ be a set of distributions over $\mathbb{R}\times\mathbb{R}\times\dots$.
-A nonnegative stochastic process $\{Y_{n}\}$ adapted to $\{\mathcal{F}_{n}\}$ is called an e-process for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, there exists a supermartingale $\{M_{n}^{P}\}$ adapted to $\{\mathcal{F}_{n}\}$ with $M_{0}^{P}\leqslant 1$ that upper bounds the process $\{Y_{n}\}$.
-Equivalently, $\{Y_{n}\}$ is an e-process if $\mathbb{E}_{P}[Y_{\tau}]\leqslant 1$ for any stopping time $\tau$ on $\{\mathcal{F}_{n}\}$ [Ramdas et al., 2020, Lemma 6].
-Clearly, if $\{Y_{n}\}$ is a nonnegative supermartingale for $\mathcal{P}$ (meaning for every $P\in\mathcal{P}$), the optional stopping theorem implies that it is also an e-process for $\mathcal{P}$, but e-processes are a much more general class [Ramdas et al., 2022, Ruf et al., 2023].
-A scalar e-process satisfies the following form of (VI) or (UVI):
-for $\text{for any }P\in\mathcal{P},\mathbb{P}_{P}(Y_{\tau}\geqslant Ua)\leqslant a^{-1}$
-for any stopping time $\tau$ and independent super-uniform random variable $U$ (and thus we also have that $\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$ by setting $U=1$).
-This property enables scalar e-processes to be used for testing a hypothesis that the data is drawn from some $P\in\mathcal{P}$: we may reject the null as soon as the e-process exceeds $1/\alpha$ (for some prespecified constant $\alpha\in(0,1)$) and be guaranteed that the type-I error is at most $\alpha$.
+we may say an adapted stochastic process is a martingale for $P$) to refer to probabilistic statements when the observations $X_{1},X_{2},\dots$ are distributed jointly according to $P$, where $P$ is some distribution on $\mathcal{X}\times\mathcal{X}\times\dots$.
+Let $\mathcal{P}$ be a set of distributions on $\mathcal{X}\times\mathcal{X}\times\dots$.
+An $\mathcal{F}_{\infty}$-measurable nonnegative random variable $Y$ is called an e-value for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, its mean $\mathbb{E}_{P}Y\leqslant 1$.
+An adapted nonnegative process $\{Y_{n}\}$ is called an e-process for $\mathcal{P}$ if, for any stopping time $\tau$, the stopped variable $Y_{\tau}$ is an e-value for $\mathcal{P}$.
+It is known that, due to Ramdas et al.
+[2020, Lemma 6], $\{Y_{n}\}$ is an e-process if and only if for any $P\in\mathcal{P}$ it is upper bounded by a supermartingale $\{M_{n}^{P}\}$ with $M_{0}^{P}\leqslant 1$.
+Clearly, if $\{Y_{n}\}$ is a nonnegative supermartingale for $\mathcal{P}$ (i.e.
+for every $P\in\mathcal{P}$), the optional stopping theorem implies that it is also an e-process for $\mathcal{P}$, but e-processes are a much more general class [Ramdas et al., 2022, Ruf et al., 2023].
+Scalar e-processes also satisfy (VI), both time-uniformly and at a stop time:
+$\text{For any $P\in\mathcal{P}$}:\quad\mathbb{P}_{P}(\exists n,\ Y_{n}\geqslant a)\leqslant a^{-1};$ (30)
+and for any stopping time $\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}.$ (31)
+This property enables scalar e-processes to also be used for testing the null hypothesis that the data is drawn from some $P\in\mathcal{P}$: we may reject the null as soon as the e-process exceeds $1/\alpha$ (for some prespecified constant $\alpha\in(0,1)$) and be guaranteed that the type-I error is at most $\alpha$.
 E-processes are particularly useful for composite sequential testing problems for which nonnegative supermartingales are powerless [Ramdas et al., 2022, Ruf et al., 2023].
-We can generalize the scalar e-processes to our matrix setting as well, with the sequence of observations that generates the filtration $\{\mathcal{F}_{n}\}$ now denoted by $\mathbf{X}_{1},\mathbf{X}_{2},\dots$.
+We define matrix e-values and e-processes similarly.
+Definition 3.13 (Matrix e-Value).
+Let $\mathcal{P}$ be a set of distributions on $\mathcal{X}\times\mathcal{X}\times\dots$.
+An $\mathcal{S}_{d}^{+}$-valued, $\mathcal{F}_{\infty}$-measurable random matrix $\mathbf{Y}$ is called a (matrix) e-value for $\mathcal{P}$ if $\mathbb{E}_{P}\mathbf{Y}\preceq\mathbf{I}$ for any $P\in\mathcal{P}$.
 Definition 3.14 (Matrix e-Process).
-Let $\mathfrak{P}$ be a set of distributions over $\mathcal{S}_{d}\times\mathcal{S}_{d}\times\dots$.
-An $\mathcal{S}_{d}^{+}$-valued stochastic process $\{\mathbf{Y}_{n}\}$ adapted to $\{\mathcal{F}_{n}\}$ is called a matrix e-process for $\mathfrak{P}$ if, for any $\mathfrak{p}\in\mathfrak{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{\mathfrak{p}}\}$ adapted to $\{\mathcal{F}_{n}\}$ with $\mathbf{M}_{0}^{\mathfrak{p}}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{\mathfrak{p}}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $\mathfrak{p}$.
-Similar to the scalar case, matrix e-processes satisfy their (UMVI).
-Lemma 4.12 (Uniformly Randomized Matrix Ville’s Inequality for Matrix e-Processes).
-Let $\{\mathbf{Y}_{n}\}$ be a matrix e-process for $\mathfrak{P}$.
-Then, for any $\mathfrak{p}\in\mathfrak{P}$, for any stopping time $\tau$ on $\{\mathcal{F}_{n}\}$, any trace super-uniform random matrix $\mathbf{U}$ taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ and any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
-$\mathbb{P}_{\mathfrak{p}}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1}).$ (40)
-Proof.
-Consider the upper bounding matrix supermartingale $\{\mathbf{M}_{n}^{\mathfrak{p}}\}$.
-By (UMVI), $\mathbb{P}(\mathbf{M}_{\tau}^{\mathfrak{p}}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1})$.
-Since $\mathbf{M}_{n}^{\mathfrak{p}}\succeq\mathbf{Y}_{n}$ almost surely for any $n$, $\mathbf{M}_{\tau}^{\mathfrak{p}}\succeq\mathbf{Y}_{\tau}$ almost surely as well.
-Therefore, $\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathbb{P}(\mathbf{M}_{\tau}^{\mathfrak{p}}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1}).$
-Simple examples of matrix e-processes following the previous subsection include lower bounds of $\mathcal{S}_{d}^{+}$-valued supermartingales.
-First, we define the following “matrix minimum”: for $\mathbf{A},\mathbf{B}\in\mathcal{S}_{d}$, we define $\mathbf{A}\curlyvee\mathbf{B}$ to be $\mathbf{A}$ if $\mathbf{A}\preceq\mathbf{B}$, and $\mathbf{A}-\lambda_{\max}(\mathbf{A}-\mathbf{B})\cdot\mathbf{I}$ if otherwise.
-Note that $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{A}$ and $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{B}$, and we remark that in general there is no largest symmetric matrix lower bound of $\mathbf{A}$ and $\mathbf{B}$ in the $\preceq$ order [Kadison, 1951, Theorem 6].
+Let $\mathcal{P}$ be a set of distributions on $\mathcal{X}\times\mathcal{X}\times\dots$.
+An $\mathcal{S}_{d}^{+}$-valued adapted stochastic process $\{\mathbf{Y}_{n}\}$ is called a (matrix) e-process for $\mathcal{P}$ if, for any stopping time $\tau$, the stopped random matrix $\mathbf{Y}_{\tau}$ is a matrix e-value for $\mathcal{P}$.
+It can be seen from the matrix optional stopping theorem (Theorem 3.3) that if $\{\mathbf{Y}_{n}\}$ is a PSD supermartingale for $\mathcal{P}$ then it is also an e-process for $\mathcal{P}$.
+More generally, $\{\mathbf{Y}_{n}\}$ is a matrix e-process for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{P}\}$ with $\mathbf{M}_{0}^{P}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{P}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $P$.
+Similar to the scalar case above, matrix e-processes satisfy (MVI), both time-uniform and at a stopping time:
+$\text{For any $P\in\mathcal{P}$}:\quad\mathbb{P}_{P}(\exists n,\ \mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1});$ (32)
+and for any stopping time $\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1}),$ (33)
+which follows directly from the proof of Theorem 3.4. Therefore, one can run a procedure similar to Algorithm 1 with an e-process instead for sequential hypothesis testing.
+Simple examples of matrix e-processes following the previous subsection therefore include lower bounds of $\mathcal{S}_{d}^{+}$-valued supermartingales.
+In the scalar case, an e-process is often constructed for a composite null hypothesis by taking an infimum over the test supermartingales under each of the null distributions.
+While a similar construction carries over to the matrices, it is important to note that in general the matrix minimum does not exist: for most pairs $\mathbf{A},\mathbf{B}\in\mathcal{S}_{d}$, there is no largest lower bound of $\mathbf{A}$ and $\mathbf{B}$ in the $\preceq$ order within $\mathcal{S}_{d}$ [Kadison, 1951, Theorem 6].
+However, we can define a matrix minimum, for example as follows:
+Definition 3.15.
+We define the following minimum operation $\curlyvee:\mathcal{S}_{d}\times\mathcal{S}_{d}\to\mathcal{S}_{d}$:
+$\mathbf{A}\curlyvee\mathbf{B}=\begin{cases}\mathbf{A},&\text{if }\mathbf{A}\preceq\mathbf{B};\\ \mathbf{A}-\lambda_{\max}(\mathbf{A}-\mathbf{B})\cdot\mathbf{I},&\text{otherwise}.\end{cases}$ (34)
+Note that $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{A}$ and $\mathbf{A}\curlyvee\mathbf{B}\preceq\mathbf{B}$, and that in general $\mathbf{A}\curlyvee\mathbf{B}\neq\mathbf{B}\curlyvee\mathbf{A}$.
+[…]
+$\mathbf{C}\preceq\mathbf{A},\mathbf{C}\preceq\mathbf{B},\mathbf{C}\npreceq\mathbf{A}\curlyvee\mathbf{B}.$ (35)
+For example, $\mathbf{B}\curlyvee\mathbf{A}$ is often an example of $\mathbf{C}$ that satisfies above.
+On the other hand, however, it is easy to verify that there does not exist a matrix $\mathbf{C}$ that satisfies
+$\mathbf{C}\preceq\mathbf{A},\mathbf{C}\preceq\mathbf{B},\mathbf{C}\succ\mathbf{A}\curlyvee\mathbf{B},$ (36)
+in the sense that there are several choices of “matrix minima”, Definition 3.15 being one among them, and these choices do not dominate one another.
+Finally, we note that the matrix minimum enables us to construct matrix e-processes for null hypotheses that are finite unions of several conditions under which a matrix test supermartingale is constructible:
+Proposition 3.16.
+Suppose $\mathcal{P}=\bigcup_{k=1}^{K}\mathcal{P}_{k}$ and $\{\mathbf{Y}^{k}_{n}\}$ is a matrix e-process for $\mathcal{P}_{k}$.
+Then, $\{\mathbf{Y}_{n}^{1}\curlyvee\mathbf{Y}_{n}^{2}\curlyvee\dots\curlyvee\mathbf{Y}_{n}^{K}\}$ is a matrix e-process for $\mathcal{P}$.
+Here, the $K$-fold minimum $\mathbf{Y}_{n}^{1}\curlyvee\mathbf{Y}_{n}^{2}\curlyvee\dots\curlyvee\mathbf{Y}_{n}^{K}$ can be defined either left-associatively as $(\dots(\mathbf{Y}_{n}^{1}\curlyvee\mathbf{Y}_{n}^{2})\curlyvee\dots)\curlyvee\mathbf{Y}_{n}^{K}$ or right-associatively as $\mathbf{Y}_{n}^{1}\curlyvee(\dots\curlyvee(\mathbf{Y}_{n}^{K-1}\curlyvee\mathbf{Y}_{n}^{K})\dots)$.
+These lead to two different e-processes.
+One can take the average over all $K!$ permutations if one desires an e-process that is symmetric w.r.t.
+all $K$ subclasses.
+A simple example is as follows.
 Let $\mathbf{M},\mathbf{M}^{\prime}\preceq\mathbf{B}$ be three matrices in $\mathcal{S}_{d}^{++}$, and we define $\mathcal{B}(\mathbf{B},\mathbf{M})$ to be the set of distributions on $\mathcal{S}_{d}^{+}$ upper bounded by $\mathbf{B}$ with mean $\mathbf{M}$.
 Further, we define the distribution families
-$\mathfrak{P}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M})\text{ for all }n\}$ (41)
+$\mathcal{P}_{1}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M})\text{ for all }n\}$ (37)
 and
-$\mathfrak{P}^{\prime}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M}^{\prime})\text{ for all }n\}.$ (42)
-According to Example 4.7, we can construct a $\mathcal{S}_{d}^{+}$-valued betting process $\{\mathbf{Y}_{n}\}$ which is a supermartingale under any of $\mathfrak{p}\in\mathfrak{P}$, and another $\{\mathbf{Y}_{n}^{\prime}\}$ which is a supermartingale under any of $\mathfrak{p}\in\mathfrak{P}^{\prime}$.
-Then the minimum process $\{\mathbf{Y}_{n}\curlyvee\mathbf{Y}_{n}^{\prime}\}$ is a matrix $\mathfrak{P}\cup\mathfrak{P}^{\prime}$-e-process, useful for testing sequentially the null that the true mean is always either $\mathbf{M}$ or $\mathbf{M}^{\prime}$.
-4.4 Matrix-Valued Submartingales and Backward Supermartingales
-Next, we note that all the above are still valid if we replace “supermartingale”s with “submartingale”s.
-To wit, we say that $\{\mathbf{Y}_{n}\}$ is a submartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq\mathbf{Y}_{n-1}$.
-It is clear from the proof of Lemma 4.3 that,
+$\mathcal{P}_{2}=\{P_{1}\otimes P_{2}\otimes\dots:P_{n}\in\mathcal{B}(\mathbf{B},\mathbf{M}^{\prime})\text{ for all }n\}.$ (38)
+According to Example 3.7, we can construct a $\mathcal{S}_{d}^{+}$-valued betting process $\{\mathbf{Y}_{n}\}$ which is a supermartingale under any of $P\in\mathcal{P}_{1}$, and another $\{\mathbf{Y}_{n}^{\prime}\}$ which is a supermartingale under any of $P\in\mathcal{P}_{2}$.
+Then the minimum process $\{\mathbf{Y}_{n}\curlyvee\mathbf{Y}_{n}^{\prime}\}$ is a matrix e-process for $\mathcal{P}_{1}\cup\mathcal{P}_{2}$, useful for testing sequentially the null that the true mean is always either $\mathbf{M}$ or $\mathbf{M}^{\prime}$.
+4 Backward PSD Submartingales
+We have so far discussed the properties and applications $\mathcal{S}_{d}^{+}$-valued supermartingales on a forward filtration $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
+We now move on to a backward filtration $\{\mathcal{G}_{0}\}_{n\geqslant 0}$ and its adapted $\mathcal{S}_{d}^{+}$-valued submartingales to extend the established theory with scalars.
+This involves first considering a forward submartingale of finite length on the finite forward filtration $\{\mathcal{F}_{n}\}_{0\leqslant n\leqslant N}$, inverting the time, and extending indefinitely into the “past” (before the inversion).
+For this reason, we shall always be explicit with the filtration and the time range in this section.
+4.1 Submartingales and Maximal Inequalities: From Forward to Backward
+In the proof of Theorem 3.3, we appeal to the simple fact that if $\{\mathbf{Y}_{n}\}$ is a matrix supermartingale then $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a scalar supermartingale, which enables us to derive stopping time theorems for matrix supermartingales via those for scalar supermartingales.
+This clearly remains still valid if we replace “supermartingale”s with “submartingale”s.
+To wit, we say that $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ adapted to $\{\mathcal{F}_{n}\}_{n\geqslant 0}$ is a matrix submartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq\mathbf{Y}_{n-1}$.
+It is clear from the proof of Theorem 3.3 that,
 Lemma 4.1.
-Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued submartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$.
-Then, $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a submartingale on $\{\mathcal{F}_{n}\}$.
+Let $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ be an $\mathcal{S}_{d}$-valued submartingale, and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
+Then, $\{\mathbf{Y}_{n\wedge\tau}\}_{n\geqslant 0}$ is also a submartingale on $\{\mathcal{F}_{n}\}_{n\geqslant 0}$.
 Consequently, if $\tau\leqslant N$ almost surely for some $N\in\mathbb{N}$, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{N}$.
@@
 $\{\mathbf{Y}_{\tau}\mathbbmss{1}_{E}\npreceq\mathbf{A}\}=E,\quad\mathbf{Y}_{\tau}\mathbbmss{1}_{\overline{E}}=\mathbf{Y}_{N}\mathbbmss{1}_{\overline{E}}.$ (41)
-Thus by Lemmas 4.13 and 3.3,
+Thus by Lemmas 4.1 and 5.3,
 $\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{N})\mathbf{A}^{-1/2})\geqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})$ (42)
@@
 implying that $\mathbb{P}(E)\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}\mathbb{E}(\mathbf{Y}_{N}\mathbbmss{1}_{E})\mathbf{A}^{-1/2})$. ∎
-Next, we illustrate how Theorem 4.14 remains valid even without assuming that the process $\{\mathbf{Y}_{n}\}$ starts at $n=0$ — or starts at all; it is applicable even if $\{\mathbf{Y}_{n}\}$ is indexed on $n\in\{\dots,-2,-1,0,1,\dots,N\}$.
-We formalize this by considering a backward filtration $\{\mathcal{G}_{n}\}$, i.e., $\mathcal{G}_{0}\supseteq\mathcal{G}_{1}\supseteq\dots$, and an adapted backward submartingale $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{Y}_{n}$ is $\mathcal{G}_{n}$-measurable and $\mathbb{E}[\mathbf{Y}_{n}|\mathcal{G}_{n+1}]\succeq\mathbf{Y}_{n+1}$, while keeping the time indices $n\geqslant 0$.
-This, in the scalar regime, sets the stage for the backward Ville’s inequality which has been shown equivalent to the (forward) Doob’s inequality (hence the duality between Ville’s and Doob’s inequality; see the second proof of Theorem 2 in Manole and Ramdas [2023]), for which we have the following matrix version.
+The backward Ville’s inequality comes from the idea that Doob’s inequality remains valid even without assuming that time starts at $n=0$ — or that it starts at all.
+In the scalar regime, this idea proves the backward Ville’s inequality (7) and proves that it is equivalent to the (forward) Doob’s inequality (hence the “duality” between Ville’s and Doob’s inequality; see the second proof of Theorem 2 in Manole and Ramdas [2023]).
+This idea is valid for matrices as well; that is, Theorem 4.2 is applicable even if $\{\mathbf{Y}_{n}\}$ and $\{\mathcal{F}_{n}\}$ are indexed on $n\in\{\dots,-2,-1,0,1,\dots,N\}$.
+We formalize this by recalling that a backward filtration $\{\mathcal{G}_{n}\}$ is one that satisfies $\mathcal{G}_{0}\supseteq\mathcal{G}_{1}\supseteq\dots$ and an adapted backward submartingale $\{\mathbf{Y}_{n}\}$ is one such that each $\mathbf{Y}_{n}$ is $\mathcal{G}_{n}$-measurable and $\mathbb{E}[\mathbf{Y}_{n}|\mathcal{G}_{n+1}]\succeq\mathbf{Y}_{n+1}$, while we formally keep the time indices $n\geqslant 0$.
 Theorem 4.3 (Matrix Backward Ville’s Inequality).
-Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued backward submartingale adapted to the backward filtration $\{\mathcal{G}_{n}\}$. For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
+Let $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ be an $\mathcal{S}_{d}^{+}$-valued backward submartingale adapted to the backward filtration $\{\mathcal{G}_{n}\}_{n\geqslant 0}$.
+For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
 $\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (45)
@@
 Take an arbitrary $N\in\mathbb{N}$ and consider, indexed by $0\leqslant n\leqslant N$, the process $\mathbf{P}_{n}=\mathbf{Y}_{N-n}$ on the (forward) filtration $\mathcal{F}_{n}=\mathcal{G}_{N-n}$.
-It is clear that $\{\mathbf{P}_{n}\}$ is a (forward) submartingale on $\{\mathcal{F}_{n}\}$, to which we can apply Theorem 4.14,
+It is clear that $\{\mathbf{P}_{n}\}_{0\leqslant n\leqslant N}$ is a (forward) submartingale on $\{\mathcal{F}_{n}\}_{0\leqslant n\leqslant N}$, to which we can apply Theorem 4.2,
 $\mathbb{P}\left(\exists n\leqslant N,\mathbf{P}_{n}\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{P}_{N})\mathbf{A}^{-1}).$ (46)
@@
 The desired inequality is immediately proved by appealing to the fact that
-$\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}=\liminf_{N\to\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}=\bigcup_{k=1}^{\infty}\bigcap_{N=k}^{\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}.$ (52)
-4.5 Exchangeable Matrices
-We now have the tools to prove a matrix version of the exchangeable Chebyshev inequality (XCI) we mentioned earlier.
-Recall from Fact 2 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $21\leqslant p<$ 2) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
+$\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}=\liminf_{N\to\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}=\bigcup_{k=1}^{\infty}\bigcap_{N=k}^{\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}.\qed$ (48)
+4.2 Application: Concentration of Exchangeable Matrices
+An immediate application of Theorem 4.3 is numerous matrix Chebyshev-type concentration bounds for exchangeable random matrices.
+We recall that a sequence of random variables $X_{1},X_{2},\dots$ taking values in a Polish space is exchangeable if the joint distribution of $(X_{i_{1}},\dots,X_{i_{k}})$ equals that of $(X_{j_{1}},\dots,X_{j_{k}})$ for any distinct $i_{1},\dots,i_{k}$ and distinct $j_{1},\dots,j_{k}$.
+Repetitions of a single random variable, the Pólya urn model, sampling from a finite set without replacement, as well as i.i.d.
+random variables are all examples of exchangeable sequences.
+The scalar exchangeable Chebyshev inequality was recently proved by Ramdas and Manole [2024, Theorem 2].
+Recall from Fact 1 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $21\leqslant p<$ 2) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
 $\frac{1}{n}(\mathbf{X}_{1}^{p}+\dots+\mathbf{X}_{n}^{p})\succeq\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}\right)^{p}.$ (49)
@@
 concluding the proof that $\{\mathbf{R}_{n}\}$ is a backward submartingale. ∎
-Applying the matrix backward Ville’s inequality (Theorem 4.15) to this process $\{\mathbf{R}_{n}\}$, we immediately have the following two matrix extensions of (XCI).
+Applying the matrix backward Ville’s inequality (Theorem 4.3) to this process $\{\mathbf{R}_{n}\}$, we immediately have the following two exchangeable matrix Chebyshev inequalities.
+First, the following holds when a finite variance exists.
+Recall that the anticommutator between two sqaure matrices $\mathbf{A},\mathbf{B}$ is defined as $\{\mathbf{A},\mathbf{B}\}=\mathbf{A}\mathbf{B}+\mathbf{B}\mathbf{A}$.
 Theorem 4.5 (Exchangeable Matrix Chebyshev Inequality).
@@
 Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
-$\mathbb{P}(\exists n,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (XMCI)
+$\mathbb{P}(\exists n,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (54)
 Further, if $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$, we have for any fixed $N$,
@@
 $\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{R}_{1})\mathbf{A}^{-2})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}).$ (56)
-Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/2}$,
+Now we use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/2}$ (Fact 1),
 $\mathbb{P}(\exists n,\;\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2}),$ (57)
-concluding the proof of (XMCI). To show (58), simply apply Theorem 4.15 to the submartingale $\mathbf{R}_{n}=(\overline{\mathbf{X}_{n}}-\mathbf{M})^{2}$ with $n\geqslant N$, obtaining
+concluding the proof of (54). To show (55), simply apply Theorem 4.3 to the submartingale $\mathbf{R}_{n}=(\overline{\mathbf{X}_{n}}-\mathbf{M})^{2}$ with $n\geqslant N$, obtaining
 $\mathbb{P}(\exists n\geqslant N,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{R}_{N})\mathbf{A}^{-2})\leqslant N^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ (58)
-due to Theorem 3.5. ∎
+where the last step is due to Lemma B.2. ∎
+We remark that the assumption $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for $i\neq j$ is satisfied when $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ are pairwise uncorrelated, which is in turn satisfied under martingale difference assumption (i.e., $\{\mathbf{X}_{i}\}$ is adapted to some filtration $\{\mathcal{F}_{i}\}$ with $\mathbb{E}(\mathbf{X}_{i}|\mathcal{F}_{i-1})=\mathbf{M}$ and $\mathbf{Var}(\mathbf{X}_{i}|\mathcal{F}_{i-1})\preceq\mathbf{V}$), which is in turn implied if they are i.i.d.
+The inequality (54) generalizes the matrix Chebyshev inequality by Ahlswede and Winter [2002, Theorem 14] which states that $\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}-\mathbf{M})\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ for one square-integrable matrix $\mathbf{X}$ with variance $\mathbf{V}$.
+It is only by assuming $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ that we may obtain a “shrinking” factor $N^{-1}$ on the tail probability that shows up in the right hand side of (55).
+Without this assumption, these exchangeable matrices can all be the repetition of a single random matrix $\mathbf{X}_{1}=\mathbf{X}_{2}=\dots$ in which case (54) is reduced to the matrix Chebyshev inequality by Ahlswede and Winter [2002, Theorem 14].
+On the other hand, with the stated additional assumption $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$, the inequality (55) generalizes the matrix “weak law of large numbers” by Ahlswede and Winter [2002, Corollary 16] which we already quoted as (10), that $\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{N}-\mathbf{M})\npreceq\mathbf{A})\leqslant N^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ for $N$ i.i.d.
+matrices $X_{1},\dots,X_{N}$.
+Both generalizations, we note, offer the exact same concentration at the sample sizes (1 and $N$ respectively) compared to the original results by Ahlswede and Winter [2002], while also extending infinitely into all future sample sizes.
+These generalizations generalize those in the scalar case by Ramdas and Manole [2024, Section 2.2].
+We next present results for exchangeable matrices with heavier tails.
 Theorem 4.6 (Exchangeable Matrix $p$-Chebyshev Inequality).
@@
 Proof.
-Applying Theorem 4.15 to $\{\mathbf{R}_{n}\}$, with $\mathbf{C}$ being 00, in Lemma 4.16,
+We define $\mathbf{R}_{n}=(\overline{\mathbf{X}}_{n})^{p}$, which, since $p\in[1,2]$, is a matrix backward submartingale due to Lemma 4.4. Applying Theorem 4.3 to $\{\mathbf{R}_{n}\}$, we see that
 $\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n})^{p}\npreceq\mathbf{A}^{p})\leqslant\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (60)
-Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/p}$,
+Now use the operator monotonicity of $\mathbf{X}\mapsto\mathbf{X}^{1/p}$ (Fact 1),
 $\mathbb{P}(\exists n,\;\overline{\mathbf{X}}_{n}\npreceq\mathbf{A})\leqslant\mathbb{P}(\exists n,\;(\overline{\mathbf{X}}_{n})^{p}\npreceq\mathbf{A}^{p}),$ (61)
 concluding the proof. ∎
-Indeed, we can weaken the second-moment assumption to a $p$th moment, leading to, however, a concentration not around the mean but a Markov-type upper bound.
-This is largely because the scalar map $x\mapsto|x|^{p}$ is convex but not its matrix extension, and so the raw $p$th moment is used above.
-Nonetheless, due to Fact 4, the following trace inequality holds.
-$\mathrm{tr}\left(\frac{1}{n}((\operatorname{abs}\mathbf{X}_{1})^{p}+\dots+(\operatorname{abs}\mathbf{X}_{n})^{p})\right)\geqslant\mathrm{tr}\left(\left(\operatorname{abs}\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}\right)^{p}\right).$ (65)
-Therefore,
-$r_{n}=\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{C})^{p}\right)$ (66)
-is a real-valued backward submartingale. With the scalar backward Ville’s inequality, this leads to the following anytime-valid extension of Lemma 3.7.
+We remark that, from Theorem 4.5 to Theorem 4.6, we weaken the second moment to a $p$th moment assumption, but in an imperfect manner as we obtain a concentration not around the mean $\mathbf{M}$ but a Markov-type “upper tail bound” for $\mathcal{S}_{d}^{+}$-valued matrices instead.
+This is largely because the scalar map $x\mapsto|x|^{p}$ is convex on (,)($\infty$,$\infty$)( - ∞ , ∞ ) but not its matrix extension, and so the raw $p$th moment is used above.
+Also, Theorem 4.6 fails if $p>2$ because the scalar map $x\mapsto x^{p}$ is convex on $[0,\infty)$ but not its matrix extension.
+Nonetheless, thanks to Fact 3, we can resort to trace scalarization to obtain the following spectral and trace inequalities that hold for all $p\geqslant 1$.
 Theorem 4.7 (Exchangeable Matrix $p$-Chebyshev Trace Inequality).
-Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with mean $\mathbf{M}$ and central $p$th moment $\mathbf{V}_{p}=\mathbb{E}\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M})^{p}$, where $p\geqslant 1$.
+Let $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with mean $\mathbf{M}$ and central $p$th moment $\mathbf{V}_{p}=\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M}))^{p}$, where $p\geqslant 1$.
 Then, for any $a>0$,
-$\mathbb{P}(\exists n,\;\|\overline{\mathbf{X}}_{n}-\mathbf{M}\|_{\mathrm{op}}\geqslant a)\leqslant\mathbb{P}\left(\exists n,\;\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})^{p}\right)\geqslant a^{p}\right)\leqslant a^{-p}\mathrm{tr}\mathbf{V}_{p}.$ (67)
-The bound above differs from Theorem 4.18 in that the concentration is now around the mean $\mathbf{M}$.
-5 Randomized Matrix Bounds via Scalar e-Processes
-5.1 A Conversion Lemma from Matrix to Scalar Supermartingales
-We have studied the concentration of random matrices via matrix e-statistics, i.e.
-random matrices with bounded expectations, matrix supermartingales, matrix e-values, and matrix backward submartingales, since these quantities yield concentration inequalities via our Markov and Ville’s inequalities for matrices, oftentimes in randomized forms.
-Some further randomized matrix concentration inequalities can be obtained by revisiting some bounds obtained via scalar e-quantities that already exist in the literature.
+$\mathbb{P}(\exists n,\;\|\overline{\mathbf{X}}_{n}-\mathbf{M}\|\geqslant a)\leqslant\mathbb{P}\left(\exists n,\;\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})^{p}\right)\geqslant a^{p}\right)\leqslant a^{-p}\mathrm{tr}\mathbf{V}_{p}.$ (62)
+The inequality above, which we defer its straightforward proof to Section B.3, differs from Theorem 4.6 in that the concentration is now around the mean $\mathbf{M}$.
+5 Remarks and Extensions
+5.1 Tightness of Matrix Markov’s and Ville’s Inequalities
+[…]
+$\mathcal{C}_{\mathrm{tr}}(\alpha)=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathrm{tr}\mathbf{X}\leqslant d/\alpha\};$ (63)
+by $\mathcal{C}_{\mathbf{v}}(\alpha)$ the concentration region by applying (MI) to $\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v}$ where $\|\mathbf{v}\|=1$:
+$\mathcal{C}_{\mathbf{v}}(\alpha)=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v}\leqslant 1/\alpha\}.$ (64)
+Let $\mathcal{A}_{\alpha}$ be the following set of matrices:
+$\mathcal{A}_{\alpha}=\{\mathbf{A}\in\mathcal{S}_{d}^{++}:\mathrm{tr}(\mathbf{A}^{-1})=\alpha\},$ (65)
+and we denote by $\mathcal{C}(\mathbf{A})$ the concentration region by applying (MMI) to $\mathbf{X}$ with threshold $\mathbf{A}\in\mathcal{A}_{\alpha}$:
+$\mathcal{C}({\mathbf{A}})=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathbf{X}\preceq\mathbf{A}\}.$ (66)
+Following e.g.
+Minsker [2017], we denote by $\operatorname{r}(\mathbf{A})$ the effective rank $\mathrm{tr}\mathbf{A}/\|\mathbf{A}\|\in[1,d]$ of a PD matrix.
+We can establish the following relations among these concentration regions via elementary linear algebra.
+Proposition 5.1.
+Suppose $\alpha\in(0,1)$ and $d\geqslant 2$.
+  1. 1.
+    $\mathcal{C}_{\mathrm{tr}}(\alpha)\subsetneq\mathcal{C}((d/\alpha)\mathbf{I})$;
+  2. 2.
+    If $\mathbf{A}\in\mathcal{A}_{\alpha}$ and $\mathbf{A}\neq(d/\alpha)\mathbf{I}$, then $\mathcal{C}_{\mathrm{tr}}(\alpha)\nsubseteq\mathcal{C}(\mathbf{A})$ and $\mathcal{C}(\mathbf{A})\nsubseteq\mathcal{C}_{\mathrm{tr}}(\alpha)$;
+  3. 3.
+    If $\mathbf{A}\in\mathcal{A}_{\alpha}$, then $\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})\subseteq\mathcal{C}(\mathbf{A})$ if and only if $\alpha^{\prime}\geqslant\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$;
+  4. 4.
+    If $\mathbf{A}\in\mathcal{A}_{\alpha}$, then $\mathcal{C}(\mathbf{A})\subseteq\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ if and only if $\alpha^{\prime}\leqslant d(\mathrm{tr}\mathbf{A})^{-1}\in[\frac{\alpha}{d},\frac{\alpha}{\operatorname{r}(\mathbf{A}^{-1})}]$;
+  5. 5.
+    For any $\mathbf{v}$ such that $\|\mathbf{v}\|=1$ and any $\alpha^{\prime}>\alpha$, there always exists an $\mathbf{A}\in\mathcal{A}_{\alpha^{\prime}}$ such that $\mathcal{C}(\mathbf{A})\subsetneq\mathcal{C}_{\mathbf{v}}(\alpha)$;
+  6. 6.
+    For any $\mathbf{A}\in\mathcal{A}_{\alpha}$ and any $\alpha^{\prime}\in(0,1)$, there exists no $\mathbf{v}$ such that $\|\mathbf{v}\|=1$ and $\mathcal{C}_{\mathbf{v}}(\alpha^{\prime})\subseteq\mathcal{C}(\mathbf{A})$.
+Let us briefly unpack these statements above.
+The first two state that (MMI) and trace-(MI) are mutually incomparable unless one takes the threshold matrix $\mathbf{A}$ to be a multiple of $\mathbf{I}$, in which case (MMI) is weaker than trace-(MI).
+Statements 3 and 4 indicate that (MMI) and trace-(MI) can simulate each other only by increasing the error probability $\alpha$, and that if $\mathbf{A}^{-1}$ is chosen to have a low effective rank, (MMI) can easily simulate trace-(MI) whereas trace-(MI) pays a price up to $d$ to simulate (MMI).
+Statements 5 and 6 imply respectively that, while one can always simulate quadratic form-(MI) with (MMI) by infinitesimally increasing the error probability $\alpha$, the converse is never true.
+Proposition 5.1 is proved in Section B.4.
+Similar reasoning holds for (MVI) versus (VI) as well because the concentration regions, while holding time-uniformly or at a stopping time, remain in the same forms of $\mathcal{C}_{\mathrm{tr}}(\alpha)$, $\mathcal{C}_{\mathbf{v}}(\alpha)$, and $\mathcal{C}(\mathbf{A})$.
+We shall spell out the statistical consequences of these linear algebra results in Section 6.
+5.2 Stepwise Scalarization
 As we mentioned earlier in Section 3.3, Howard et al.
-[2020, Section 6.5] prove, under different conditions, numerous inequalities for symmetric matrix martingale difference $\{\mathbf{Z}_{n}\}$ in the style of
+[2020, Section 6.5] prove, under different conditions, numerous inequalities for $\mathcal{S}_{d}$-valued martingale difference $\{\mathbf{Z}_{n}\}$ in the style of
 $\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\exp(\mathbf{C}_{n}^{\prime}),$ (67)
 for some respectively $\mathcal{F}_{n}$-, $\mathcal{F}_{n-1}$-measurable symmetric matrices $\mathbf{C}_{n},\mathbf{C}_{n}^{\prime}$.
-At this step, our previous methods and that of Howard et al.
-[2020] had diverged.
-We have been able to directly construct matrix-valued supermartingales in Section 4.2, leading potentially to computable sequential tests by virtue of our new (UMVI).
+At this step, our methods in Section 3 and that of Howard et al. [2020] diverged. We have been able to directly construct the matrix-valued supermartingale
+$\exp\left(-\frac{\mathbf{C}_{n}^{\prime}}{2}\right)\exp\left(\frac{\mathbf{Z}_{n}-\mathbf{C}_{n}}{2}\right)\dots\exp\left(\frac{\mathbf{Z}_{n}-\mathbf{C}_{n}}{2}\right)\exp\left(-\frac{\mathbf{C}_{n}^{\prime}}{2}\right)$ (68)
+via Lemma 3.5, leading to a computable sequential test by our new (MVI).
 Howard et al.
-[2020], on the other hand, construct scalar e-processes from the condition (68).
+[2020], on the other hand, construct scalar e-processes from the condition (67) by applying a trace scalarization technique.
 Let us rephrase their “master lemma” behind this construction.
@@
 $L_{n}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\mathbf{Z}_{i}-\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)$ (69)
-is a (scalar) supermartingale. Further,
+is a supermartingale. Further,
 $L_{n}\geqslant\exp\left(\lambda_{\max}\left(\sum_{i=1}^{n}\mathbf{Z}_{i}\right)-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\right).$ (70)
-The proof of this lemma can be found in Section B.5.
-Note that the upper bounded expression (70) is an instance of e-processes, introduced earlier in Section 4.3.
-We shall see soon that both (69) and (70) can be useful.
-5.2 Examples: Self-Normalization, Bernstein, and Hoeffding
-Recall from Example 4.8 the self-normalized bound, that if $\{\mathbf{X}_{n}\}$ are symmetric matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\{\mathbf{M}_{n}\}$ and predictable conditional variance upper bounds $\{\mathbf{V}_{n}\}$,
-$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\mid\mathcal{F}_{n-1}\right)\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (71)
-an instance of (68). By Lemma 5.1, this implies the following statement whose weaker version is stated as Lemma 3(f) by Howard et al. [2020].
-Example 5.2 (Matrix Self-Normalized Scalar Supermartingale).
-Under the same assumptions as Example 4.8, the process
-$L_{n}^{\textsf{SN}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)$ (72)
-is a supermartingale.
-The proof is simply taking $\mathbf{Z}_{n}=\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})$, $\mathbf{C}_{n}=\frac{\gamma_{n}^{2}}{6}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}$, and $\mathbf{C}_{n}^{\prime}=\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}$ in Lemma 5.1.
-Some remarks on this result are in order.
-First, the expression of the scalar supermartingale $\{L_{n}^{\textsf{SN}}\}$ explains the name “self-normalized”, as an empirical variance term exists within the bound.
-Second, unlike the scalar case where a closed-form confidence sequence can be solved from the supermartingale combining (VI) if the means are a constant [Wang and Ramdas, 2023a, Section 5], in our matrix case, we only have an implicit bound when applying (UVI).
-Corollary 5.3 (Uniformly Randomized Stopped Matrix Self-Normalized Bound).
-Under the same assumptions as Example 4.8, for any stopping time $\tau$, $\alpha\in(0,1)$, and a super-uniform $U$ independent from $\mathcal{F}_{\infty}$,
-$\mathbb{P}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{\tau}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{\tau}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)\geqslant dU/\alpha\right)\leqslant\alpha.$ (73)
-The bound above, as those matrix supermartingales in Section 4.2, does not bear closed-form confidence sets but still implies a computable randomized sequential test for the null, e.g., that the means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})$ all equal $\mathbf{M}$ with marginal type I error rate within $\alpha$.
-However, a closed-form concentration can be computed when further assumption is introduced on top of the finite variance in Example 4.8.
-This is the Hoeffding-type e-process by Howard et al.
-[2020, Lemma 3], which we present as a Bernstein-type result akin to Mackey et al.
-[2014, Corollary 4.2] for generality.
-Example 5.4 (Matrix Bernstein Scalar e-Process).
-Under the same assumptions as Example 4.8, suppose further that $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}\preceq\mathbf{B}_{n}$ almost surely for some predictable sequence $\{\mathbf{B}_{n}\}$.
-Then, the self-normalized supermartingale $\{L_{n}^{\textsf{SN}}\}$ upper bounds the following process:
-$L_{n}^{\textsf{SN}}\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)\right\}.$ (74)
+The proof of this lemma can be found in Section B.5. We have numerous remarks regarding this lemma.
+First, the upper bounded expression (70) is an instance of e-processes (if divided by $L_{0}=d$), introduced earlier in Section 3.4.
+The larger process (69), a supermartingale, can be used for a sequential test; whereas the smaller process (70), an e-process, leads to closed-form spectral time-uniform concentration inequalities (i.e.
+confidence sequences), but as a test process it is less powerful than (69).
+Second, the divergence from the common assumption (67) to the matrix-valued (68) and scalar-valued (69) test supermartingale indicates that for some testing problems, two different routes can be pursued: our “matrix-native” approach, and the scalarization approach to the test process.
+Examples in Section 3.3 that fit in the assumption (67) include Examples 3.8 and 3.9 which Howard et al.
+[2020, Section 6.5] originally prove, but also Example 3.10: To see that Example 3.10 is in the form of (67),
+$\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}))=\exp(\underbrace{\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})}_{:=\mathbf{Z}_{n}}+\underbrace{\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})}_{:=\mathbf{C}_{n}}),$ (71)
+which is exactly what is done by Wang and Ramdas [2024, Theorem 4.2] to derive a sharp closed-form matrix empirical Bernstein inequality.
+One of the crucial ways the two approaches differ is that with the scalar-valued test process (69), one rejects the null when the process surpasses $d/\alpha$; whereas with the matrix-valued test process (68), one rejects when the process $\npreceq\mathbf{A}$ where $\mathbf{A}$ is a matrix chosen freely beforehand, only subject to the “size constraint” that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$.
+Indeed, the extra degrees of freedom that come with $\mathbf{A}$ play a crucial role as we compare the two approaches empirically with the self-normalizing example (Example 3.8) in Section 6.
+Third, the assumption (67) for scalarization being more restrictive than our Lemma 3.5, for some testing problems our concept of matrix-valued test supermartingales is unavoidable.
+These include Examples 3.7 and 3.12, to which the stepwise scalarization Lemma 5.2 is not applicable.
+Their one-step condition $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}$ leads naturally to a matrix-valued test supermartingale, and scalar-valued test supermartingales might be obtained by taking the trace or quadratic form after the formation of the matrix-valued one, which is not ideal according to Section 5.1.
+5.3 Randomized Ahlswede-Winter Bounds
+The uniformly randomized Ville’s inequality (8) by Ramdas and Manole [2024] is obtained by applying the uniformly randomized Markov’s inequality [Ramdas and Manole, 2024, Theorem 1.2]
+$\mathbb{P}(X\leqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (72)
+where $U\sim\operatorname{Unif}_{(0,1)}$ independent from $X$. Here, we state a matrix version of (72), or a randomized version of the matrix Markov inequality (MMI).
+Theorem 5.3 (Uniformly Randomized Matrix Markov’s Inequality).
+Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{X}\subseteq\mathcal{S}_{d}^{+}$, and $U\sim\operatorname{Unif}_{(0,1)}$ independent from $\mathbf{X}$.
+For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
+$\mathbb{P}(\mathbf{X}\npreceq U\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}).$ (73)
+It is clear that (73) generalizes the scalar (72) and tightens (MMI).
+The randomization factor $U$, we note, can be any random variable that is stochastically larger than $\operatorname{Unif}_{(0,1)}$ (“super-uniform”), that is, $\mathbb{P}(U\leqslant x)\leqslant x$ for all $x\geqslant 0$.
+More generally, we shall present and prove a stronger version of Theorem 5.3 in Appendix A, where the scalar randomization factor $U$ is further replaced by a PSD random matrix whose distribution belongs to a class of “super-uniform” measures on $\mathcal{S}_{d}^{+}$.
+The generalized version of Theorem 5.3 in Appendix A also reveals the equality condition for (MMI), randomized or not.
+Theorem 5.3 implies that various matrix concentration inequalities by Ahlswede and Winter [2002] can be tightened via a randomization factor as well.
+For example, the Chebyshev inequality (10) for $n$ i.i.d.
+matrices can be randomized as follows.
+Corollary 5.4 (Uniformly Randomized Matrix Chebyshev’s Inequality).
+Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{i}=\mathbf{M}$ and variance matrix upper bound $\mathbf{Var}(\mathbf{X}_{i})\preceq\mathbf{V}$, such that $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$; $U$ a super-uniform random scalar independent from them all.
+Then, letting $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$.
+$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\sqrt{U}\mathbf{A})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}),$ (74)
+We omit the proof since it is a straightforward application of Theorem 5.3 to the moment bound Lemma B.2.
+We next discuss the implication for (MVI) as the randomization here raises some subtleties in its two equivalent forms, the time-uniform one and the stopped one.
+Theorem 5.5 (Uniformly Randomized Matrix Ville’s Inequality).
+Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale and $\tau$ be a stopping time on $\{\mathcal{F}_{n}\}$.
+Let $U$ be a super-uniform random scalar independent from $\mathcal{F}_{\infty}$.
+Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
+$\mathbb{P}(\mathbf{Y}_{\tau}\npreceq U\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1});$ (75)
+and
+$\mathbb{P}(\exists n<\tau,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\text{ or }\mathbf{Y}_{\tau}\npreceq U\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (76)
 Proof.
-Taking $\psi(x)=x^{2}/2$, $\mathbf{C}_{n}=\frac{1}{3}(\mathbf{X}_{n}-\mathbf{M})^{2}$, and $\mathbf{C}_{n}^{\prime}=\frac{2}{3}\mathbf{V}_{n}$ in Lemma 5.1, from (70),
-$L_{n}^{\textsf{SN}}$ $\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}((\mathbf{X}_{i}-\mathbf{M}_{i})^{2}+2\mathbf{V}_{i})\right)\right\}$ (75)
-$\geqslant\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\frac{\gamma_{i}^{2}}{6}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)\right\},$ (76)
-concluding the proof. ∎
-A closed-form, randomized, stopping time-valid matrix Bernstein or Hoeffding inequality on the concentration of weighted sample average is immediately in force by applying (UVI) to $\{L_{n}^{\textsf{SN}}\}$.
-Below, we denote by $\overline{\mathbf{X}}_{n}^{\gamma}$ the weighted average $\frac{\gamma_{1}\mathbf{X}_{1}+\dots+\gamma_{n}\mathbf{X}_{n}}{\gamma_{1}+\dots+\gamma_{n}}$ w.r.t.
-some positive weight sequence $\{\gamma_{n}\}$.
-Theorem 5.5 (Uniformly Randomized Stopped Matrix Bernstein/Hoeffding Inequality).
-Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$, $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}\preceq\mathbf{B}_{n}$ and $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ almost surely for predictable $\{\mathbf{B}_{n}\}$ and $\{\mathbf{V}_{n}\}$.
-Then, for any stopping time $\tau$, predictable positive sequence $\{\gamma_{n}\}$, $\alpha\in(0,1)$, and super-uniform random scalar $U$ independent of $\mathcal{F}_{\infty}$,
-$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{6}\lambda_{\max}\left(\sum{\gamma_{i}^{2}}(\mathbf{B}_{i}+2\mathbf{V}_{i})\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)$$\leqslant$ (77)
-$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\frac{1}{2}\lambda_{\max}\left(\sum_{i=1}^{\tau}\gamma_{i}^{2}\mathbf{B}_{i}\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)$ $\leqslant\alpha.$ (78)
-We now remark at length on this theorem.
-First, the Hoeffding term $\frac{1}{2}\lambda_{\max}\left(\sum\gamma_{i}^{2}\mathbf{B}_{i}\right)$ comes from applying $\mathbf{V}_{i}\preceq\mathbf{B}_{i}$ to the Bernstein term $\frac{1}{6}\lambda_{\max}\left(\sum\gamma_{i}^{2}(\mathbf{B}_{i}^{2}+2\mathbf{V}_{i})\right)$, among which one may choose, as is with scalar Bernstein versus Hoeffding, depending on whether variance bounds $\{\mathbf{V}_{n}\}$ better than the total square bounds $\{\mathbf{B}_{n}\}$ are known.
-Second, the strength of this bound can be seen from the following fact: taking $\{\mathbf{M}_{n}\}$ to be a constant $\mathbf{M}$, $\{\mathbf{B}_{n}\}$ a constant $\mathbf{B}$, $\{\mathbf{V}_{n}\}$ a constant $\mathbf{V}$, $\tau$ a deterministic sample size $n$ and letting $U=1$, $\gamma_{1}=\dots=\gamma_{n}=\sqrt{\frac{6\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B}+2\mathbf{V})}}$, (77) reads
-$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\right)\geqslant\sqrt{\frac{2\log(d/\alpha)\lambda_{\max}(\mathbf{B}+2\mathbf{V})}{3n}}\right)\leqslant\alpha,$ (79)
-where $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$.
-This even improves the optimal matrix Hoeffding inequality by Mackey et al.
-[2014, Corollary 4.2] whose dependence on $\mathbf{B}$ and $\mathbf{V}$ is $\frac{1}{2}\lambda_{\max}(\mathbf{B}+\mathbf{V})$ compared to our smaller $\frac{1}{3}\lambda_{\max}(\mathbf{B}+2\mathbf{V})$, [^4] The optimality of Mackey et al.
-[2014, Corollary 4.2] is in terms of the overall multiplicative constant which (79) achieves as well, not the distribution of the constant between $\mathbf{B}$ and $\mathbf{V}$.
-Even tighter dependence in this regard shall be presented soon in Section 5.3.
-and recovers (MHI), both via $\mathbf{V}\preceq\mathbf{B}$.
-The uncompromised inequality (77) further allows for stopping time-validity, time-varying and adaptive weighting, and randomized tightening simultaneously.
-Third, in the i.i.d.
-(or homoscedastic more generally) sequential (sample size $\tau$ not fixed) case, following Waudby-Smith and Ramdas [2020], Wang and Ramdas [2023a], the weights $\{\gamma_{n}\}$ can be set as a vanishing sequence such that each $\gamma_{n}$ matches the fixed-time optimal choice of $\gamma_{n}$ with sample size $n$, so we recommend $\gamma_{n}=\sqrt{\frac{6\log(d/\alpha)}{n\lambda_{\max}(\mathbf{B}+2\mathbf{V})}}$.
-By doing so, the concentration tightens itself as time goes on; or, the corresponding sequential test has asymptotic power 1 [Waudby-Smith and Ramdas, 2023, Table 1].
-Moreover, it can be seen that the closed-form concentration of $\overline{\mathbf{X}}_{\tau}$ here is achieved via multiple steps of scalarization, each step loosening the bound by a little.
-Therefore, if one does not want to quantify uncertainty but is only interested in sequentially testing if the means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}$, the more powerful test in the current subsection is still with the largest process, the self-normalized supermartingale $\{L^{\textsf{SN}}_{n}\}$ (i.e., rejecting whenever $L^{\textsf{SN}}_{n}$ surpasses $Ud/\alpha$), with variance upper bounds $\mathbf{V}_{n}$ supplied by the upper bounds $\mathbf{B}_{n}$.
-5.3 Matrix Empirical Bernstein Inequality
-We now present a closed-form empirical Bernstein inequality that enjoys the same stopping time-validity, adaptive weighting, and randomization as well, foreshadowed by a matrix expectation bound in Proposition 4.10.
-Empirical Bernstein inequalities for scalars differ from “classical” Bernstein inequalities, e.g.
-our (77) when $d=1$, in that the population variance is replaced by an empirical variance in the deviation bound.
-They therefore achieve the advantage of Bernstein’s inequalities over Hoeffding’s inequalities without requiring knowing the variance beforehand.
-Scalar empirical Bernstein inequalities are often derived in two different ways.
-First, a union bound between a classical Bernstein inequality and a concentration inequality on the empirical variance, a method that early empirical Bernstein results [Audibert et al., 2009, Maurer and Pontil, 2009] all share.
-Second, a self-normalization technique pioneered by Howard et al.
-[2021, Theorem 4] and later followed by Waudby-Smith and Ramdas [2023], Chugg et al.
-[2023], among others, which enables supermartingale-based, stopping time-valid concentration, and variance proxy by predictable estimates other than the empirical variance.
-These two methods are inherently different and as argued convincingly by Howard et al.
-[2021, Appendix A.8]: the latter’s avoidance of the union bound produces a sharper concentration.
-No matrix empirical Bernstein results, especially those in the second fashion, are previously known.
-Our matrix extension of Howard et al.
-[2021, Theorem 4] starts from the bound we realier prove as Proposition 4.10.
-Recall that $\psi_{E}(x)=-\log(1-x)-x=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$ for $x\in[0,1)$, and that
-$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\middle|\mathcal{F}_{n-1}\right)\preceq\exp(\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n}))$ (80)
-under, among other, a key assumption that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$.
-We can use Lemma 5.1 again to produce a scalar supermartingale and a closed-form spectral concentration inequality.
-Theorem 5.6 (Uniformly Randomized Stopped Matrix Empirical Bernstein Inequality).
-Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$.
-Let $\{\widehat{\mathbf{X}}_{n}\}$ be a sequence of predictable and integrable $\mathcal{S}_{d}$-valued random matrices such that $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ almost surely.
-Then, for any predictable $(0,1)$-valued sequence $\{\gamma_{n}\}$,
-$L_{n}^{\mathsf{EB}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}\right)$ (81)
-is a supermartingale. Further, for any stopping time $\tau$, $\alpha\in(0,1)$, and super-uniform scalar random variable $U$ independent of $\mathcal{F}_{\infty}$, we have
-$\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{\tau}^{\gamma}-\overline{\mathbf{M}}_{\tau}^{\gamma}\right)\geqslant\frac{\log(dU/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{\tau}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}\right)}{\gamma_{1}+\dots+\gamma_{\tau}}\right)\leqslant\alpha.$ (82)
-Before we remark on the uncompromised Theorem 5.6, let us first write down its fixed-time, non-randomized, fine-tuned special case with $\tau=n$ and $U=1$ which shall justify the “empirical Bernstein” name it bears.
-Corollary 5.7 (Matrix Empirical Bernstein Inequality).
-Suppose $\alpha\in(0,1)$.
-Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be $\mathcal{S}_{d}^{[0,1]}$-valued i.i.d.
-random matrices with mean $\mathbf{M}=\mathbb{E}\mathbf{X}_{1}$.
-Let $\overline{\mathbf{X}}_{i}=\frac{1}{i}(\mathbf{X}_{1}+\dots+\mathbf{X}_{i})$ and $\overline{\mathbf{X}}_{0}=0$.
-Then, for any predictable $(0,1)$-valued sequence $\gamma_{1},\dots,\gamma_{n}$,
-[…]
-Further, if we define the following variance proxies
-$\overline{\mathbf{V}}_{0}=0,\quad\overline{\mathbf{V}}_{k}=\frac{1}{k}\sum_{i=1}^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{k})^{2},\quad\overline{v}_{k}=\lambda_{\max}(\overline{\mathbf{V}}_{k})\vee\frac{5\log(d/\alpha)}{n},$ (84)
-$\widetilde{s}_{n}=\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}^{2})}{\overline{v}_{i-1}}\right),$ (85)
-and set $\gamma_{i}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}$ for $i=1,\dots,n$, then,
-$D_{n}^{\mathsf{EB}}\leqslant\sqrt{\frac{\log(d/\alpha)}{2n}}\frac{1+2\widetilde{s}_{n}}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}\quad\approx\sqrt{\frac{9\log(d/\alpha)\lambda_{\max}(\mathbf{V})}{2n}}$ (86)
-where $\mathbf{V}=\mathbf{Var}(\mathbf{X}_{1})$; further,
-$\lim_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}=\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}\quad\text{almost surely}.$ (87)
-We prove both Theorem 5.6 and Corollary 5.7 in Section B.4.
-The asymptotic behavior (87) of deviation bound $D_{n}^{\mathsf{EB}}$ is striking as it adapts fully to, without knowing, the true variance $\mathbf{V}$, a strict improvement even compared to the deviation bound $D_{n}^{\mathsf{B}}$ of the our previous matrix Bernstein inequality (79),
-$\sqrt{n}D_{n}^{\mathsf{B}}=\sqrt{2\log(d/\alpha){\frac{1+2\lambda_{\max}(\mathbf{V})}{3}}},$ (88)
-which requires knowing beforehand, but only adapts partially to, the variance $\mathbf{V}$.
-Indeed, when $\mathbf{V}$ is small, (88) can be arbitrarily larger than (87).
-We can also compare Corollary 5.7 to another matrix Bernstein inequality in the literature by Tropp [2012, Theorem 1.4], who proves that $\mathbb{P}\left(\lambda_{\max}\left(\overline{\mathbf{X}}_{n}-\mathbf{M}\right)\geqslant D_{n}^{\mathsf{TB}}\right)\leqslant\alpha$ under the same assumptions (plus knowing the variance $\mathbf{V}$), where its deviation bound
-$D_{n}^{\mathsf{TB}}=\frac{\log(d/\alpha)+\sqrt{\log^{2}(d/\alpha)+18\log(d/\alpha)n\lambda_{\max}(\mathbf{V})}}{3n}$ (89)
-satisfies the same asymptotic $\lim_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{TB}}=\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}$ as (87), which we achieve without requiring knowing $\mathbf{V}$.
-Other matrix Bernstein bounds in the literature, like the ones by Mackey et al.
-[2014, Corollary 5.2] and by Oliveira [2009], are looser than the Tropp [2012, Theorem 1.4] bound we discuss here as acknowledged by Mackey et al.
-[2014] themselves.
-Our Theorem 5.6 and Corollary 5.7 owe much to the techniques developed by Waudby-Smith and Ramdas [2023, Theorem 2 and Remark 1] in the scalar case (who in turn build on the earlier result by Howard et al.
-[2021, Theorem 4] via the “predictable mixing” sequence $\{\gamma_{n}\}$).
-We recover exactly as special cases those two known results when $d=1$.
-We reiterate our remark following Proposition 4.10 that $\widehat{\mathbf{X}}_{n}$ in Theorem 5.6 can usually be taken as the empirical average from $\mathbf{X}_{1}$ to $\mathbf{X}_{n-1}$, which is exactly what is done in Corollary 5.7.
-In the i.i.d.
-(or homoscedastic more generally) sequential case without the sample size $n$ fixed in advance, we can again (recalling a similar discussion following Theorem 5.5) set the weights $\{\gamma_{n}\}$ to be a vanishing sequence such that each $\gamma_{n}$ matches the fixed-time optimal choice of $\gamma_{n}$ with sample size $n$, so we recommend $\gamma_{n}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{n-1}}}$.
-Under this weight sequence, we see the choice of a weighted average $\widehat{\mathbf{X}}_{n}=\overline{\mathbf{X}}_{n-1}^{\psi_{E}(\gamma)}$ is more reasonable as it roughly minimizes the weighted sum of squares $\sum_{i=1}^{\tau}\psi_{E}(\gamma_{i})(\mathbf{X}_{i}-\widehat{\mathbf{X}}_{i})^{2}$ in (82) into a weighted empirical variance — a heuristic which should work for Proposition 4.10 as well.
-Obviously, as long as $\widehat{\mathbf{X}}_{n}$ is any average, weighted or not, of $\mathbf{X}_{1},\dots,\mathbf{X}_{n-1}$, the condition $\lambda_{\min}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\geqslant-1$ is met when $\{\mathbf{X}_{n}\}$ all take values in $\mathcal{S}_{d}^{[0,1]}$.
+First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Theorem 3.3.
+We apply Theorem 5.3 to obtain the first inequality.
+For the second inequality, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$ and apply (75) to the stopping time $\nu\wedge\tau$.
+∎
+These two inequalities above, we note, generalize respectively the stopped and time-uniform statements of (MVI), just as the two inequalities by Ramdas and Manole [2024, Theorem 4.1, Corollary 4.1.1] generalize the two forms of the scalar (VI).
+It is important to note that while it is tempting to hope for the following randomized time-uniform (MVI):
+$\mathbb{P}(\exists n,\;\mathbf{Y}_{n}\npreceq U\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}),$ (77)
+the temptation must be resisted as it is known to fail in the scalar case; see the discussion before Corollary 4.1.1 by Ramdas and Manole [2024].
+Theorem 3.4 provides the following improvement over Algorithm 1 where the user may prespecify a stopping criterion $\tau$ beforehand.
+As is already stated in Algorithm 1, whenever $\mathbf{L}\mathbf{R}\npreceq\mathbf{A}$, $\mathcal{H}_{0}$ is rejected; what is now different is that when the stopping criterion $\tau$ is met and $\mathcal{H}_{0}$ is not yet rejected, the user can draw a $U\sim\operatorname{Unif}_{(0,1)}$ to see if $\mathbf{L}\mathbf{R}\npreceq U\mathbf{A}$, and reject $\mathcal{H}_{0}$ when this happens.
 6 Experiments: To Scalarize or Not to Scalarize
-We have developed thus far two categories of matrix concentration bounds: those obtained via the pure matrix approach, typically expressed in the $\preceq$ order, in Sections 3 and 4; and those obtained via scalarization, i.e.
-taking the trace, in Section 5.
-It is therefore a natural question to ask if either of these is tighter than the other.
-A fair comparison can be raised between the self-normalized matrix and scalar test supermartingales, Example 4.8 and Example 5.2 — for both are under the assumption of finite variance and based on the key one-step self-normalized bound (71).
-We place them in the setting of the following sequential testing problem: let $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ be an i.i.d.
-sequence of matrices in $\mathcal{S}_{d}$, with mean $\mathbf{M}$ and variance upper bound $\mathbf{I}$.
-We would like to sequentially test the null hypothesis $H_{0}:\mathbf{M}=0$.
-Let us call the tests derived from Example 4.8 and Example 5.2 for this null Matrix and Scalar respectively.
-Ignoring the randomizer for simplicity (plug in $\mathbf{U}=\mathbf{I}$), Matrix computes the matrix-valued process
+We now return to our discussion in Section 5.2 that from the condition
+$\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\exp(\mathbf{C}_{n}^{\prime})$ (78)
+that holds under the null, one can either use our approach to construct a matrix test supermartingale via Lemma 3.5, or use the stepwise scalarization technique formalized as Lemma 5.2 to construct a scalar test supermartingale.
+We now compare these two empirically.
+We conduct the comparison with the self-normalized Example 3.8, and consider the two corresponding sequential test algorithms, Matrix and Scalar.
+More specifically, both receive the stream of i.i.d matrices $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ in $\mathcal{S}_{d}$ with mean $\mathbf{M}$ and variance upper bound $\mathbf{I}$ and sequentially test the null hypothesis $H_{0}:\mathbf{M}=0$.
+Matrix computes the matrix-valued process
 $\mathbf{Y}_{n}^{\textsf{SN}}=\exp\left(-\frac{\gamma_{1}^{2}+\dots+\gamma_{n}^{2}}{3}\right)\cdot\exp\left(\frac{\gamma_{1}\mathbf{X}_{1}}{2}-\frac{\gamma_{1}^{2}\mathbf{X}_{n}^{2}}{12}\right)\dots\exp\left(\frac{\gamma_{n}\mathbf{X}_{n}}{2}-\frac{\gamma_{n}^{2}\mathbf{X}_{n}^{2}}{12}\right)\\ \exp\left(\frac{\gamma_{n}\mathbf{X}_{n}}{2}-\frac{\gamma_{n}^{2}\mathbf{X}_{n}^{2}}{12}\right)\dots\exp\left(\frac{\gamma_{1}\mathbf{X}_{1}}{2}-\frac{\gamma_{1}^{2}\mathbf{X}_{1}^{2}}{12}\right),$ (79)
-rejecting when $\mathbf{Y}_{n}^{\textsf{SN}}\npreceq\mathbf{A}$ for some threshold matrix $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$; Scalar computes the scalar-valued process
+rejecting when $\mathbf{Y}_{n}^{\textsf{SN}}\npreceq\mathbf{A}$ for some threshold matrix $\mathbf{A}\in\mathcal{S}_{d}^{++}$ such that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$ picked beforehand; Scalar computes the scalar-valued process
 $L_{n}^{\textsf{SN}}=\exp\left(-\frac{\gamma_{1}^{2}+\dots+\gamma_{n}^{2}}{3}\right)\cdot\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}\mathbf{X}_{i}-\sum_{i=1}^{n}\frac{\gamma_{i}^{2}\mathbf{X}_{i}^{2}}{6}\right),$ (80)
@@
 When only $n=1$ observation is revealed, $L_{1}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{1}^{\textsf{SN}})$.
-We shall see that the comparison depends heavily on the rejection threshold matrix $\mathbf{A}$ that Matrix picks beforehand.
+We shall see that the comparison depends heavily on the rejection threshold matrix $\mathbf{A}$ that Matrix picks beforehand, as rejection happens on the complements of the “concentration regions” discussed in Section 5.1.
+The following are direct corollaries of Proposition 5.1.
 Proposition 6.1.
@@
 $\mathbf{A}=\varepsilon^{-1}(d-1)\mathbf{P}_{1}+\dots\varepsilon^{-1}(d-1)\mathbf{P}_{d-1}+(\alpha-\varepsilon)^{-1}\mathbf{P}_{d}$ (83)
-with $\varepsilon\in(0,\alpha-\lambda_{d}^{-1})$. The condition $\lambda_{d}>1/\alpha$ here is strictly weaker than the condition under which Scalar rejects: $\lambda_{1}+\dots+\lambda_{d}=d/\alpha$.
-Indeed, as we can see from these examples with $n=1$ (thus we have only compared the matrix Markov inequality with the scalar Markov inequality on the trace), for Matrix to outpower Scalar, apart from prior knowledge, it is favorable that $\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})$ is an outlier in the spectrum of $\mathbf{Y}_{1}^{\mathsf{SN}}$, i.e., ${\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{1}^{\mathsf{SN}})}$ is close to 1.
-This is exactly what we would predict from the discussion on (UMMI) early on, where we mention that (UMMI) is tighter when the random matrix has (approximately) a lower rank.
-In the very next subsection, we shall see that whereas the condition ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\approx 1$ seems artificial and contrived here with $n=1$, it becomes an almost certain phenomenon with large $n$.
+with $\varepsilon\in(0,\alpha-\lambda_{d}^{-1})$.
+The condition $\lambda_{d}>1/\alpha$ here is strictly weaker than the condition under which Scalar rejects: $\lambda_{1}+\dots+\lambda_{d}\geqslant d/\alpha$.
+Indeed, as we can see from these examples with $n=1$ (thus we have only compared the matrix Markov inequality with the scalar Markov inequality on the trace), for Matrix to outpower Scalar, apart from prior knowledge, it is favorable that $\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})$ is an outlier in the spectrum of $\mathbf{Y}_{1}^{\mathsf{SN}}$, i.e., its effective rank $\operatorname{r}(\mathbf{Y}_{1}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{1}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{1}^{\mathsf{SN}})}$ is close to 1.
+This is exactly what we would predict from the discussion in Section 5.1, where we mention that (MMI) is tighter when the rejection threshold matrix $\mathbf{A}$ (thus also the random matrix itself) has a lower effective rank.
+In the very next subsection, we shall see that whereas the low-rank condition $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})\approx 1$ seems artificial and contrived here with $n=1$, it becomes an almost certain phenomenon with large $n$.
 6.2 Comparison with a Sequence of Observations
@@
 The scalar component process $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows exponentially fast if $\pi_{k}$ has mean $\lambda_{k}\neq 0$, and shrinks to 0 if $\lambda_{k}=0$; moreover, the rate in which $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows depends exponentially on the size $|\lambda_{k}|$.
-Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will be approximately low-rank with ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})\mathbf{P}_{d}$ will exponentially outgrow the rest.
+Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will have a low effective rank with $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})\mathbf{P}_{d}$ will exponentially outgrow the rest.
 In this case if Matrix correctly anticipates the largest component $\mathbf{P}_{d}$, by picking again the threshold matrix
@@
 We consider the following three sequential tests: smart Matrix, one that follows the recipe (88) with $\varepsilon=\alpha/20$; naive Matrix, one that is agnostic about the real distribution and picks $\mathbf{A}=(d/\alpha)\mathbf{I}$; and Scalar, all with the parameters $\gamma_{n}=n^{-1/2}$ and $\alpha=0.05$.
-We record the times at which these tests correctly reject the null, as well as the evolution of ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$.
-The findings are displayed in Figure 1: smart Matrix indeed outpowers the other two, whereas naive Matrix and Scalar are similar in power, explained by the quick convergence of ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$ to 1.
+We record the times at which these tests correctly reject the null, as well as the evolution of $\operatorname{r}^{-1}(\mathbf{Y}_{n}^{\mathsf{SN}})={\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}$.
+The findings are displayed in Figure 1: smart Matrix indeed outpowers the other two, whereas naive Matrix and Scalar are similar in power, explained by the quick convergence of $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})$ to 1.
 Figure 1: Empirical findings with commuting i.i.d.
@@
 Right.
-Evolution of the spectral distribution of the matrix-valued test process under the alternative.
+Evolution of the inverse of the effective rank of the matrix-valued test process under the alternative.
 While this indicates that
@@
 Besides the unsurprising fact that the smart Matrix is again the most powerful, rejecting the null earlier than the other two, we also observe the phenomenon that, while in the $n=1$ case, the matrix-valued test process with “naive” $\mathbf{A}=(d/\alpha)\mathbf{I}$ is always less powerful than the scalar-valued test process, this is reversed under non-commutativity with large $n$ as naive Matrix rejects slightly earlier than Scalar in our simulation.
-As also shown in Figure 2, the combination of two auxiliary findings contributes to this phenomenon: ${\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$ even when observations do not commute; and, while $L_{n}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ in all our previous examples, now $\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ grows slightly faster than $L_{n}^{\textsf{SN}}$.
+As also shown in Figure 2, the combination of two auxiliary findings contributes to this phenomenon: $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$ even when observations do not commute; and, while $L_{n}^{\textsf{SN}}=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ in all our previous examples, now $\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}})$ grows slightly faster than $L_{n}^{\textsf{SN}}$.
 It takes further theoretical investigations to understand why this happens (either in this particular example or in general), but to some extent it hints nonetheless at the possibility that in practice where the effect size is small (requiring large $n$) and matrix observations are non-commuting, there is an inherent advantage using the matrix-valued test process we propose.
@@
 Middle.
-Evolution of the spectral distribution of the matrix-valued test process under the alternative.
+Evolution of the inverse of the effective rank of the matrix-valued test process under the alternative.
 Right.
@@
 7 Conclusion
-We prove in this paper several novel concentration inequalities for random symmetric matrices.
-Some of these bounds are tighter than previous ones in the literature via randomization or Loewner order, others enjoy the virtues of validity at stopping times and/or the more general applicability to non-i.i.d.
-exchangeable matrices.
-These results shall find applications in, for example, covariance estimation and testing.
-We also leave open some questions for future studies to address.
-These include a deeper understanding of the matrix randomizer $\mathbf{U}$ in our paper, which belongs to the “trace super-uniform” distribution family; tighter $p$-Chebyshev inequalities; further investigation into matrix-valued e-processes; as well as the theoretical explanation of our comparison between matrix-valued test supermartingales and scalar-valued test supermartingales when testing sequentially the same null hypotheses.
-Our experiments show that matrix-valued test processes are often more powerful in these sequential tests.
+In this paper, we study positive semidefinite matrix supermartingales and backward submartingales.
+In particular, we present several maximal inequalities that generalize existing results in the well-known scalar case.
+We demonstrate the statistical applications of these results by noting that for data and hypotheses that inherently take matrix values, our “matrix-native” approach is beneficial and can outperform previous methods that often require “scalarization”.
+In particular, sequential testing with matrix-valued test supermartingales and e-processes allows a freely chosen rejection threshold matrix $\mathbf{A}$ subject to the constraint $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$, one that can be chosen according to the prior knowledge to boost power without sacrificing the type 1 error guarantee.
+On the other hand, taking the threshold matrix $\mathbf{A}$ to be a multiple of the identity matrix, a degenerate choice that it appears, can still lead to a more powerful test in the large sample regime compared to scalarization.
+This is an empirical finding of our paper that calls for further understanding.
+Our study shall lead to more powerful and flexible sequential hypothesis testing procedures for matrix inference problems including covariance and graph testing.
 Acknowledgement
-The authors thank Joel A. Tropp, Lester Mackey, and Tudor Manole for suggestions and discussions, and acknowledge support from NSF grant DMS-2310718.
-Appendix A $p$-Chebyshev Inequalities
-A.1 The Scalar (and Vector) $p$-Chebyshev Inequality
-The following statement shows that for i.i.d. random scalars, the central $p$th moment of the sum grows linearly in the number of scalars just like the variance.
-Proposition A.1 (Contraction of the scalar $p$th moment, Lemma 4.2 of Cherapanamjeri et al. [2022]).
-Suppose $1\leqslant p\leqslant 2$.
-Let $X_{1},\dots,X_{n}$ be i.i.d.
-scalar with mean $\mu=\mathbb{E}X_{1}$ and central $p$th moment $v_{p}=\mathbb{E}|X_{1}-\mu|^{p}$.
-Then, $\mathbb{E}|X_{1}+\dots+X_{n}-n\mu|^{p}\leqslant 2^{2-p}nv_{p}$.
-Wang et al. [2021] generalize the statement to vectors under the martingale difference assumption.
-Proposition A.2 (Concentration of the vector $p$th moment, Lemma 7 of Wang et al. [2021]).
-Suppose $1\leqslant p\leqslant 2$.
-Let $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ be $d$-dimensional random vectors adapted to a filtration $\mathcal{F}_{0},\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ with conditional mean $\mu=\mathbb{E}(\mathbf{x}_{i}|\mathcal{F}_{i-1})$ and conditional central $p$th moment upper bound $v_{p}\geqslant\mathbb{E}(\|\mathbf{x}_{i}-\mu\|^{p}|\mathcal{F}_{i-1})$.
-Then, $\mathbb{E}\|\mathbf{x}_{1}+\dots+\mathbf{x}_{n}-n\mu\|^{p}\leqslant 2^{2-p}d^{1-p/2}nv_{p}$.
+The authors thank Joel A. Tropp, Lester Mackey, Tudor Manole, and Reihaneh Malekian for suggestions and discussions, and acknowledge support from NSF grant DMS-2310718.
+Appendix A On the Randomized Matrix Markov Inequality
+Instead of randomizing the bound with the isotropic diagonal matrix $U\mathbf{I}$ where $U$ is a uniform or super-uniform random scalar, let us seek an inherently matrix-born definition of uniformity and super-uniformity.
+Scalar super-uniformity has two viable extensions to matrices, one in terms of the trace and the other of the spectral norm.
+Definition A.1.
+We call an $\mathcal{S}_{d}^{+}$-valued random matrix $\mathbf{U}$ spectrally super-uniform if $\mathbb{P}(\mathbf{U}\nsucceq\mathbf{Y})\leqslant\lambda_{\max}(\mathbf{Y})$ for all $\mathbf{Y}\in\mathcal{S}_{d}^{+}$, and $\mathcal{Y}$-spectrally uniform if, further, equality holds for all $\mathbf{Y}\in\mathcal{Y}\subseteq\mathcal{S}_{d}^{+}$.
+We say $\mathbf{U}$ is trace super-uniform if $\mathbb{P}(\mathbf{U}\nsucceq\mathbf{Y})\leqslant\mathrm{tr}\mathbf{Y}$ for all $\mathbf{Y}\in\mathcal{S}_{d}^{+}$, and $\mathcal{Y}$-trace uniform if, further, equality holds for all $\mathbf{Y}\in\mathcal{Y}\subseteq\mathcal{S}_{d}^{+}$.
+Note that every spectrally super-uniform matrix is a trace super-uniform matrix since $\lambda_{\max}(\mathbf{Y})\leqslant\mathrm{tr}\mathbf{Y}$.
+If $U$ is a super-uniform random scalar, then $U\mathbf{I}$ is a spectrally (hence trace) super-uniform random matrix, because $\mathbb{P}(U\mathbf{I}\nsucceq\mathbf{Y})=\mathbb{P}(U\leqslant\lambda_{\max}(\mathbf{Y}))\leqslant\lambda_{\max}(\mathbf{Y}).$
+Straightforwardly the identity matrix $\mathbf{I}$ is spectrally (hence trace) super-uniform.
+If $U\sim\operatorname{Unif}_{(0,1)}$, $U\mathbf{I}$ is $\mathcal{S}_{d}^{[0,1]}$-spectrally uniform, and $\left\{\mathbf{v}\mathbf{v}^{\mathsf{T}}:\|\mathbf{v}\|\leqslant 1\right\}$-trace uniform.
+Having defined trace super-uniformity, we can now state a matrix extension of the uniformly randomized Markov’s inequality (72).
+Theorem A.2 (Generalized Uniformly Randomized Matrix Markov Inequality).
+Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{X}\subseteq\mathcal{S}_{d}^{+}$, and $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$.
+For any $\mathbf{A}\in\mathcal{S}_{d}^{++}$,
+$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1}).$ (92)
+The above holds with equality if and only if $\mathbf{U}$ is $\mathbf{A}^{-1/2}\mathcal{X}\mathbf{A}^{-1/2}$-trace uniform.
+Proof.
+By our assumptions,
+$\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2})=\mathbb{P}(\mathbf{A}^{-1/2}\mathbf{X}\mathbf{A}^{-1/2}\npreceq\mathbf{U})=\mathbb{E}(\mathbb{P}(\mathbf{A}^{-1/2}\mathbf{X}\mathbf{A}^{-1/2}\npreceq\mathbf{U}|\mathbf{X}))$ (93)
 […]
-Corollary A.3 (Vector $p$-Chebyshev Inequality).
-Suppose $\mathbf{x}_{1},\dots,\mathbf{x}_{n}$ satisfy the same assumptions as in Proposition A.2. Then, for any $a>0$,
-$\mathbb{P}\left(\left\|\overline{\mathbf{x}}_{n}-\mu\right\|\geqslant a\right)\leqslant\frac{2^{2-p}d^{1-p}v_{p}}{n^{p-1}a^{p}}.$ (103)
-A.2 Matrix $p$-Chebyshev Inequalities
-The following conjecture is an ideal extension of Propositions A.2 and A.3.
-Conjecture A.4 (Matrix $p$-Chebyshev Inequality).
-Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d.
-random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and $p$th central moment matrix $\mathbf{V}_{p}:=\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}-\mathbf{M}))^{p}$.
-Then, for any $\mathbf{A}\in\mathcal{S}_{d}^{++}$, there exists a function $f:[1,2]\times\mathbb{N}\to(0,\infty)$ that grows sublinearly in its second argument, and bounded in its first argument, such that the $p$th central moment contracts,
-$\mathbb{E}(\operatorname{abs}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n}-n\mathbf{M})^{p})\preceq n\cdot f(p,d)\cdot\mathbf{V}_{p};$ (104)
-and consequently, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$,
-$\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\mathbf{A})\leqslant n^{1-p}\cdot f(p,d)\cdot\mathrm{tr}(\mathbf{V}_{p}\mathbf{A}^{-p}).$ (105)
-If it holds, the $\mathbf{A}$ above can be randomized into $(\mathbf{A}^{p/2}\mathbf{U}\mathbf{A}^{p/2})^{1/p}$ as well.
-We do have the following heavily scalarized and dimension-dependent bound that scales correctly w.r.t.
-the sample size $n$.
-We find this bound trivial as it regards the matrix as a mere collection of $d^{2}$ scalars and is strictly looser than the “standard” vector bound (Proposition A.2) applied to the flattened matrix.
-Proposition A.5 (Matrix $p$-Chebyshev Spectral Inequality).
-Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be i.i.d.
-random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{1}=\mathbf{M}$ and $p$th spectral central moment $v_{p}:=\mathbb{E}\|\mathbf{X}_{1}-\mathbf{M}\|^{p}$.
+concluding the proof. ∎
+When taking $\mathbf{U}$ to be the deterministic $\mathbf{I}$, (92) recovers (MMI).
+It is not hard to see that (92) is strictly more general than the randomized scalar Markov’s inequality (72) and the scalar-randomized matrix Markov’s inequality (73).
+We also see from the monotonicity of the trace and the trace-spectral norm inequality that scalarized trace and spectral bounds are easily obtainable via
+$\mathbb{P}(\lambda_{\max}(\mathbf{X})>\mathrm{tr}(\mathbf{A}\mathbf{U}))\leqslant\mathbb{P}(\mathrm{tr}\mathbf{X}>\mathrm{tr}(\mathbf{A}\mathbf{U}))\leqslant\mathbb{P}(\mathbf{X}\npreceq\mathbf{A}^{1/2}\mathbf{U}\mathbf{A}^{1/2}),$ (95)
+subject to the same upper bound $\mathrm{tr}((\mathbb{E}\mathbf{X})\mathbf{A}^{-1})$.
+The following is an example where the inequality of (92) holds with equality.
+We take $\mathbf{U}=U\mathbf{I}$ where $U\sim\operatorname{Unif}_{(0,1)}$, and $\mathbf{A}\in\mathcal{S}_{d}^{++}$.
+Let $\mathbf{x}$ be a random vector in the ellipsoid $\{\mathbf{x}:\mathbf{x}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{x}\leqslant 1\}\subseteq\mathbb{R}^{d}$.
 Then,
-$\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}-n\mathbf{M}\|^{p}\leqslant 2^{2-p}d^{2-p/2}\cdot n\cdot v_{p}.$ (106)
-Proof.
-Without loss of generality suppose $\mathbf{M}=0$.
-Denote by $\|\cdot\|_{L_{p}}$ the entrywise matrix $p$-norm and $\|\cdot\|_{\mathrm{F}}=\|\cdot\|_{L_{2}}$ the Frobenius norm.
-Recall that $\|\mathbf{A}\|\leqslant\|\mathbf{A}\|_{\mathrm{F}}\leqslant\sqrt{d}\|\mathbf{A}\|$ and $\|\mathbf{A}\|_{\mathrm{F}}\leqslant\|\mathbf{A}\|_{L_{p}}\leqslant d^{\frac{2}{p}-1}\|\mathbf{A}\|_{\mathrm{F}}$.
-Then, $\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}\|^{p}$ is upper bounded by
-$\mathbb{E}\|\mathbf{X}_{1}+\dots+\mathbf{X}_{n}\|_{L_{p}}^{p}=\mathbb{E}\left[\sum_{1\leqslant j,k\leqslant d}\left|\sum_{i=1}^{n}X_{i(jk)}\right|^{p}\right]=\sum_{1\leqslant j,k\leqslant d}\mathbb{E}\left|\sum_{i=1}^{n}X_{i(jk)}\right|^{p}$ (107)
-$\stackrel{{\scriptstyle(*)}}{{\leqslant}}\,$ $\sum_{1\leqslant j,k\leqslant d}2^{2-p}n\mathbb{E}\left|X_{1(jk)}\right|^{p}=2^{2-p}n\mathbb{E}\|\mathbf{X}_{1}\|_{L_{p}}^{p}\leqslant 2^{2-p}d^{2-p}\cdot n\cdot\mathbb{E}\|\mathbf{X}_{1}\|^{p}_{\mathrm{F}}$ (108)
-[…]
-where the inequality (*)(*)( * ) comes from the contraction of scalar $p$th moment (Proposition A.1). ∎
-Appendix B Additional Lemmas and Omitted Proofs
-B.1 Proofs of Statements in Table 1
-Proof.
-For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a direct Taylor expansion argument.
-Since the matrix log is monotone (Fact 2), $\log\mathbf{G}(\gamma/n)\preceq\gamma^{2}\mathbf{C}^{2}/2n^{2}$.
-Finally, invoking the monotonicity of $\mathbf{A}\mapsto\mathrm{tr}\exp\mathbf{A}$ (Fact 3) we see that $\mathrm{tr}(\mathrm{e}^{n\log\mathbf{G}(\gamma/n)})\leqslant\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$.
-Bennett I follows from a similar reasoning applied to the Bennett MGF bound [Tropp, 2012, Lemma 6.7].
-For Bennett II, we can apply the variance bound at last: $\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{Var}(\mathbf{X}_{1})})\leqslant\mathrm{tr}(\mathrm{e}^{n(\mathrm{e}^{\gamma/n}-\gamma/n-1)\mathbf{V}})$.
-The uni-Gaussian MGF bound is also proved by Tropp [2015, Lemma 4.6.2] via a direct Taylor expansion argument.
-For symmetric Hoeffding, introducing an independent $R\sim\operatorname{rad}_{1/2}$,
-$\log\mathbf{G}(\gamma)=\log\mathbb{E}\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbf{M})}=\log\mathbb{E}\mathrm{e}^{\gamma R(\mathbf{X}_{1}-\mathbf{M})}\preceq\gamma^{2}\mathbf{B}/2$ (110)
-due to [Tropp, 2012, Lemma 7.7], leading to the corresponding upper bound on $\mathrm{tr}(\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2n})$ as before. ∎
-B.2 Dominated Integrability Lemma
+$\mathbb{P}(\|\mathbf{A}^{-1/2}\mathbf{x}\|^{2}\leqslant U)=\mathbb{P}(\mathbf{x}\mathbf{x}^{\mathsf{T}}\npreceq U\mathbf{A})=\mathrm{tr}(\mathbf{A}^{-1}\mathbb{E}(\mathbf{x}\mathbf{x}^{\mathsf{T}}))=\mathbb{E}\|\mathbf{A}^{-1/2}\mathbf{x}\|^{2},$ (96)
+where the second equality above is (92) applied on the random matrix $\mathbf{x}\mathbf{x}^{\mathsf{T}}$.
+It is not hard to see here the condition for equality is met.
+Intuitively, this hints that (92) is tighter when the random matrix $\mathbf{X}$ is approximately of lower rank, a point that we shall return to later in this paper with statistical examples.
+Concluding this section, we remark on trace super-uniformity that clearly $U\mathbf{I}+\mathbf{Y}$ is trace super-uniform when $U\sim\operatorname{Unif}_{(0,1)}$ and $\mathbf{Y}\succeq 0$, and this provides an interesting class for our purposes.
+It may be interesting to fully characterize the set of all trace super-uniform matrices, but we leave it as future work.
+Appendix B Omitted Proofs
+B.1 Dominated Integrability Lemma
 It is well-known that if $X$ and $Y$ are scalar random variables such that $c\leqslant X\leqslant Y$ almost surely for some constant $c$ and that $\mathbb{E}|Y|<\infty$, it follows that $\mathbb{E}|X|<\infty$ as well, and $\mathbb{E}X\leqslant\mathbb{E}Y$.
@@
 ∎
-B.3 Proofs of Self-Normalization Statements
+B.2 Proofs for Forward Supermartingales
 The following transfer rule [Tropp, 2012, Equation 2.2] shall commonly be used: suppose $I\subseteq\mathbb{R}$ and $f,g:I\to\mathbb{R}$ satisfies $f(x)\leqslant g(x)$, then, since the function $g-f:I\to\mathbb{R}_{\geqslant 0}$ identifies canonically with a matrix-to-matrix function $g-f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{+}$, it follows that $f(\mathbf{X})\preceq g(\mathbf{X})$ for any $\mathbf{X}\in\mathcal{S}_{d}^{I}$.
+Proof of Lemma 3.1.
+Note that
+$\{\mathbf{Y}_{n}\}$ is a supermartingale (99)
+$\iff$ For all $n$, $\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq 0$ (100)
+$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}(\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1}))\mathbf{v}\geqslant 0$ (101)
+$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n-1}\mathbf{v}-\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}|\mathcal{F}_{n-1})\geqslant 0$ (102)
+[…]
 Proof of Example 3.8.
@@
 Note that the function $f(x)=\frac{\exp(x-x^{2}/2)+\exp(-x-x^{2}/2)}{2}$ satisfies $f(x)\leqslant 1$ for all $x\in\mathbb{R}$, we see from the transfer rule and Lemma B.1 that
-$\mathbb{E}\left(\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)=\mathbb{E}\left(f(\mathbf{X}_{n}-\mathbf{M}_{n})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbf{I}.$ (121)
-B.4 Proofs of Empirical Bernstein Statements
-Proof of Proposition 4.10.
-Recall that $\psi_{E}(\gamma)=-\log(1-\gamma)-\gamma$. An inequality by Fan et al. [2015] quoted by Howard et al. [2021, Appendix A.8] states that, for all $10\leqslant\gamma<$ 1 and $\xi\geqslant-1$,
-$\exp(\gamma\xi-\psi_{E}(\gamma)\xi^{2})\leqslant 1+\gamma\xi.$ (122)
-Since $\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}\in\mathcal{S}_{d}^{[-1,\infty)}$, we can apply the transfer rule mentioned in Section B.3, replacing the scalar $\xi$ above by the matrix $\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}$, and plugging in $\gamma=\gamma_{n}\in(0,1)$,
-$\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\preceq 1+\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n}).$ (123)
-Lemma B.1 then guarantees the integrability of the left hand side, and that
-$\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})-\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbb{E}\left(1+\gamma_{n}(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})\middle|\mathcal{F}_{n-1}\right)$ (124)
-$=$ $1+\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})\preceq\exp(\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})),$ (125)
-where in the final step we use the transfer rule again with $1+x\leqslant\exp(x)$ for all $x\in\mathbb{R}$. This concludes the proof. ∎
+$\mathbb{E}\left(\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)=\mathbb{E}\left(f(\mathbf{X}_{n}-\mathbf{M}_{n})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbf{I}.\qed$ (112)
+Proof of Example 3.12.
+Using the transfer rule on the upper $p$-logarithmic contraction,
+$\mathbb{E}\left(\exp(\phi(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})))|\mathcal{F}_{n-1}\right)$ (113)
+$\preceq$ $\mathbb{E}\left(\mathbf{I}+\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})+(1/p)\gamma_{n}^{p}\left(\operatorname{abs}(\mathbf{X}_{n}-\mathbf{M}_{n})\right)^{p}|\mathcal{F}_{n-1}\right)$ (114)
+$\preceq$ $\mathbf{I}+\frac{\gamma_{n}^{p}}{p}\mathbf{V}_{n};$ (115)
+and on the lower $p$-logarithmic contraction,
+$\mathbb{E}\left(\exp(-\phi(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})))|\mathcal{F}_{n-1}\right)$ (116)
+$\preceq$ $\mathbb{E}\left(\mathbf{I}-\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})+(1/p)\gamma_{n}^{p}\left(\operatorname{abs}(\mathbf{X}_{n}-\mathbf{M}_{n})\right)^{p}|\mathcal{F}_{n-1}\right)$ (117)
+$\preceq$ $\mathbf{I}+\frac{\gamma_{n}^{p}}{p}\mathbf{V}_{n}.$ (118)
+This concludes the proof. ∎
+B.3 Proofs for Backward Submartingales
+Lemma B.2 (Contraction of the Matrix Variance).
+Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{i}=\mathbf{M}$ and variance matrix upper bound $\mathbf{Var}(\mathbf{X}_{i})\preceq\mathbf{V}$, such that $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$.
+Then, for $\overline{\mathbf{X}}_{n}=\frac{1}{n}(\mathbf{X}_{1}+\dots+\mathbf{X}_{n})$, we have
+$\mathbf{Var}(\overline{\mathbf{X}}_{n})\preceq n^{-1}\mathbf{V}.$ (119)
+Proof.
+Note that
+$\small\mathbf{Var}(\overline{\mathbf{X}}_{n})=n^{-2}\mathbb{E}\left(\sum_{i=1}^{n}(\mathbf{X}_{i}-\mathbf{M})\right)^{2}=n^{-2}\left(\sum_{i=1}^{n}\mathbf{Var}(\mathbf{X}_{i})+\sum_{i<j}\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}\right)\preceq n^{-1}\mathbf{V}.\qed$ (120)
 Proof of Theorem 4.7.
-We apply Lemma 5.1 with $\mathbf{Z}_{n}=\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})$, $\mathbf{C}_{n}=\gamma_{n}(\widehat{\mathbf{X}}_{n}-\mathbf{M}_{n})+\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}$, and $\mathbf{C}_{n}^{\prime}=\gamma_{n}(\mathbf{M}_{n}-\widehat{\mathbf{X}}_{n})$ to see that
-$L_{n}^{\mathsf{EB}}=\mathrm{tr}\exp\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})-\sum_{i=1}^{n}\psi_{E}(\gamma_{i})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)$ (126)
-is a supermartingale, which upper bounds
-$\exp\left\{\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)\right\}.$ (127)
-Applying (UVI),
-$\mathbb{P}\left(\lambda_{\max}\left(\sum_{i=1}^{n}\gamma_{i}(\mathbf{X}_{i}-\mathbf{M}_{i})\right)-\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}(\gamma_{n})(\mathbf{X}_{n}-\widehat{\mathbf{X}}_{n})^{2}\right)\geqslant\log(dU/\alpha)\right)\leqslant\alpha,$ (128)
-and the desired result follows from rearranging. ∎
-Lemma B.2.
-Let $\psi_{E}(x)=-\log(1-x)-x$. When $0\leqslant x\leqslant\sqrt{2/5}$, $\psi_{E}(x)\leqslant x^{2}$.
-Proof.
-Let $g(x)=\psi_{E}(x)-x^{2}$. The claim follows from $g^{\prime\prime}(x)=(1-x)^{-2}-2\geqslant 0$ for $x\in[0,\sqrt{2/5}]$, and $g(0)=0$, $g(\sqrt{2/5})<0$. ∎
-Proof of Corollary 5.7.
-First, it is straightforward that $\lambda_{\max}({\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}})\geqslant-1$ for every $i=1,\dots,n$ since both $\mathbf{X}_{i}$ and $\overline{\mathbf{X}}_{i-1}$ take values in $\mathcal{S}_{d}^{[0,1]}$, so Theorem 5.6 is applicable.
-Let us prove the two claims about the deviation bound $D_{n}^{\mathsf{EB}}$ under $\gamma_{i}=\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}$.
-Recall that
-$\overline{\mathbf{V}}_{0}=0,\quad\overline{\mathbf{V}}_{k}=\frac{1}{k}\sum_{i=1}^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{k})^{2},\quad\overline{v}_{k}=\lambda_{\max}(\overline{\mathbf{V}}_{k})\vee\frac{5\log(d/\alpha)}{n},$ (129)
-$\widetilde{s}_{n}=\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1}^{2})}{\overline{v}_{i-1}}\right).$ (130)
-First, via our definition, $\overline{v}_{i}\geqslant\frac{5\log(d/\alpha)}{n}$, therefore $\gamma_{i}\leqslant\sqrt{2/5}$.
-So Lemma B.2 implies that $\psi_{E}(\gamma_{i})\leqslant\gamma_{i}^{2}$.
+Due to Fact 3,
+$\mathrm{tr}\left(\frac{1}{n}((\operatorname{abs}\mathbf{X}_{1})^{p}+\dots+(\operatorname{abs}\mathbf{X}_{n})^{p})\right)\geqslant\mathrm{tr}\left(\left(\operatorname{abs}\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}\right)^{p}\right).$ (121)
 Therefore,
-$D_{n}^{\mathsf{EB}}$ $=\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}}$ (131)
-$\leqslant\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}}$ (132)
-$=\sqrt{\frac{\log(d/\alpha)}{2n}}\frac{1+2\widetilde{s}_{n}}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}.$ (133)
-Then, let us show that $\lim_{n\to\infty}D_{n}^{\mathsf{EB}}/D_{n}^{\mathsf{B}}$ almost surely.
-Via the boundedness of $\mathbf{X}_{1}$ and the strong law of large numbers, we can see the following limits hold almost surely (cf.
-Waudby-Smith and Ramdas [2023, Lemmas 4-6]):
-$\lim_{k\to\infty}\overline{\mathbf{X}}_{k}=\mathbf{M},\quad\lim_{k\to\infty}\overline{\mathbf{V}}_{k}=\mathbf{V}=\mathbf{Var}(\mathbf{X}_{1}),$ (134)
-$\lim_{k\to\infty}\overline{v}_{k}=\lambda_{\max}(\mathbf{V}),\quad\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}=\mathbf{V},$ (135)
-$\lim_{n\to\infty}\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}\right)=\lambda_{\max}(\mathbf{V}),$ (136)
-$\lim_{n\to\infty}\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}{\overline{v}_{i-1}}\right)=1.$ (137)
-Therefore, we have, via the expansion $\psi_{E}(x)=\sum_{k=2}^{\infty}\frac{x^{k}}{k}$,
-$\limsup_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}$ (138)
-$=$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\psi_{E}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (139)
-$\leqslant$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{2}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{2}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (140)
-$+$ $\sum_{k=3}^{\infty}\limsup_{n\to\infty}\frac{\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{k}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{k}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (141)
-$=$ $\limsup_{n\to\infty}\frac{\log(d/\alpha)+\lambda_{\max}\left(\sum_{i=1}^{n}\frac{1}{2}\left(\sqrt{\frac{2\log(d/\alpha)}{n\overline{v}_{i-1}}}\right)^{2}(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}\right)}{\frac{1}{n}\sum_{i=1}^{n}\sqrt{\frac{2\log(d/\alpha)}{\overline{v}_{i-1}}}}$ (142)
-$=$ $\limsup_{n\to\infty}\sqrt{\frac{\log(d/\alpha)}{2}}\frac{\left(1+\lambda_{\max}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{(\mathbf{X}_{i}-\overline{\mathbf{X}}_{i-1})^{2}}{\overline{v}_{i-1}}\right)\right)}{\frac{1}{n}\sum_{i=1}^{n}\overline{v}_{i-1}^{-1/2}}$ (143)
-$=$ $\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}.$ (144)
-Similarly, one can show that $\liminf_{n\to\infty}\sqrt{n}D_{n}^{\mathsf{EB}}\geqslant\sqrt{2\log(d/\alpha)\lambda_{\max}(\mathbf{V})}$, concluding the proof. ∎
+$r_{n}=\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{C})^{p}\right)$ (122)
+is a scalar nonnegative backward submartingale. Using the scalar backward Ville’s inequality (7),
+$\mathbb{P}\left(\exists n,\;\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})^{p}\right)\geqslant a^{p}\right)\leqslant a^{-p}\mathbb{E}r_{1}=a^{-p}\mathrm{tr}\mathbf{V}_{p}.$ (123)
+Finally, noting that $\|\mathbf{A}\|^{p}\leqslant\mathrm{tr}((\operatorname{abs}\mathbf{A})^{p})$ concludes the proof. ∎
+B.4 Proof of Proposition 5.1
+Recall for convenience the definition of sets
+$\mathcal{C}_{\mathrm{tr}}(\alpha)=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathrm{tr}\mathbf{X}\leqslant d/\alpha\},$ (124)
+$\mathcal{C}_{\mathbf{v}}(\alpha)=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathbf{v}^{\mathsf{T}}\mathbf{X}\mathbf{v}\leqslant 1/\alpha\},$ (125)
+$\mathcal{A}_{\alpha}=\{\mathbf{A}\in\mathcal{S}_{d}^{++}:\mathrm{tr}(\mathbf{A}^{-1})=\alpha\},$ (126)
+$\mathcal{C}({\mathbf{A}})=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathbf{X}\preceq\mathbf{A}\}.$ (127)
+Proof of Statement 1.
+Let $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha)$.
+Then $\lambda_{\max}(\mathbf{X})\leqslant\mathrm{tr}\mathbf{X}\leqslant d/\alpha$.
+Therefore $\mathbf{X}\in\mathcal{C}((d/\alpha)\mathbf{I})$.
+To see that there exists an $\mathbf{X}\in\mathcal{C}((d/\alpha)\mathbf{I})$ such that $\mathbf{X}\notin\mathcal{C}_{\mathrm{tr}}(\alpha)$, one can simply take $\mathbf{X}=(d/\alpha)\mathbf{I}$.
+∎
+Proof of Statement 2.
+To see that there exists an $\mathbf{X}\in\mathcal{C}(\mathbf{A})$ such that $\mathbf{X}\notin\mathcal{C}_{\mathrm{tr}}(\alpha)$, one can take $\mathbf{X}=\mathbf{A}$ as $\mathrm{tr}(\mathbf{A})\geqslant d^{2}/\alpha>d/\alpha$.
+To see that there exists an $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha)$ such that $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$, suppose $\lambda_{\min}(\mathbf{A})=d/\alpha-2\delta$ where $\delta>0$.
+Take $\mathbf{X}$ to be a matrix with the same spectrum as $\mathbf{A}$ and eigenvalues $d/\alpha-\delta,\delta/(d-1),\dots,\delta/(d-1)$.
+∎
+Proof of Statement 3.
+First, when $\alpha^{\prime}\geqslant\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ then $\lambda_{\max}(\mathbf{X})\leqslant\mathrm{tr}(\mathbf{X})\leqslant d/\alpha^{\prime}\leqslant 1/\lambda_{\max}(\mathbf{A}^{-1})=\lambda_{\min}(\mathbf{A})$ so $\mathbf{X}\in\mathcal{C}(\mathbf{A})$.
+Second, when $\alpha^{\prime}<\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\delta=(d/\alpha^{\prime}-\lambda_{\min}(\mathbf{A}))/2$ and take $\mathbf{X}$ as a matrix with the same spectrum as $\mathbf{A}$ and eigenvalues $\lambda_{\min}(\mathbf{A})+\delta,\delta/(d-1),\dots,\delta/(d-1)$ to see that $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ but $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$.
+∎
+Proof of Statement 4.
+First, when $\alpha^{\prime}\leqslant d(\mathrm{tr}\mathbf{A})^{-1}$, let $\mathbf{X}\in\mathcal{C}(\mathbf{A})$ then $\mathrm{tr}\mathbf{X}\leqslant\mathrm{tr}\mathbf{A}\leqslant d/\alpha^{\prime}$ so $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$.
+Second, when $\alpha^{\prime}>d(\mathrm{tr}\mathbf{A})^{-1}$, take $\mathbf{X}=\mathbf{A}$ to see that $\mathbf{X}\in\mathcal{C}(\mathbf{A})$ but $\mathbf{X}\notin\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$.
+∎
+Proof of Statement 5.
+Complete the orthonormal basis $\{\mathbf{v}_{k}\}_{1\leqslant k\leqslant d}$ where $\mathbf{v}_{1}=\mathbf{v}$ and let $\mathbf{A}=\alpha^{-1}\mathbf{v}\mathbf{v}^{\mathsf{T}}+\sum_{k=2}^{d}\delta\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{T}}$ for $\delta=\frac{\alpha^{\prime}-\alpha}{d-1}$.
+∎
+Proof of Statement 6.
+Let $\mathbf{A}$, $\alpha^{\prime}$, and $\mathbf{v}$ be arbitrary.
+Complete the orthonormal basis as in the previous proof.
+To find an $\mathbf{X}\in\mathcal{C}_{\mathbf{v}}(\alpha^{\prime})$ such that $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$, take $\mathbf{X}={\alpha^{\prime}}^{-1}\mathbf{v}\mathbf{v}^{\mathsf{T}}+\sum_{k=2}^{d}L\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{T}}$ for some sufficiently large $L$.
 B.5 Proof of Lemma 5.2
 Proof.
-Due to the monotonicity of $\log$ (Fact 2), the condition (68) implies
+Due to the monotonicity of $\log$ (Fact 1), the condition (67) implies
 $\log\mathbb{E}(\exp(\mathbf{Z}_{n}-\mathbf{C}_{n})|\mathcal{F}_{n-1})\preceq\mathbf{C}_{n}^{\prime}.$ (128)
@@
 […]
-(by (145) and Fact 3)
+(by (128) and Fact 2)
 […]
@@
 concluding the proof. ∎
-Appendix C Further Notes on Matrix-Valued Martingales
-C.1 Doob’s $L^{p}$ Inequality on Trace
+Appendix C Miscellaneous Technical Results
+C.1 Matrix MGF Bounds
+In Table 1, we state several results in the random concentration literature on controlling the matrix MGF $\mathbf{G}(\gamma)=\mathbb{E}(\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})})$ for a random matrix $\mathbf{X}_{1}$ satisfying certain conditions.
+If these conditions are met conditionally instead, bounds hold conditionally as well.
+Name Condition $\mathbf{G}(\gamma)$
+Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $=\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$
+Rademacher $\mathbf{X}_{1}-\mathbf{M}=R_{1}\mathbf{C},R_{1}\sim\operatorname{rad}_{1/2}$ $\preceq\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$
+Bennett $\lambda_{\max}(\mathbf{X}_{1}-\mathbf{M})\leqslant 1,\mathbf{Var}(\mathbf{X}_{1})=\mathbf{V}$ $\preceq\mathrm{e}^{(\mathrm{e}^{\gamma}-\gamma-1)\mathbf{V}}$
+Bernstein $\mathbb{E}(\mathbf{X}_{1}-\mathbf{M})^{p}\preceq\frac{p!}{2}\mathbf{C}^{2}$ for $p=2,3,\dots$ $\preceq\mathrm{e}^{\frac{\gamma^{2}}{2(1-\gamma)}\mathbf{C}^{2}/2},\gamma\in(0,1)$
+Table 1: Some important matrix MGFs. Here $\mathbf{C},\mathbf{V}$ are deterministic matrices, and $\mathbf{M}=\mathbb{E}\mathbf{X}_{1}$.
+References for Table 1.
+The uni-Gaussian MGF bound is proved by Tropp [2015, Lemma 4.6.2] via a Taylor expansion argument.
+For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a scalar inequality similar to the Hoeffding’s inequality plus the transfer rule (see Section B.2).
+Bennett [Tropp, 2012, Lemma 6.7] and Bernstein [Tropp, 2012, Lemma 6.8] follow from a similar reasoning applied to the respective series expansion.
+∎
+C.2 A Matrix Doob’s $L^{p}$ Inequality
 Applying an operator-monotone and concave function to a matrix supermartingale, we get another matrix supermartingale. In particular,
@@
 The same proof applies for the logarithm as well. ∎
-Fact 3 and Fact 4, on the other hand, convert a matrix-valued (super/sub)martingale to a scalar-valued one. For example,
+Fact 2 and Fact 3, on the other hand, convert a matrix-valued (super/sub)martingale to a scalar-valued one by trace scalarization.
+For example, denoting by $\|\mathbf{X}\|_{p}$ the Schatten $p$-norm of $\mathbf{X}\in\mathcal{S}_{d}$; that is, $\|\mathbf{X}\|_{p}=(\mathrm{tr}((\operatorname{abs}\mathbf{X})^{p}))^{1/p}$.
 Lemma C.2.
-Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$.
-Then, $\{\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})^{p}\}$ is a $\mathbb{R}_{\geqslant 0}$-valued submartingale.
+Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then, $\{\|\mathbf{M}_{n}\|^{p}_{p}\}$ is a submartingale.
 Proof.
-Using Fact 4,
+Using Fact 3,
 $\mathbb{E}(\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})^{p}|\mathcal{F}_{n-1})=\mathrm{tr}\mathbb{E}((\operatorname{abs}\mathbf{M}_{n})^{p}|\mathcal{F}_{n-1})\geqslant\mathrm{tr}(\operatorname{abs}\mathbb{E}(\mathbf{M}_{n}|\mathcal{F}_{n-1}))^{p}.$ (138)
-If $\{\mathbf{M}_{n}\}$ is a martingale, the right hand side above is $\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n-1})^{p}$. If $\mathbf{M}_{n}\succeq 0$ is a submartingale, then we have by Fact 3,
+If $\{\mathbf{M}_{n}\}$ is a martingale, the right hand side above is $\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n-1})^{p}$. If $\mathbf{M}_{n}\succeq 0$ is a submartingale, then we have by Fact 2,
 $\mathrm{tr}(\operatorname{abs}\mathbb{E}(\mathbf{M}_{n}|\mathcal{F}_{n-1}))^{p}\geqslant\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n-1})^{p},$ (139)
 since the function $x\mapsto|x|^{p}$ is monotone on $x\geqslant 0$. ∎
-Proposition C.3 (Trace Doob’s $L^{p}$ Inequality).
-Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then,
-$\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{\mathrm{op}}^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}(\mathrm{tr}\operatorname{abs}\mathbf{M}_{n})^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\mathrm{tr}\operatorname{abs}\mathbf{M}_{N})^{p}\leqslant\left(\frac{pd}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|_{\mathrm{op}}^{p}.$ (157)
+In particular, when $p=1$ in above, the nuclear norm process $\{\|\mathbf{M}_{n}\|_{1}\}$ is a submartingale.
+We now have the following generalization of Doob’s $L^{p}$ maximal inequality (e.g.
+Theorem 12.1 in Klenke [2013]).
+Proposition C.3 (Doob-type Nuclear $L^{p}$ Inequality).
+Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p>1$. Then,
+$\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{1}^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|^{p}_{1}=\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\|\mathbf{M}_{N}\|\cdot\operatorname{r}(\mathbf{M}_{N}))^{p}.$ (140)
 Proof.
-The second inequality is directly from applying Doob’s $L^{p}$ inequality on the nonnegative submartingale $\{\mathrm{tr}(\operatorname{abs}\mathbf{M}_{n})\}$.
-The first and the third $\leqslant$’s are both trivial.
+The second inequality is directly from applying Doob’s $L^{p}$ inequality on the nonnegative submartingale $\{\|\mathbf{M}_{n}\|_{1}\}$.
+The first $\leqslant$ and the final $=$ follow from straightward relations between the spectral and nuclear norms, and the effective rank.
 ∎
-Appendix D Non-symmetric, Non-square Matrices
+C.3 Non-symmetric, Non-square Matrices
 We finally remark on the extension of our methods to non-symmetric and possibly non-square matrices.
@@
 In Table 2, we see that quadratic assumptions on the possibly non-square random matrix $\mathbf{X}_{1}$ translate directly to those on the symmetric random matrix $D(\mathbf{X}_{1})$.
+Conditional distribution assumptions are translated similarly.
 Assumption on $\mathbf{X}_{1}$ Assumption on $D(\mathbf{X}_{1})$
-$\mathbf{X}_{1}^{\mathsf{T}}\mathbf{X}_{1}\preceq\mathbf{B}_{1}$ and $\mathbf{X}_{1}\mathbf{X}_{1}^{\mathsf{T}}\preceq\mathbf{B}_{2}$ $D^{2}(\mathbf{X}_{1})\preceq\left(\begin{matrix}\mathbf{B}_{1}&\\ &\mathbf{B}_{2}\end{matrix}\right)$
+$\mathbf{X}_{1}^{\mathsf{T}}\mathbf{X}_{1}\preceq\mathbf{B}_{1}$ and $\mathbf{X}_{1}\mathbf{X}_{1}^{\mathsf{T}}\preceq\mathbf{B}_{2}$ $D^{2}(\mathbf{X}_{1})\preceq\left(\begin{matrix}\mathbf{B}_{1}&\\ &\mathbf{B}_{2}\end{matrix}\right)$ and thus
+$\left(\begin{matrix}-\mathbf{B}_{1}^{1/2}&\\ &-\mathbf{B}_{2}^{1/2}\end{matrix}\right)\preceq D(\mathbf{X}_{1})\preceq\left(\begin{matrix}\mathbf{B}_{1}^{1/2}&\\ &\mathbf{B}_{2}^{1/2}\end{matrix}\right)$
+$\|\mathbf{X}_{1}\|\leqslant b$ $\|D(\mathbf{X}_{1})\|\leqslant b$
 $\mathbf{Cov}(\mathbf{X}_{1}^{\mathsf{T}})\preceq\mathbf{V}_{1}$ and $\mathbf{Cov}(\mathbf{X}_{1})\preceq\mathbf{V}_{2}$ $\mathbf{Var}(D(\mathbf{X}_{1}))\preceq\left(\begin{matrix}\mathbf{V}_{1}&\\ &\mathbf{V}_{2}\end{matrix}\right)$
@@
 For example, the self-normalized matrix supermartingale (Example 3.8) can be generalized as follows.
-Proposition D.1.
+Example C.4.
 Let $\{\mathbf{X}_{n}\}$ be $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional covariance upper bounds $\mathbf{Cov}(\mathbf{X}_{n}^{\mathsf{T}}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{1n}$, $\mathbf{Cov}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{2n}$ that are predictable.
@@
 the process
-$\mathbf{Y}_{n}=\mathrm{e}^{\mathbf{F}_{1}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{2}/2}\dots\mathrm{e}^{\mathbf{F}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{F}_{n}/2}\dots\mathrm{e}^{\mathbf{E}_{2}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{1}/2}.$ (160)
-is a $\mathcal{S}_{d_{1}+d_{2}}^{+}$-valued supermartingale.
+$\mathbf{Y}_{n}=\mathrm{e}^{\mathbf{F}_{1}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{2}/2}\dots\mathrm{e}^{\mathbf{F}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{E}_{n}/2}\mathrm{e}^{\mathbf{F}_{n}/2}\dots\mathrm{e}^{\mathbf{E}_{2}/2}\mathrm{e}^{\mathbf{F}_{2}/2}\mathrm{e}^{\mathbf{E}_{1}/2}\mathrm{e}^{\mathbf{F}_{1}/2}$ (143)
+is an $\mathcal{S}_{d_{1}+d_{2}}^{+}$-valued supermartingale.

### v6

@@
-Positive Semidefinite Matrix Supermartingales
+Positive Semidefinite Matrix Supermartingales [^1] Accepted for publication in the Electronic Journal of Probability
 Abstract
@@
 The process $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ is called a (matrix-valued) supermartingale if
-for $\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}.$ (1)
+$\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}.$ (1)
 Here, crucially, the order $\preceq$ is the Loewner order, where $\mathbf{A}\preceq\mathbf{B}$ if and only if $\mathbf{B}-\mathbf{A}$ is a positive semidefinite matrix.
-This introduces the rich spectral theory of matrices to the familiar $d=1$ scalar special case as well as the matrix martingale special case where the $\preceq$” in (1) takes the entry-wise “$=$”.
+This introduces the rich spectral theory of matrices to the familiar $d=1$ scalar special case as well as the matrix martingale special case where the $\preceq$” in (1) takes the entry-wise “==”.
 Similarly, if the matrices $\{\mathbf{Y}_{n}\}_{n\geqslant 0}$ are adapted instead to a backward filtration $\{\mathcal{G}_{n}\}_{n\geqslant 0}$, that is, $\mathcal{G}_{0}\supseteq\mathcal{G}_{1}\supseteq\dots$, it is called a (matrix-valued) backward submartingale if
-for $\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n-1}|\mathcal{G}_{n})\succeq\mathbf{Y}_{n}.$ (2)
+$\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n-1}|\mathcal{G}_{n})\succeq\mathbf{Y}_{n}.$ (2)
 In this paper, we primarily study matrix-valued supermartingales and backward submartingales that are almost surely positive semidefinite (PSD).
@@
   1. 1.
-    The convergence and optional stopping theorems, and the Ville’s inequality for PSD matrix supermartingales (Theorems 3.2, 3.3 and 3.4) that generalize the corresponding scalar results.
+    The convergence and optional stopping theorems, and the Ville’s inequality for PSD matrix supermartingales (Theorems 3.2, 3.3, and 3.4) that generalize the corresponding scalar results.
   2. 2.
-    Sequential testing procedure with a matrix test supermartingale (Algorithm 1) that can be applied to the meaning testing of bounded (Examples 3.7 and 3.10), symmetric (Example 3.9), or heavy-tailed (Examples 3.8 and 3.12) martingale-dependent matrices.
+    Sequential testing procedure with a matrix test supermartingale (Algorithm 1) that can be applied to the meaning testing of bounded (Examples 3.7 and 3.11), symmetric (Example 3.10), or heavy-tailed (Example 3.9 and 3.13) martingale-dependent matrices.
   3. 3.
-    Matrix e-values and e-processes (Definitions 3.13 and 3.14), generalizing the corresponding scalar concepts that have recently been extensively studied.
+    Matrix e-values and e-processes (Definitions 3.14 and 3.15), generalizing the corresponding scalar concepts that have recently been extensively studied.
   4. 4.
@@
 We then study backward PSD matrix submartingales, their Ville’s inequality and applications in Section 4.
-In Section 5 we provide several perspectives comparing our inherent matrix approach to some existing scalar-born or scalarized results, and discuss a randomized improvement that is further developed in Appendix A.
+In Section 5 we provide several perspectives comparing our inherent matrix approach to some existing scalar-born or scalarized results, and discuss a randomized improvement that is further developed in Section 7.
 An interesting empirical study is presented in Section 6, including a problem where our matrix method yields higher out-of-the-box power compared to the previous scalar method.
-Omitted proofs are found in Appendix B; some further references and independent theoretical results in Appendix C.
+Omitted proofs are found in Section 8; some further references and independent theoretical results in Section 9.
 2 Preliminaries and Related Work
@@
 Let $\mathcal{S}_{d}$ denote the set of all $d\times d$ real-valued symmetric matrices, which is the only class of matrices considered in this paper.
-[^1] We only briefly discuss the case with non-symmetric and non-square matrices in Section C.3.
+[^2] We only briefly discuss the case with non-symmetric and non-square matrices in Section 9.3.
 These matrices are denoted by bold upper-case letters $\mathbf{A},\mathbf{B}$, etc; whereas $d$ dimensional (column) vectors by bold lower-case letters $\mathbf{u},\mathbf{v}$, etc.
@@
 Any function $f:I\to J$ (where $I,J\subseteq\mathbb{R}$) is identified canonically with a function $f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$ in the following way: for a diagonal matrix $\mathbf{\Sigma}\in\mathcal{S}_{d}^{I}$, $f(\mathbf{\Sigma})$ is obtained by applying $f$ entrywise; otherwise it is obtained by applying $f$ after orthogonal diagonalization, $f(\mathbf{U}^{\mathsf{T}}\mathbf{\Sigma}\mathbf{U})=\mathbf{U}^{\mathsf{T}}f(\mathbf{\Sigma})\mathbf{U}$.
-Notable matrix-to-matrix operators that recur in this paper defined this way include the matrix exponential $\exp:\mathcal{S}_{d}\to\mathcal{S}^{++}_{d}$, the matrix logarithm $\log:\mathcal{S}^{++}_{d}\to\mathcal{S}_{d}$, the matrix absolute value [^2] We use $\operatorname{abs}(\,\cdot\,)$ instead of ||$\cdot$|| ⋅ | to distinguish it from the determinant and other norms.
+Notable matrix-to-matrix operators that recur in this paper defined this way include the matrix exponential $\exp:\mathcal{S}_{d}\to\mathcal{S}^{++}_{d}$, the matrix logarithm $\log:\mathcal{S}^{++}_{d}\to\mathcal{S}_{d}$, the matrix absolute value [^3] We use $\operatorname{abs}(\,\cdot\,)$ instead of $|\cdot|$ to distinguish it from the determinant and other norms.
 $\operatorname{abs}:\mathcal{S}_{d}\to\mathcal{S}^{+}_{d}$, and powers $\mathbf{X}\mapsto\mathbf{X}^{k}$ for any $k\in\mathbb{R}$.
-[…]
+Important functionals $\mathcal{S}_{d}\to\mathbb{R}$ we use throughout the paper include $\lambda_{\max}$, the largest eigenvalue of a matrix; $\|\cdot\|$, the spectral norm of a matrix which equals $\lambda_{\max}\circ\operatorname{abs}$; and $\mathrm{tr}$, the trace.
+Note that $\mathrm{tr}$ preserves the orders (viz., $\mathbf{A}\preceq\mathbf{B}$ implies $\mathrm{tr}\mathbf{A}\leqslant\mathrm{tr}\mathbf{B}$) and is lower bounded by $\lambda_{\max}=\|\cdot\|$ on $\mathcal{S}_{d}^{+}$.
 It is rare to find matrix-to-matrix functions that are monotone (i.e.
@@
 confidence intervals) valid at fixed finite sample sizes, Ville’s inequality (VI) has been the workhorse for many recent time-uniform concentration inequalities and accordingly sequential statistical methods which are valid at data-dependent sample sizes.
-For example, a $(1-\alpha)$-confidence sequence, which is derived almost exclusively via (VI) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b], is a sequence of confidence sets $\{C_{n}\}_{n\geqslant 1}$ for the parameter $\theta$ such that $\mathbb{P}(\theta\in C_{n}\text{ for all }n)\geqslant 1-\alpha$, or equivalently $\mathbb{P}(\theta\in C_{\tau})\geqslant 1-\alpha$ for any stopping time $\tau$ that can possibly depend on the data.
-More generally, the stopped version of (VI) is used in sequential hypothesis testing, as it implies that to construct a conservative sequential test of $\alpha$ level it suffices to construct a process $\{Y_{n}\}$ with $Y_{0}=1$ that is a nonnegative supermartingale under the null hypothesis, and reject it when $Y_{n}$ first surpasses $1/\alpha$.
-Various examples of sequential tests constructed via nonnegative supermartingales can be found in the recent review article by Ramdas et al.
-[2023, Sections 4 and 5].
-We return to this topic of test processes with a more involved discussion in Section 3.4.
+These methods enable conducting hypothesis testing and estimation with uncertainty quantification without being committed to an a priori fixed sample size, allowing the experimenter to stop the experiment for any (possibly data-dependent) reason — formally, at a stopping time $\tau$ on the filtration $\mathcal{F}_{n}=\sigma(X_{1},\dots,X_{n})$ generated by the data sequence $X_{1},X_{2},\dots$.
+  • A $(1-\alpha)$-confidence sequence for the parameter $\theta$ is a sequence of sets $\{C_{n}\}_{n\geqslant 1}$ (each $C_{n}$ computed from $X_{1},\dots,X_{n}$) such that $\mathbb{P}(\theta\in C_{n}\text{ for all }n)\geqslant 1-\alpha$, or equivalently $\mathbb{P}(\theta\in C_{\tau})\geqslant 1-\alpha$ for any stopping time $\tau$.
+They are usually derived via equation (VI) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b].
+  • When testing the null hypothesis that the true data-generating distribution belongs to the set $\mathcal{H}_{0}$, a test supermartingale for $\mathcal{H}_{0}$ is a process $\{Y_{n}\}$ (each $Y_{n}$ computed from $X_{1},\dots,X_{n}$) with $Y_{0}=1$ that is a nonnegative supermartingale, i.e.
+$\mathbb{E}_{P}(Y_{n}|X_{1},\dots,X_{n-1})\leqslant Y_{n-1}$, under any distribution $P\in\mathcal{H}_{0}$.
+One rejects $\mathcal{H}_{0}$ when the observed $Y_{n}$ first surpasses $1/\alpha$.
+This ensures, due to (VI), that the type-I error rate of ever falsely rejecting $\mathcal{H}_{0}$ is at most $\alpha$.
+These (VI)-driven sequential methods are widely applied in real-world cases where it is much more economical (see e.g.
+the argument made by Ham et al.
+[2023]) to leave the sample size unspecified and allow flexible sequential experiments.
+See the recent review article by Ramdas et al.
+[2023, Sections 4 and 5] for numerous examples.
+We revisit this topic with a more involved discussion on test processes in Section 3.4.
 Numerous variations and extensions on (VI) exist.
@@
 In particular, we show that this Loewner-ordered approach pioneered by Ahlswede and Winter [2002] does lead to sharp matrix concentration inequalities and tests, fixed-time and time-uniform alike, comparable to and sometimes better than the results cited in Section 2.3.1.
-We provide a self-contained proof of (MMI) in Appendix A that also comes with the extension of randomization.
+We provide a self-contained proof of (MMI) in Section 7 that also comes with the extension of randomization.
 3 Forward PSD Supermartingales
@@
 Recall our definition of a matrix-valued supermartingale from the beginning of the paper that an adapted $\mathcal{S}_{d}$-valued stochastic process $\{\mathbf{Y}_{n}\}$ is a supermartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}$ holds for every $n$.
-Before we proceed, note that the order $\preceq$ is preserved by addition, limit, and Lebesgue integration (hence including convergent infinite sums, expectations, conditional expectations, etc., essentially because $\mathcal{S}^{+}=\{\mathbf{A}:\mathbf{A}\succeq 0\}$ is closed both topologically and under $+$).
-See also a “dominated integrability” lemma we prove in Section B.1.
+Before we proceed, note that the order $\preceq$ is preserved by addition, limit, and Lebesgue integration (hence including convergent infinite sums, expectations, conditional expectations, etc., essentially because $\mathcal{S}^{+}=\{\mathbf{A}:\mathbf{A}\succeq 0\}$ is closed both topologically and under ++).
+See also a “dominated integrability” lemma we prove in Section 8.1.
 Matrix supermartingales are related to scalar supermartingales in the following ways.
 First, the monotonicity of trace implies that if $\{\mathbf{Y}_{n}\}$ is an $\mathcal{S}_{d}$-valued supermartingale, its trace process $\{\mathrm{tr}(\mathbf{Y}_{n})\}$ is a real-valued supermartingale.
-Second, we have the following lemma.
+Second, we have the following lemma ($\mathbb{Q}$ being the set of rational numbers).
 Lemma 3.1.
-An $\mathcal{S}_{d}$-valued adapted process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every non-random vector $\mathbf{v}\in\mathbb{R}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale.
-The very straightforward proof of Lemma 3.1 can be found in Section B.2. It leads to an analog of Doob’s martingale convergence theorem for PSD matrix supermartingales.
+(1) An $\mathcal{S}_{d}$-valued adapted process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every non-random vector $\mathbf{v}\in\mathbb{R}^{d}$ or $\mathbf{v}\in\mathbb{Q}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale.
+(2) Similar to the scalar case, an $\mathcal{S}_{d}$-valued integrable process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if the predictable component of its Doob decomposition $\mathbf{A}_{n}=\sum_{i=1}^{n}(\mathbb{E}(\mathbf{Y}_{i}|\mathcal{F}_{i-1})-\mathbf{Y}_{i-1})$ is decreasing in the $\preceq$ order.
+The proof of Lemma 3.1 can be found in Section 8.2. It leads to an analog of Doob’s martingale convergence theorem for PSD matrix supermartingales.
 Theorem 3.2 (Matrix Supermartingale Convergence Theorem).
@@
 Theorem 3.3 (Matrix Optional Stopping).
-Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\tau$ be a stopping time. Then,
+Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\sigma\leqslant\tau$ be two stopping times. Then:
   1. 1.
-    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale.
+    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale on both $\{\mathcal{F}_{n}\}$ and $\{\mathcal{F}_{n\wedge\tau}\}$.
   2. 2.
-    If $\tau\leqslant N$ almost surely for some $N$, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
-  3. 3.
-    If $\{\mathbf{Y}_{n}\}$ is $\mathcal{S}_{d}^{+}$-valued, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
+    If (a) $\tau<N$ for some finite $N$, or (b) $\tau<\infty$ and $\{\mathbf{Y}_{n}\}$ is $\mathcal{S}_{d}^{+}$-valued, then $\mathbb{E}(\mathbf{Y}_{\tau}|\mathcal{F}_{\sigma})\preceq\mathbf{Y}_{\sigma}$.
+Consequently, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{\sigma}\preceq\mathbb{E}\mathbf{Y}_{0}$.
 Proof.
-1.
-For any $\mathbf{v}\in\mathbb{R}^{d}$, the scalar process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale on $\{\mathcal{F}_{n}\}$ and therefore so is the stopped process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n\wedge\tau}\mathbf{v}\}$ due to the optional stopping theorem (e.g.
+The proofs of two statements are as follows.
+  1. 1.
+    For any $\mathbf{v}\in\mathbb{R}^{d}$, the scalar process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale on $\{\mathcal{F}_{n}\}$ and therefore so is the stopped process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n\wedge\tau}\mathbf{v}\}$ on both $\{\mathcal{F}_{n}\}$ and $\{\mathcal{F}_{n\wedge\tau}\}$ due to the optional stopping theorem (e.g.
 Theorem 10.15 of Klenke [2013]).
-This implies that $\{\mathbf{Y}_{n\wedge\tau}\}$ is a matrix supermartingale on $\{\mathcal{F}_{n}\}$.
-2.
+This implies that $\{\mathbf{Y}_{n\wedge\tau}\}$ is a matrix supermartingale on both $\{\mathcal{F}_{n}\}$ and $\{\mathcal{F}_{n\wedge\tau}\}$.
+  2. 2.
     For any $\mathbf{v}\in\mathbb{R}^{d}$, we apply the optional sampling theorem (e.g.
-Theorem 10.11(i) of Klenke [2013]) to the scalar-valued supermartingale $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ to obtain that $\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{v}=\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\tau}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{0}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{0})\mathbf{v}$.
-Therefore $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
-3.
-Applying Theorem 3.2 to the $\mathcal{S}_{d}^{+}$-valued stopped supermartingale $\{\mathbf{Y}_{\tau\wedge n}\}$, we have $\mathbb{E}\mathbf{Y}_{\tau\wedge\infty}\preceq\mathbb{E}\mathbf{Y}_{\tau\wedge 0}$, i.e., $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
+Theorem 10.11(i,ii) of Klenke [2013]) to the scalar-valued supermartingale $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ to obtain that $\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\sigma}\mathbf{v}\geqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\tau}\mathbf{v}|\mathcal{F}_{\sigma})=\mathbf{v}^{\mathsf{T}}\mathbb{E}(\mathbf{Y}_{\tau}|\mathcal{F}_{\sigma})\mathbf{v}$.
+Therefore $\mathbf{Y}_{\sigma}\succeq\mathbb{E}(\mathbf{Y}_{\tau}|\mathcal{F}_{\sigma})$.
 ∎
@@
 or equivalently,
-for any stopping time $\text{for any stopping time }\tau,\quad\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
+$\text{for any stopping time }\tau,\quad\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
 Proof.
-We prove the stopped inequality first. Due to Part 3 of Theorem 3.3, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. Now by (MMI) and the monotinicity of the trace,
+We prove the stopped inequality first. Due to Part 3 of Theorem 3.3, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. Now by (MMI) and the monotonicity of the trace,
 $\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})\leqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1/2}),$ (12)
@@
     While the one-time Markov case reveals the potential tightness benefit of (MVI), in real data the conditions for such benefit often arise naturally in the sequential, anytime-valid testing case that necessitates Ville-type inequalities.
-As an aside, we present the additional result of a matrix analog of Doob’s $L^{p}$ inequality for forward matrix martingales in Section C.2.
+As an aside, we present the additional result of a matrix analog of Doob’s $L^{p}$ inequality for forward matrix martingales in Section 9.2.
 It is not necessarily relevant to our upcoming discussion in statistical contexts but might be of independent interest.
 3.2 Matrix Test Supermartingales
-Scalar-valued nonnegative supermartingales are often constructed incrementally: if one can find an adapted nonnegative process $\{E_{n}\}$ that has the conditional expectation bound $\mathbb{E}(E_{n}|\mathcal{F}_{n})\leqslant 1$, then their product $Y_{n}=\prod_{i=1}^{n}E_{i}$ is a nonnegative supermartingale.
-In particular, the supermartingale satisfies $Y_{0}=1$ and thus satisfies (VI) with $\mathbb{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$.
-It is called a test supermartingale.
+Scalar-valued nonnegative supermartingales are often constructed incrementally: if one can find an adapted nonnegative process $\{E_{n}\}$ that has the conditional expectation bound $\mathbb{E}(E_{n}|\mathcal{F}_{n})\leqslant 1$, then their product $Y_{n}=\prod_{i=1}^{n}E_{i}$ is a nonnegative supermartingale which satisfies $Y_{0}=1$ and thus satisfies (VI) with $\mathbb{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}$.
 We propose the following analogous way to construct $\mathcal{S}_{d}^{+}$-valued supermartingales that involves a bit more flexibility.
-Lemma 3.5 (Construction of Matrix Test Supermartingales).
+Lemma 3.5 (Incremental Construction of Matrix Supermartingales).
 If an adapted $\mathcal{S}_{d}^{+}$-valued process $\{\mathbf{E}_{n}\}_{n\geqslant 1}$ and a predictable $\mathcal{S}_{d}^{++}$-valued process $\{\mathbf{F}_{n}\}_{n\geqslant 1}$ satisfy $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}^{-1}$ almost surely for all $n\geqslant 1$.
@@
 $\mathbf{Y}_{n}=\sqrt{\mathbf{F}_{1}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{2}}\dots\sqrt{\mathbf{F}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{F}_{n}}\dots\sqrt{\mathbf{E}_{2}}\sqrt{\mathbf{F}_{2}}\sqrt{\mathbf{E}_{1}}\sqrt{\mathbf{F}_{1}},$ (15)
-with $\mathbf{Y}_{0}=\mathbf{I}$. It is called a matrix test supermartingale.
+with $\mathbf{Y}_{0}=\mathbf{I}$.
 In some cases, we can take $\mathbf{F}_{1}=\mathbf{F}_{2}=\dots=\mathbf{I}$, where the process has a simpler expression $\mathbf{Y}_{n}=\sqrt{\mathbf{E}_{1}}\dots\sqrt{\mathbf{E}_{n}}\sqrt{\mathbf{E}_{n}}\dots\sqrt{\mathbf{E}_{1}}$.
@@
 end for
+Recall from Section 2.2 the concept of testing a null hypothesis $\mathcal{H}_{0}$ with a scalar test supermartingale $\{Y_{n}\}$: a process with $Y_{0}=1$ that is a nonnegative supermartingale under $\mathcal{H}_{0}$.
 With Lemma 3.5, we can recover matrix versions of several scalar test supermartingales in the recent sequential statistics literature.
@@
 and one tests if their common mean equals a specific value.
+It is not straightforward to see that some of these examples satisfy the assumptions of Lemma 3.5, so we prove them in Section 8.2.
 First, we have the moment generating function (MGF)-based supermartingales whose well-known scalar case is mentioned by e.g. Howard et al. [2020, Fact 1(a)].
@@
 These log-MGF upper bounds only lead to scalar-valued supermartingales (9) instead of matrix-valued supermartingales.
-We review some existing results of matrix MGFs by Tropp [2012, 2015] in Section C.1.
+We review some existing results of matrix MGFs by Tropp [2012, 2015] in Section 9.1.
 Our second example extends the wealth process of betting against the fairness of a coin.
@@
 where
-$-\frac{1}{\lambda_{\max}(\mathbf{B}_{n}-\mathbf{M}_{n})}<\gamma_{n}<\frac{1}{\lambda_{\max}(\mathbf{M}_{n})}$ (21)
+$\frac{1}{\lambda_{\max}(\mathbf{B}_{n}-\mathbf{M}_{n})}<\gamma_{n}<\frac{1}{\lambda_{\max}(\mathbf{M}_{n})}$ (21)
 is predictable.
 To see that this is a valid $(\mathbf{E}_{n},\mathbf{F}_{n})$ pair for Lemma 3.5, it is clear that $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})=\mathbf{I}$ and (21) ensures $\mathbf{E}_{n}\succeq 0$.
+The resulting $\{\mathbf{Y}_{n}\}$ is therefore an $\mathcal{S}_{d}^{+}$-valued matrix martingale.
+Remark 3.8.
+Using Example 3.7 to test the null $\mathcal{H}_{0}:\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ with Algorithm 1, while any choice of predictable $\{\gamma_{n}\}$ that satisfies (21) makes sure the type-I error $\leqslant\alpha$ under $\mathcal{H}_{0}$, different choices of the sequence lead to difference in power when $\mathcal{H}_{0}$ does not hold.
+A simple example is stated below.
+Suppose $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}+c\mathbf{I}$ for some unknown $c$ ($c=0$ under null $\mathcal{H}_{0}$).
+Since $\mathbb{E}[\mathbf{E}_{n}|\mathcal{F}_{n-1}]=(1+\gamma_{n}c)\mathbf{I}$, we can see from the proof of Lemma 3.5 that the “growth” event $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq\mathbf{Y}_{n-1}$ happens if and only if the predictable $\gamma_{n}$ and the ground-truth $c$ have matching signs: $\gamma_{n}c\geqslant 0$.
+If the sign of $c$ is not known beforehand, there are two ways to achieve this:
+  • Since $\gamma_{n}$ can be picked predictably, one can estimate $c$ from $\mathbf{X}_{1},\dots,\mathbf{X}_{n-1}$ and match the sign of $\gamma_{n}$ with this estimator.
+For example, $\hat{c}_{n-1}=\mathrm{tr}((\mathbf{X}_{1}-\mathbf{M}_{1})+\cdots+(\mathbf{X}_{n-1}-\mathbf{M}_{n-1}))/(d(n-1))$, and $\gamma_{n}=\operatorname{sgn}(\hat{c}_{n-1})\gamma$ for some $0<\gamma<\min\left\{\frac{1}{\lambda_{\max}(\mathbf{B}_{n}-\mathbf{M}_{n})},\frac{1}{\lambda_{\max}(\mathbf{M}_{n})}\right\}$.
+  • (Hedging).
+On the other hand, one may launch two test martingales, $\{\mathbf{Y}_{n}^{+}\}$ with $\gamma_{n}>0$, and $\{\mathbf{Y}_{n}^{-}\}$ with $\gamma_{n}<0$ for all $n$.
+Since one of them must match the correct sign of $c$ if $c\neq 0$, one can use $\{(\mathbf{Y}_{n}^{+}+\mathbf{Y}_{n}^{-})/2\}$ as the test process (which is still a test martingale for the null $\mathcal{H}_{0}:c=0$) to have power in detecting any $c\neq 0$.
+The same logic applies to many of the subsequent examples.
 Besides these two examples where we extend prior scalar results to matrices, we also note that Howard et al.
@@
 [2020] do not consider combining these conditional expectation bounds into matrix-valued supermartingales via (15); they pursue scalar supermartingales instead, with methods we discuss later in Section 5.2.
-Of their eight inequalities, we will introduce two interesting cases now and (re-)prove them in Section B.2, omitting the other six due to their similar construction.
+Of their eight inequalities, we will introduce two interesting cases now and (re-)prove them in Section 8.2, omitting the other six due to their similar construction.
 First, the self-normalized supermartingale that only assumes the second moment exists [Howard et al., 2020, Part (f) of Section 6.5].
-Example 3.8 (Self-Normalized Matrix Supermartingale).
+Example 3.9 (Finite-Variance Self-Normalized Matrix Supermartingale).
 Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional variance upper bounds $\mathbb{E}((\mathbf{X}_{n}-\mathbf{M}_{n})^{2}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{n}$ that are predictable.
@@
 where $\{\gamma_{n}\}$ are predictable positive scalars.
+As a bibliographical remark, we follow the usage of De la Pena et al.
+[2009], Bercu and Touati [2008, 2019], among others, in adopting the name “self-normalized”.
+These authors refer to a martingale $\{M_{n}\}$ on $\{\mathcal{F}_{n}\}$ as “self-normalized” if maximal inequalities of $\{Y_{n}\}$ are derived via the predictable quadratic variation $\sum\mathbb{E}((Y_{n}-Y_{n-1})^{2}|\mathcal{F}_{n-1})$ and the total quadratic variaiton $\sum(Y_{n}-Y_{n-1})^{2}$.
+The summands of these two quadratic variations correspond loosely to the $\mathbf{V}_{n}$ and $(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}$ terms in (23).
+The next example further demonstrates this concept.
 Second, a bound for symmetric random matrices that have a symmetric (around $0\in\mathcal{S}_{d}$), potentially heavy-tailed distribution [Howard et al., 2020, Part (d) of Section 6.5].
-Example 3.9.
+Example 3.10 (Symmetric Self-Normalized Matrix Supermartingale).
 Let $\{\mathbf{X}_{n}\}$ be an adapted sequence of $\mathcal{S}_{d}$-valued random matrices with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and satisfy $\mathbf{X}_{n}-\mathbf{M}_{n}|\mathcal{F}_{n-1}\stackrel{{\scriptstyle d}}{{=}}\mathbf{M}_{n}-\mathbf{X}_{n}|\mathcal{F}_{n-1}$.
@@
 The following is a rephrasing of a recent result by Wang and Ramdas [2024, Lemma 4.1] in the language of our Lemma 3.5.
-Example 3.10 (Empirical Bernstein-Type Matrix Supermartingale).
+Example 3.11 (Empirical Bernstein Matrix Supermartingale).
 Let $\{\mathbf{X}_{n}\}$ be $\mathcal{S}_{d}$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$.
@@
 Definition 3.12.
-An increasing function $\phi:\mathbb{R}\to\mathbb{R}$ is upper $p$-logarithmically contractive if $\phi(x)\leqslant\log(1+x+|x|^{p}/p)$; lower $p$-logarithmically contractive if $-\log(1-x+|x|^{p}/p)\leqslant\phi(x)$.
+An increasing function $\phi:\mathbb{R}\to\mathbb{R}$ is upper $p$-logarithmically contractive if $\phi(x)\leqslant\log(1+x+|x|^{p}/p)$; lower $p$-logarithmically contractive if $\log(1-x+|x|^{p}/p)\leqslant\phi(x)$.
 For example, the function
@@
 Above, $\{\gamma_{n}\}$ are predictable positive scalars.
-Example 3.12 is proved in Section B.2.
+Example 3.13 is proved in Section 8.2.
 We remark that, in the scalar case, the two supermartingales constructed this way remain supermartingales under one-sided nulls, one for each side.
@@
 Throughout this subsection, we consider $X_{1},X_{2},\dots$ a sequence of random observations taking values in some space $\mathcal{X}$ and they generate the filtration $\{\mathcal{F}_{n}\}$.
-We use notations $\mathbb{P}_{P}$, $\mathbb{E}_{P}$ and the general terminology “for $P$” (e.g.
+We use notations $\mathbb{P}_{P}$, $\mathbb{E}_{P}$ and the general terminology “for PP” (e.g.
 we may say an adapted stochastic process is a martingale for $P$) to refer to probabilistic statements when the observations $X_{1},X_{2},\dots$ are distributed jointly according to $P$, where $P$ is some distribution on $\mathcal{X}\times\mathcal{X}\times\dots$.
@@
 $\text{For any $P\in\mathcal{P}$}:\quad\mathbb{P}_{P}(\exists n,\ Y_{n}\geqslant a)\leqslant a^{-1};$ (30)
-and for any stopping time $\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}.$ (31)
+$\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(Y_{\tau}\geqslant a)\leqslant a^{-1}.$ (31)
 This property enables scalar e-processes to also be used for testing the null hypothesis that the data is drawn from some $P\in\mathcal{P}$: we may reject the null as soon as the e-process exceeds $1/\alpha$ (for some prespecified constant $\alpha\in(0,1)$) and be guaranteed that the type-I error is at most $\alpha$.
@@
 It can be seen from the matrix optional stopping theorem (Theorem 3.3) that if $\{\mathbf{Y}_{n}\}$ is a PSD supermartingale for $\mathcal{P}$ then it is also an e-process for $\mathcal{P}$.
-More generally, $\{\mathbf{Y}_{n}\}$ is a matrix e-process for $\mathcal{P}$ if, for any $P\in\mathcal{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{P}\}$ with $\mathbf{M}_{0}^{P}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{P}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $P$.
+More generally, the following is a sufficient condition for a matrix process to be an e-process:
+Proposition 3.16.
+Suppose for any $P\in\mathcal{P}$, there exists an $\mathcal{S}_{d}^{+}$-valued supermartingale $\{\mathbf{M}_{n}^{P}\}$ for $P$ with $\mathbf{M}_{0}^{P}\preceq\mathbf{I}$ that upper bounds the process $\{\mathbf{Y}_{n}\}$, i.e., $\mathbf{M}_{n}^{P}\succeq\mathbf{Y}_{n}$ for all $n$ almost surely for $P$.
+Then, $\{\mathbf{Y}_{n}\}$ is a matrix e-process for $\mathcal{P}$.
+This can be seen by simply noting that $\mathbb{E}_{P}\mathbf{Y}_{\tau}\preceq\mathbb{E}_{P}\mathbf{M}_{\tau}^{P}\preceq\mathbf{I}$ for any $P,\tau$.
+While the converse to Proposition 3.16 (that it is a necessary condition for e-processes as well) is true in the scalar case [Ramdas et al., 2020, Lemma 6], the argument fails to extend to matrices as the construction for scalars relies on the Snell envelope which we have yet to find a suitable generalization for matrices due to the non-existence of maximum and minimum under $\preceq$ within $\mathcal{S}_{d}$.
 Similar to the scalar case above, matrix e-processes satisfy (MVI), both time-uniform and at a stopping time:
 $\text{For any $P\in\mathcal{P}$}:\quad\mathbb{P}_{P}(\exists n,\ \mathbf{Y}_{n}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1});$ (32)
-and for any stopping time $\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1}),$ (33)
+$\text{and for any stopping time $\tau$},\quad\mathbb{P}_{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}(\mathbf{A}^{-1}),$ (33)
 which follows directly from the proof of Theorem 3.4. Therefore, one can run a procedure similar to Algorithm 1 with an e-process instead for sequential hypothesis testing.
@@
 $\{\mathbf{Y}_{\tau}\mathbbmss{1}_{E}\npreceq\mathbf{A}\}=E,\quad\mathbf{Y}_{\tau}\mathbbmss{1}_{\overline{E}}=\mathbf{Y}_{N}\mathbbmss{1}_{\overline{E}}.$ (41)
-Thus by Lemmas 4.1 and 5.3,
+Thus by Lemma 4.1 and (MMI),
 $\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{N})\mathbf{A}^{-1/2})\geqslant\mathrm{tr}(\mathbf{A}^{-1/2}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{A}^{-1/2})$ (42)
@@
 $\mathbb{P}\left(\exists n\leqslant N,\mathbf{P}_{n}\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{P}_{N})\mathbf{A}^{-1}).$ (46)
-Substituting $M_{n}$ back to the inequality above, we have,
+Substituting $\mathbf{Y}_{n}$ back to the inequality above, we have,
 $\mathbb{P}\left(\exists n\leqslant N,\mathbf{Y}_{n}\npreceq\mathbf{A}\right)\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (47)
-The desired inequality is immediately proved by appealing to the fact that
-$\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}=\liminf_{N\to\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}=\bigcup_{k=1}^{\infty}\bigcap_{N=k}^{\infty}\left\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\right\}.\qed$ (48)
+The desired inequality follows from the monotone convergence theorem: the random variable $\mathbbmss{1}_{\{\exists n,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}}$ is the pointwise limit of the increasing sequence of random variables $\{\mathbbmss{1}_{\{\exists n\leqslant N,\;\mathbf{Y}_{n}\npreceq\mathbf{A}\}}\}_{N\geqslant 1}$.
+∎
 4.2 Application: Concentration of Exchangeable Matrices
@@
 The scalar exchangeable Chebyshev inequality was recently proved by Ramdas and Manole [2024, Theorem 2].
-Recall from Fact 1 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $21\leqslant p<$ 2) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
+Recall from Fact 1 that when $p\in[1,2]$, the function $\mathbf{X}\mapsto\mathbf{X}^{p}$ is operator convex within the domain $\mathcal{S}_{d}^{+}$ (for $1\leqslant p<2$) or $\mathcal{S}_{d}$ (for $p=2$) respectively,
 $\frac{1}{n}(\mathbf{X}_{1}^{p}+\dots+\mathbf{X}_{n}^{p})\succeq\left(\frac{\mathbf{X}_{1}+\dots+\mathbf{X}_{n}}{n}\right)^{p}.$ (48)
@@
 Lemma 4.4.
-Let $p\in[1,2]$, and $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued, $L^{p}$ random matrices.
+Let $p\in[1,2]$, and $\{\mathbf{X}_{n}\}$ be an exchangeable sequence of $\mathcal{S}_{d}$-valued random matrices with $\mathbb{E}\|\mathbf{X}_{1}\|^{p}<\infty$.
 Suppose $\mathbf{C}\in\mathcal{S}_{d}$ is deterministic, and either of the following holds:
@@
 Recalling that the conditional expectation w.r.t.
-the exchangeable $\sigma$-algebra $\mathcal{E}_{n+1}$ equals the average over all $(n+1)!$ permutations of the indices $11,2,\dots,n+$ 1 (see e.g.
+the exchangeable $\sigma$-algebra $\mathcal{E}_{n+1}$ equals the average over all $(n+1)!$ permutations of the indices $1,2,\dots,n+1$ (see e.g.
 Klenke [2013, Theorem 12.10]), we have
 $\mathbb{E}[\mathbf{R}_{n}|\mathcal{E}_{n+1}]=\frac{1}{(n+1)!}\sum_{\sigma\in\mathrm{S}_{n+1}}\left(\frac{\mathbf{X}_{\sigma(1)}+\dots+\mathbf{X}_{\sigma(n)}}{n}-\mathbf{C}\right)^{p}$ (51)
-(${{\succeq}}$ $\left(\frac{\sum_{\sigma\in\mathrm{S}_{n+1}}\left(\frac{\mathbf{X}_{\sigma(1)}+\dots+\mathbf{X}_{\sigma(n)}}{n}-\mathbf{C}\right)}{(n+1)!}\right)^{p}=\mathbf{R}_{n+1},$ (53)
+[…]
 concluding the proof that $\{\mathbf{R}_{n}\}$ is a backward submartingale. ∎
@@
 $\mathbb{P}(\exists n\geqslant N,\;(\overline{\mathbf{X}}_{n}-\mathbf{M})^{2}\npreceq\mathbf{A}^{2})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{R}_{N})\mathbf{A}^{-2})\leqslant N^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2})$ (57)
-where the last step is due to Lemma B.2. ∎
+where the last step is due to Lemma 8.2. ∎
 We remark that the assumption $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for $i\neq j$ is satisfied when $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ are pairwise uncorrelated, which is in turn satisfied under martingale difference assumption (i.e., $\{\mathbf{X}_{i}\}$ is adapted to some filtration $\{\mathcal{F}_{i}\}$ with $\mathbb{E}(\mathbf{X}_{i}|\mathcal{F}_{i-1})=\mathbf{M}$ and $\mathbf{Var}(\mathbf{X}_{i}|\mathcal{F}_{i-1})\preceq\mathbf{V}$), which is in turn implied if they are i.i.d.
@@
 We remark that, from Theorem 4.5 to Theorem 4.6, we weaken the second moment to a $p$th moment assumption, but in an imperfect manner as we obtain a concentration not around the mean $\mathbf{M}$ but a Markov-type “upper tail bound” for $\mathcal{S}_{d}^{+}$-valued matrices instead.
-This is largely because the scalar map $x\mapsto|x|^{p}$ is convex on (,)($\infty$,$\infty$)( - ∞ , ∞ ) but not its matrix extension, and so the raw $p$th moment is used above.
-Also, Theorem 4.6 fails if $p>2$ because the scalar map $x\mapsto x^{p}$ is convex on $[0,\infty)$ but not its matrix extension.
+The raw $p$th moment is used above.
+Also, $p>2$ is not allowed.
+The reason is summarized in Table 1.
+$p$ $x\mapsto x^{p}$, $x\geqslant 0$ $\mathbf{X}\mapsto\mathbf{X}^{p}$, $\mathbf{X}\succeq 0$ $\mapsto|x|^{p}, xx\in\mathbb{R}$ $\mathbf{X}\mapsto(\operatorname{abs}\mathbf{X})^{p}$, $\mathbf{X}\in\mathcal{S}_{d}$
+$[1,2)$ convex convex convex non-convex
+22 convex convex convex convex
+$(2,\infty)$ convex non-convex convex non-convex
+Table 1: Convexity of scalar-to-scalar vs. matrix-to-matrix power maps.
 Nonetheless, thanks to Fact 3, we can resort to trace scalarization to obtain the following spectral and trace inequalities that hold for all $p\geqslant 1$.
@@
 $\mathbb{P}(\exists n,\;\|\overline{\mathbf{X}}_{n}-\mathbf{M}\|\geqslant a)\leqslant\mathbb{P}\left(\exists n,\;\mathrm{tr}\left(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})^{p}\right)\geqslant a^{p}\right)\leqslant a^{-p}\mathrm{tr}\mathbf{V}_{p}.$ (61)
-The inequality above, which we defer its straightforward proof to Section B.3, differs from Theorem 4.6 in that the concentration is now around the mean $\mathbf{M}$.
+The inequality above, which we defer its straightforward proof to Section 8.3, differs from Theorem 4.6 in that the concentration is now around the mean $\mathbf{M}$.
 5 Remarks and Extensions
 5.1 Tightness of Matrix Markov’s and Ville’s Inequalities
-[…]
+Recall that a PSD supermartingale $\{\mathbf{Y}_{n}\}$ can be converted into a scalar nonnegative supermartingale via either taking the trace $\{\mathrm{tr}\mathbf{Y}_{n}\}$ or a quadratic form $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$.
+We thus discuss the tightness of (MVI) in this subsection in comparison to the scalar (VI) on these transformations.
+Since (MVI) is proved by applying (MMI) on a stopped random matrix, it suffices to compare (MMI) versus (MI) on a PSD random matrix $\mathbf{X}$.
+Without loss of generality, we assume that $\mathbb{E}\mathbf{X}=\mathbf{I}$.
+Denote by $\mathcal{C}_{\mathrm{tr}}(\alpha)$ the concentration region by applying (MI) to $\mathrm{tr}\mathbf{X}$:
 $\mathcal{C}_{\mathrm{tr}}(\alpha)=\{\mathbf{X}\in\mathcal{S}_{d}^{+}:\mathrm{tr}\mathbf{X}\leqslant d/\alpha\};$ (62)
@@
   4. 4.
-    If $\mathbf{A}\in\mathcal{A}_{\alpha}$, then $\mathcal{C}(\mathbf{A})\subseteq\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ if and only if $\alpha^{\prime}\leqslant d(\mathrm{tr}\mathbf{A})^{-1}\in[\frac{\alpha}{d},\frac{\alpha}{\operatorname{r}(\mathbf{A}^{-1})}]$;
+    If $\mathbf{A}\in\mathcal{A}_{\alpha}$, then $\mathcal{C}(\mathbf{A})\subseteq\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ if and only if $\alpha^{\prime}\leqslant d(\mathrm{tr}\mathbf{A})^{-1}$ $\in(0,\frac{\alpha}{\operatorname{r}(\mathbf{A})}]$;
   5. 5.
@@
 The first two state that (MMI) and trace-(MI) are mutually incomparable unless one takes the threshold matrix $\mathbf{A}$ to be a multiple of $\mathbf{I}$, in which case (MMI) is weaker than trace-(MI).
-Statements 3 and 4 indicate that (MMI) and trace-(MI) can simulate each other only by increasing the error probability $\alpha$, and that if $\mathbf{A}^{-1}$ is chosen to have a low effective rank, (MMI) can easily simulate trace-(MI) whereas trace-(MI) pays a price up to $d$ to simulate (MMI).
+Statements 3 and 4 indicate that (MMI) and trace-(MI) can simulate each other only by increasing the error probability $\alpha$ substantially.
 Statements 5 and 6 imply respectively that, while one can always simulate quadratic form-(MI) with (MMI) by infinitesimally increasing the error probability $\alpha$, the converse is never true.
-Proposition 5.1 is proved in Section B.4.
+Proposition 5.1 is proved in Section 8.4.
 Similar reasoning holds for (MVI) versus (VI) as well because the concentration regions, while holding time-uniformly or at a stopping time, remain in the same forms of $\mathcal{C}_{\mathrm{tr}}(\alpha)$, $\mathcal{C}_{\mathbf{v}}(\alpha)$, and $\mathcal{C}(\mathbf{A})$.
@@
 $L_{n}\geqslant\exp\left(\lambda_{\max}\left(\sum_{i=1}^{n}\mathbf{Z}_{i}\right)-\lambda_{\max}\left(\sum_{i=1}^{n}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})\right)\right).$ (69)
-The proof of this lemma can be found in Section B.5. We have numerous remarks regarding this lemma.
+The proof of this lemma can be found in Section 8.5. We have numerous remarks regarding this lemma.
 First, the upper bounded expression (69) is an instance of e-processes (if divided by $L_{0}=d$), introduced earlier in Section 3.4.
@@
 Second, the divergence from the common assumption (66) to the matrix-valued (67) and scalar-valued (68) test supermartingale indicates that for some testing problems, two different routes can be pursued: our “matrix-native” approach, and the scalarization approach to the test process.
-Examples in Section 3.3 that fit in the assumption (67) include Examples 3.8 and 3.9 which Howard et al.
+Examples in Section 3.3 that fit in the assumption (66) include Examples 3.9 and 3.10 which Howard et al.
 [2020, Section 6.5] originally prove, but also Example 3.11: To see that Example 3.11 is in the form of (66),
@@
 One of the crucial ways the two approaches differ is that with the scalar-valued test process (68), one rejects the null when the process surpasses $d/\alpha$; whereas with the matrix-valued test process (67), one rejects when the process $\npreceq\mathbf{A}$ where $\mathbf{A}$ is a matrix chosen freely beforehand, only subject to the “size constraint” that $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$.
-Indeed, the extra degrees of freedom that come with $\mathbf{A}$ play a crucial role as we compare the two approaches empirically with the self-normalizing example (Example 3.8) in Section 6.
+Indeed, the extra degrees of freedom that come with $\mathbf{A}$ play a crucial role as we compare the two approaches empirically with the finite-variance self-normalizing example (Example 3.9) in Section 6.
 Third, the assumption (66) for scalarization being more restrictive than our Lemma 3.5, for some testing problems our concept of matrix-valued test supermartingales is unavoidable.
-These include Examples 3.7 and 3.12, to which the stepwise scalarization Lemma 5.2 is not applicable.
+These include Example 3.7 and 3.13, to which the stepwise scalarization Lemma 5.2 is not applicable.
 Their one-step condition $\mathbb{E}(\mathbf{E}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{F}_{n}$ leads naturally to a matrix-valued test supermartingale, and scalar-valued test supermartingales might be obtained by taking the trace or quadratic form after the formation of the matrix-valued one, which is not ideal according to Section 5.1.
@@
 The uniformly randomized Ville’s inequality (8) by Ramdas and Manole [2024] is obtained by applying the uniformly randomized Markov’s inequality [Ramdas and Manole, 2024, Theorem 1.2]
-$\mathbb{P}(X\leqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (72)
+$\mathbb{P}(X\geqslant Ua)\leqslant(\mathbb{E}X)a^{-1}$ (71)
 where $U\sim\operatorname{Unif}_{(0,1)}$ independent from $X$. Here, we state a matrix version of (71), or a randomized version of the matrix Markov inequality (MMI).
@@
 The randomization factor $U$, we note, can be any random variable that is stochastically larger than $\operatorname{Unif}_{(0,1)}$ (“super-uniform”), that is, $\mathbb{P}(U\leqslant x)\leqslant x$ for all $x\geqslant 0$.
-More generally, we shall present and prove a stronger version of Theorem 5.3 in Appendix A, where the scalar randomization factor $U$ is further replaced by a PSD random matrix whose distribution belongs to a class of “super-uniform” measures on $\mathcal{S}_{d}^{+}$.
-The generalized version of Theorem 5.3 in Appendix A also reveals the equality condition for (MMI), randomized or not.
+More generally, we shall present and prove a stronger version of Theorem 5.3 in Section 7, where the scalar randomization factor $U$ is further replaced by a PSD random matrix whose distribution belongs to a class of “super-uniform” measures on $\mathcal{S}_{d}^{+}$.
+The generalized version of Theorem 5.3 in Section 7 also reveals the equality condition for (MMI), randomized or not.
 Theorem 5.3 implies that various matrix concentration inequalities by Ahlswede and Winter [2002] can be tightened via a randomization factor as well.
@@
 $\mathbb{P}(\operatorname{abs}(\overline{\mathbf{X}}_{n}-\mathbf{M})\npreceq\sqrt{U}\mathbf{A})\leqslant n^{-1}\mathrm{tr}(\mathbf{V}\mathbf{A}^{-2}),$ (73)
-We omit the proof since it is a straightforward application of Theorem 5.3 to the moment bound Lemma B.2.
+We omit the proof since it is a straightforward application of Theorem 5.3 to the moment bound Lemma 8.2.
 We next discuss the implication for (MVI) as the randomization here raises some subtleties in its two equivalent forms, the time-uniform one and the stopped one.
@@
 Proof.
-First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Theorem 3.3.
-We apply Theorem 5.3 to obtain the first inequality.
+First, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$ due to Part 3 of Theorem $3.$ We apply Theorem 5.3 to obtain the first inequality.
 For the second inequality, we define the stopping time $\nu:=\inf\{n:\mathbf{Y}_{n}\npreceq\mathbf{A}\}$ and apply (74) to the stopping time $\nu\wedge\tau$.
@@
 We now compare these two empirically.
-We conduct the comparison with the self-normalized Example 3.8, and consider the two corresponding sequential test algorithms, Matrix and Scalar.
+We conduct the comparison with the finite-variance self-normalized Example 3.9, and consider the two corresponding sequential test algorithms, Matrix and Scalar.
 More specifically, both receive the stream of i.i.d matrices $\mathbf{X}_{1},\mathbf{X}_{2},\dots$ in $\mathcal{S}_{d}$ with mean $\mathbf{M}$ and variance upper bound $\mathbf{I}$ and sequentially test the null hypothesis $H_{0}:\mathbf{M}=0$.
@@
 For power comparison, we shall first discuss the case when only $n=1$ observation is revealed, and then move on to the case with an indefinite sequence of observations, with examples where Matrix outpowers Scalar under decreasingly restrictive conditions and finally no restriction at all.
+We shall specify the choice of the $\{\gamma_{n}\}$ sequence as needed.
 6.1 Comparison with a Single Observation
@@
 with common deterministic projection matrices $\mathbf{P}_{1},\dots,\mathbf{P}_{d}$.
-For each $k$, $\lambda_{k1},\lambda_{k2},\dots\overset{\mathrm{iid}}{\sim}\pi_{k}$ with mean $\mathbb{E}\lambda_{k1}=\lambda_{k}$ and variance $\mathbf{Var}\lambda_{k1}=\sigma_{k}^{2}\leqslant 1$.
+For each $k$, $\lambda_{k1},\lambda_{k2},\dots\overset{\mathrm{iid}}{\sim}\pi_{k}$ with mean $\mathbb{E}\lambda_{k1}=\lambda_{k}\geqslant 0$ and variance $\mathbf{Var}\lambda_{k1}=\sigma_{k}^{2}\leqslant 1$.
 Then, the variance $\mathbf{Var}(\mathbf{X}_{1})=\sigma_{1}^{2}\mathbf{P}_{1}+\dots+\sigma_{d}^{2}\mathbf{P}_{d}\preceq\mathbf{I}$.
-The matrix-valued process that Matrix computes is
-$\mathbf{Y}_{n}^{\textsf{SN}}=\sum_{k=1}^{d}Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})\mathbf{P}_{k},$ (85)
+We fix a deterministic sequence of positive scalars $\gamma_{n}=n^{-1/2}$.
+For each $k$, we define the auxiliary “component process” $\{F_{kn}\}_{n\geqslant 0}$ as the scalar finite-variance self-normalized test process with $\gamma_{n}=n^{-1/2}$ as the weight sequence when observing data $\lambda_{k1},\lambda_{k2},\dots$:
+$F_{kn}=\exp\left(-\frac{1}{3}\sum_{i=1}^{n}\gamma_{i}^{2}+\sum_{i=1}^{n}\gamma_{i}\lambda_{ki}-\frac{1}{6}\sum_{i=1}^{n}\gamma_{i}^{2}\lambda_{ki}^{2}\right).$ (84)
+The matrix-valued process that Matrix computes can then be written as
+$\mathbf{Y}_{n}^{\textsf{SN}}=\sum_{k=1}^{d}F_{kn}\mathbf{P}_{k},$ (85)
 and the scalar-valued process that Scalar computes is
-$L_{n}^{\textsf{SN}}=\sum_{k=1}^{d}Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}}).$ (86)
-In both expressions the function $Y_{n}^{\textsf{SN}}(\cdot)$ denotes the self-normalized test process for scalar observations,
-$Y_{n}^{\textsf{SN}}(x_{1},\dots,x_{n})=\exp\left(-\frac{1}{3}\sum_{i=1}^{n}\gamma_{i}^{2}+\sum_{i=1}^{n}\gamma_{i}x_{i}-\frac{1}{6}\sum_{i=1}^{n}\gamma_{i}^{2}x_{i}^{2}\right).$ (87)
-The scalar component process $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows exponentially fast if $\pi_{k}$ has mean $\lambda_{k}\neq 0$, and shrinks to 0 if $\lambda_{k}=0$; moreover, the rate in which $Y_{n}^{\textsf{SN}}(\lambda_{k1},\dots,\lambda_{kn})$ grows depends exponentially on the size $|\lambda_{k}|$.
-Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will have a low effective rank with $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})\mathbf{P}_{d}$ will exponentially outgrow the rest.
+$L_{n}^{\textsf{SN}}=\sum_{k=1}^{d}F_{kn}=\mathrm{tr}(\mathbf{Y}_{n}^{\textsf{SN}}).$ (86)
+Note that $\{F_{kn}\}$ satisfies
+$F_{kn}=F_{k(n-1)}\cdot\exp\left(\frac{\lambda_{kn}}{\sqrt{n}}-\frac{2+\lambda_{kn}^{2}}{6n}\right).$ (87)
+For large $n$, the first $n^{-1/2}$ term inside the $\exp(\cdot)$ above dominates the second $n^{-1}$ term.
+We see that each scalar component process $\{F_{kn}\}$ eventually grows exponentially fast if $\lambda_{k}=\mathbb{E}\lambda_{kn}>0$, and shrinks to 0 if $\lambda_{k}=0$; moreover, the exponential growth rate of $\{F_{kn}\}$ depends polynomially on the effect size $\lambda_{k}$.
+Consequently, as long as there is any mean disparity between the component distributions $\pi_{1},\dots,\pi_{d}$, say $\lambda_{1},\dots,\lambda_{d-1}<\lambda_{d}$, in the long run the process $\mathbf{Y}_{n}^{\textsf{SN}}$ will have a low effective rank with $\operatorname{r}(\mathbf{Y}_{n}^{\mathsf{SN}})={\mathrm{tr}(\mathbf{Y}_{n}^{\mathsf{SN}})}/{\lambda_{\max}(\mathbf{Y}_{n}^{\mathsf{SN}})}\to 1$, as the component $F_{dn}\mathbf{P}_{d}$ will exponentially outgrow the rest $F_{1n}\mathbf{P}_{1},\dots,F_{(d-1)n}\mathbf{P}_{d-1}$.
 In this case if Matrix correctly anticipates the largest component $\mathbf{P}_{d}$, by picking again the threshold matrix
 $\mathbf{A}=\varepsilon^{-1}(d-1)\mathbf{P}_{1}+\dots\varepsilon^{-1}(d-1)\mathbf{P}_{d-1}+(\alpha-\varepsilon)^{-1}\mathbf{P}_{d}$ (88)
-with small $\varepsilon$ it will outperform Scalar: it only takes $Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})$ to surpass $(\alpha-\varepsilon)^{-1}\approx 1/\alpha$ for Matrix to reject, whereas it takes $L_{n}^{\textsf{SN}}\approx Y_{n}^{\textsf{SN}}(\lambda_{d1},\dots,\lambda_{dn})$ to surpass $d/\alpha$ for Scalar to reject.
+with small $\varepsilon$ it will outperform Scalar: it only takes $F_{dn}$ to surpass $(\alpha-\varepsilon)^{-1}\approx 1/\alpha$ for Matrix to reject, whereas it takes $L_{n}^{\textsf{SN}}\approx F_{dn}$ to surpass $d/\alpha$ for Scalar to reject.
 Similar reasoning applies as long as Matrix knows a subset of the $d$ components with larger means.
+We finally remark that we can lift the assumption $\lambda_{k}\geqslant 0$ here, by using the “hedging” argument outlined in Remark 3.8.
+Let $\mathbf{Y}_{n}^{+}$ and $\mathbf{Y}_{n}^{-}$ be the processes Matrix would use with $\gamma_{n}=n^{-1/2}$ and $\gamma_{n}=-n^{-1/2}$ respectively.
+Then, if $|\lambda_{1}|,\dots,|\lambda_{d-1}|<|\lambda_{d}|$, the spectrum of $(\mathbf{Y}_{n}^{+}+\mathbf{Y}_{n}^{-})/2$ will concentrate on the $\mathbf{P}_{d}$ direction, too.
 We experiment with the following setup.
@@
 While this indicates that
-smart Matrix $>$ naive Matrix $\approx$ Scalar
+smart Matrix >> naive Matrix $\approx$ Scalar
 in power, we now come to the striking (in light of Proposition 6.1) phenomenon that with the commutativity assumption removed, it is actually possible that
-smart Matrix $>$ naive Matrix $>$ Scalar.
+smart Matrix >> naive Matrix >> Scalar.
 In the following example that involves non-commuting sequential observations, we consider the application scenario of covariance estimation [Ke et al., 2019].
@@
 Comparison between the trace of the matrix-valued test process, and the scalar-valued test process under the same alternative distribution.
-7 Conclusion
-In this paper, we study positive semidefinite matrix supermartingales and backward submartingales.
-In particular, we present several maximal inequalities that generalize existing results in the well-known scalar case.
-We demonstrate the statistical applications of these results by noting that for data and hypotheses that inherently take matrix values, our “matrix-native” approach is beneficial and can outperform previous methods that often require “scalarization”.
-In particular, sequential testing with matrix-valued test supermartingales and e-processes allows a freely chosen rejection threshold matrix $\mathbf{A}$ subject to the constraint $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$, one that can be chosen according to the prior knowledge to boost power without sacrificing the type 1 error guarantee.
-On the other hand, taking the threshold matrix $\mathbf{A}$ to be a multiple of the identity matrix, a degenerate choice that it appears, can still lead to a more powerful test in the large sample regime compared to scalarization.
-This is an empirical finding of our paper that calls for further understanding.
-Our study shall lead to more powerful and flexible sequential hypothesis testing procedures for matrix inference problems including covariance and graph testing.
-Acknowledgement
-The authors thank Joel A. Tropp, Lester Mackey, Tudor Manole, and Reihaneh Malekian for suggestions and discussions, and acknowledge support from NSF grant DMS-2310718.
-Appendix A On the Randomized Matrix Markov Inequality
+7 On the Randomized Matrix Markov Inequality
 Instead of randomizing the bound with the isotropic diagonal matrix $U\mathbf{I}$ where $U$ is a uniform or super-uniform random scalar, let us seek an inherently matrix-born definition of uniformity and super-uniformity.
 Scalar super-uniformity has two viable extensions to matrices, one in terms of the trace and the other of the spectral norm.
-Definition A.1.
+Definition 7.1.
 We call an $\mathcal{S}_{d}^{+}$-valued random matrix $\mathbf{U}$ spectrally super-uniform if $\mathbb{P}(\mathbf{U}\nsucceq\mathbf{Y})\leqslant\lambda_{\max}(\mathbf{Y})$ for all $\mathbf{Y}\in\mathcal{S}_{d}^{+}$, and $\mathcal{Y}$-spectrally uniform if, further, equality holds for all $\mathbf{Y}\in\mathcal{Y}\subseteq\mathcal{S}_{d}^{+}$.
@@
 Having defined trace super-uniformity, we can now state a matrix extension of the uniformly randomized Markov’s inequality (71).
-Theorem A.2 (Generalized Uniformly Randomized Matrix Markov Inequality).
+Theorem 7.2 (Generalized Uniformly Randomized Matrix Markov Inequality).
 Let $\mathbf{X}$ be a random matrix taking values in $\mathcal{X}\subseteq\mathcal{S}_{d}^{+}$, and $\mathbf{U}$ a trace super-uniform random matrix taking values in $\mathcal{S}_{d}^{+}$ independent from $\mathbf{X}$.
@@
 It may be interesting to fully characterize the set of all trace super-uniform matrices, but we leave it as future work.
-Appendix B Omitted Proofs
-B.1 Dominated Integrability Lemma
+8 Omitted Proofs
+8.1 Dominated Integrability Lemma
 It is well-known that if $X$ and $Y$ are scalar random variables such that $c\leqslant X\leqslant Y$ almost surely for some constant $c$ and that $\mathbb{E}|Y|<\infty$, it follows that $\mathbb{E}|X|<\infty$ as well, and $\mathbb{E}X\leqslant\mathbb{E}Y$.
@@
 Let us prove its symmetric matrix extension for the sake of self-containedness.
-Lemma B.1 (Dominated Integrability).
+Lemma 8.1 (Dominated Integrability).
 Let $\mathbf{X}$ and $\mathbf{Y}$ be $\mathcal{S}_{d}^{[c,\infty)}$-valued random matrices for some $c\in\mathbb{R}$ such that $\mathbf{X}\preceq\mathbf{Y}$ almost surely.
@@
 ∎
-B.2 Proofs for Forward Supermartingales
+8.2 Proofs for Forward Supermartingales
 The following transfer rule [Tropp, 2012, Equation 2.2] shall commonly be used: suppose $I\subseteq\mathbb{R}$ and $f,g:I\to\mathbb{R}$ satisfies $f(x)\leqslant g(x)$, then, since the function $g-f:I\to\mathbb{R}_{\geqslant 0}$ identifies canonically with a matrix-to-matrix function $g-f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{+}$, it follows that $f(\mathbf{X})\preceq g(\mathbf{X})$ for any $\mathbf{X}\in\mathcal{S}_{d}^{I}$.
 Proof of Lemma 3.1.
-Note that
+(1) Note that $\mathbf{A}\succeq 0$ if and only if $\mathbf{v}^{\mathsf{T}}\mathbf{A}\mathbf{v}\geqslant 0$ for all $\mathbf{v}\in\mathbb{Q}^{d}$ due to the continuity of the quadratic form. We have,
 $\{\mathbf{Y}_{n}\}$ is a supermartingale (99)
-$\iff$ For all $n$, $\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq 0$ (100)
-$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}(\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1}))\mathbf{v}\geqslant 0$ (101)
-$\iff$ For all $n$ and $\mathbf{v}$, $\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n-1}\mathbf{v}-\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}|\mathcal{F}_{n-1})\geqslant 0$ (102)
-[…]
+$\iff$ $\mathbb{P}\left(\bigcap_{n}\{\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\succeq 0\}\right)=1$ (100)
+$\iff$ $\mathbb{P}\left(\bigcap_{n}\bigcap_{v\in\mathbb{Q}^{d}}\{\mathbf{v}^{\mathsf{T}}(\mathbf{Y}_{n-1}-\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1}))\mathbf{v}\geqslant 0\}\right)=1$ (101)
+$\iff$ $\mathbb{P}\left(\bigcap_{n}\bigcap_{v\in\mathbb{Q}^{d}}\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n-1}\mathbf{v}-\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}|\mathcal{F}_{n-1})\geqslant 0\}\right)=1$ (102)
+$\iff$ For all $\mathbf{v}\in\mathbb{Q}^{d}$, $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale (103)
+$\iff$ For all $\mathbf{v}\in\mathbb{R}^{d}$, $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale. (104)
+(2) Since $\mathbf{A}_{n}=\sum_{i=1}^{n}(\mathbb{E}(\mathbf{Y}_{i}|\mathcal{F}_{i-1})-\mathbf{Y}_{i-1})$,
+$\{\mathbf{Y}_{n}\}$ is a supermartingale (105)
+$\iff$ For all $n$, $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})-\mathbf{Y}_{n-1}\preceq 0$ (106)
+$\iff$ $\text{For all $n$, $\mathbf{A}_{n}-\mathbf{A}_{n-1}\preceq 0$.}\qed$ (107)
 Proof of Example 3.9.
@@
 $\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\preceq 1+\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})+\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/3.$ (108)
-This allows us to use Lemma B.1,
+This allows us to use Lemma 8.1,
 $\mathbb{E}\left(\exp(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/6)\mid\mathcal{F}_{n-1}\right)$ (109)
 […]
-=\ = $1+\mathbb{E}\left(\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/3\mid\mathcal{F}_{n-1}\right)\preceq 1+\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (107)
+$=\$$1+\mathbb{E}\left(\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}/3\mid\mathcal{F}_{n-1}\right)\preceq 1+\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\preceq\exp\left(\frac{\gamma_{n}^{2}}{3}\mathbf{V}_{n}\right),$ (111)
 where in the last step we use the transfer rule again with $1+x\leqslant\exp(x)$ for all $x\in\mathbb{R}$. ∎
@@
 Let us replicate the key steps of the proof of Lemma 3(d) in Howard et al.
-[2020], which is also a self-normalization argument.
+[2020], another self-normalization-style bound derived this time by the classical symmetrization argument.
 Due to the symmetry of $\mathbf{X}_{n}-\mathbf{M}_{n}$, we introduce a Rademacher random variable $\epsilon$ independent from $\mathcal{F}_{n}$,
@@
 $=$ $\mathbb{E}\left\{\frac{\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)+\exp\left(-\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)}{2}\middle|\mathcal{F}_{n-1}\right\}.$ (115)
-Note that the function $f(x)=\frac{\exp(x-x^{2}/2)+\exp(-x-x^{2}/2)}{2}$ satisfies $f(x)\leqslant 1$ for all $x\in\mathbb{R}$, we see from the transfer rule and Lemma B.1 that
+Note that the function $f(x)=\frac{\exp(x-x^{2}/2)+\exp(-x-x^{2}/2)}{2}$ satisfies $f(x)\leqslant 1$ for all $x\in\mathbb{R}$, we see from the transfer rule and Lemma 8.1 that
 $\mathbb{E}\left(\exp\left(\gamma_{n}(\mathbf{X}_{n}-\mathbf{M}_{n})-\frac{\gamma_{n}^{2}(\mathbf{X}_{n}-\mathbf{M}_{n})^{2}}{2}\right)\middle|\mathcal{F}_{n-1}\right)=\mathbb{E}\left(f(\mathbf{X}_{n}-\mathbf{M}_{n})\middle|\mathcal{F}_{n-1}\right)\preceq\mathbf{I}.\qed$ (116)
@@
 This concludes the proof. ∎
-B.3 Proofs for Backward Submartingales
-Lemma B.2 (Contraction of the Matrix Variance).
+8.3 Proofs for Backward Submartingales
+Lemma 8.2 (Contraction of the Matrix Variance).
 Let $\mathbf{X}_{1},\dots,\mathbf{X}_{n}$ be random matrices taking values in $\mathcal{S}_{d}$ with common mean matrix $\mathbb{E}\mathbf{X}_{i}=\mathbf{M}$ and variance matrix upper bound $\mathbf{Var}(\mathbf{X}_{i})\preceq\mathbf{V}$, such that $\mathbb{E}\{\mathbf{X}_{i}-\mathbf{M},\mathbf{X}_{j}-\mathbf{M}\}=0$ for any $i\neq j$.
@@
 Finally, noting that $\|\mathbf{A}\|^{p}\leqslant\mathrm{tr}((\operatorname{abs}\mathbf{A})^{p})$ concludes the proof. ∎
-B.4 Proof of Proposition 5.1
+8.4 Proof of Proposition 5.1
 Recall for convenience the definition of sets
@@
 To see that there exists an $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha)$ such that $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$, suppose $\lambda_{\min}(\mathbf{A})=d/\alpha-2\delta$ where $\delta>0$.
-Take $\mathbf{X}$ to be a matrix with the same spectrum as $\mathbf{A}$ and eigenvalues $d/\alpha-\delta,\delta/(d-1),\dots,\delta/(d-1)$.
+Take $\mathbf{X}$ to be a matrix with the same eigenvectors as $\mathbf{A}$ and eigenvalues $d/\alpha-\delta,\delta/(d-1),\dots,\delta/(d-1)$.
 ∎
@@
 First, when $\alpha^{\prime}\geqslant\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ then $\lambda_{\max}(\mathbf{X})\leqslant\mathrm{tr}(\mathbf{X})\leqslant d/\alpha^{\prime}\leqslant 1/\lambda_{\max}(\mathbf{A}^{-1})=\lambda_{\min}(\mathbf{A})$ so $\mathbf{X}\in\mathcal{C}(\mathbf{A})$.
-Second, when $\alpha^{\prime}<\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\delta=(d/\alpha^{\prime}-\lambda_{\min}(\mathbf{A}))/2$ and take $\mathbf{X}$ as a matrix with the same spectrum as $\mathbf{A}$ and eigenvalues $\lambda_{\min}(\mathbf{A})+\delta,\delta/(d-1),\dots,\delta/(d-1)$ to see that $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ but $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$.
+Second, when $\alpha^{\prime}<\frac{d}{\operatorname{r}(\mathbf{A}^{-1})}\cdot\alpha$, let $\delta=(d/\alpha^{\prime}-\lambda_{\min}(\mathbf{A}))/2$ and take $\mathbf{X}$ as a matrix with the same eigenvectors as $\mathbf{A}$ and eigenvalues $\lambda_{\min}(\mathbf{A})+\delta,\delta/(d-1),\dots,\delta/(d-1)$ to see that $\mathbf{X}\in\mathcal{C}_{\mathrm{tr}}(\alpha^{\prime})$ but $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$.
 ∎
@@
 Proof of Statement 5.
-Complete the orthonormal basis $\{\mathbf{v}_{k}\}_{1\leqslant k\leqslant d}$ where $\mathbf{v}_{1}=\mathbf{v}$ and let $\mathbf{A}=\alpha^{-1}\mathbf{v}\mathbf{v}^{\mathsf{T}}+\sum_{k=2}^{d}\delta\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{T}}$ for $\delta=\frac{\alpha^{\prime}-\alpha}{d-1}$.
+Complete the orthonormal basis $\{\mathbf{v}_{k}\}_{1\leqslant k\leqslant d}$ where $\mathbf{v}_{1}=\mathbf{v}$ and let $\mathbf{A}=\alpha^{-1}\mathbf{v}\mathbf{v}^{\mathsf{T}}+\sum_{k=2}^{d}\delta\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{T}}$ for $\delta={\frac{d-1}{\alpha^{\prime}-\alpha}}$.
 ∎
@@
 To find an $\mathbf{X}\in\mathcal{C}_{\mathbf{v}}(\alpha^{\prime})$ such that $\mathbf{X}\notin\mathcal{C}(\mathbf{A})$, take $\mathbf{X}={\alpha^{\prime}}^{-1}\mathbf{v}\mathbf{v}^{\mathsf{T}}+\sum_{k=2}^{d}L\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{T}}$ for some sufficiently large $L$.
-B.5 Proof of Lemma 5.2
+8.5 Proof of Lemma 5.2
 Proof.
@@
 Therefore,
-$\mathbb{E}(L_{n}|\mathcal{F}_{n-1})=$ $\mathbb{E}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{n-1}\mathbf{Z}_{i}-\sum_{i=1}^{n-1}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})-\mathbf{C}_{n}^{\prime}+\log\mathrm{e}^{\mathbf{Z}_{n}-\mathbf{C}_{n}}\right)\middle|\mathcal{F}_{n-1}\right)$ (129)
+$\mathbb{E}(L_{n}|\mathcal{F}_{n-1})=\$$\mathbb{E}\left(\mathrm{tr}\exp\left(\sum_{i=1}^{n-1}\mathbf{Z}_{i}-\sum_{i=1}^{n-1}(\mathbf{C}_{i}+\mathbf{C}_{i}^{\prime})-\mathbf{C}_{n}^{\prime}+\log\mathrm{e}^{\mathbf{Z}_{n}-\mathbf{C}_{n}}\right)\middle|\mathcal{F}_{n-1}\right)$ (133)
 (Jensen’s inequality)
@@
 concluding the proof. ∎
-Appendix C Miscellaneous Technical Results
-C.1 Matrix MGF Bounds
+9 Miscellaneous Technical Results
+9.1 Matrix MGF Bounds
 In Table 2, we state several results in the random concentration literature on controlling the matrix MGF $\mathbf{G}(\gamma)=\mathbb{E}(\mathrm{e}^{\gamma(\mathbf{X}_{1}-\mathbb{E}\mathbf{X}_{1})})$ for a random matrix $\mathbf{X}_{1}$ satisfying certain conditions.
@@
 Name Condition $\mathbf{G}(\gamma)$
-Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $=\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$
+Uni-Gaussian $\mathbf{X}_{1}-\mathbf{M}=G_{1}\mathbf{C},G_{1}\sim\mathcal{N}_{0,1}$ $\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$
 Rademacher $\mathbf{X}_{1}-\mathbf{M}=R_{1}\mathbf{C},R_{1}\sim\operatorname{rad}_{1/2}$ $\preceq\mathrm{e}^{\gamma^{2}\mathbf{C}^{2}/2}$
@@
 The uni-Gaussian MGF bound is proved by Tropp [2015, Lemma 4.6.2] via a Taylor expansion argument.
-For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a scalar inequality similar to the Hoeffding’s inequality plus the transfer rule (see Section B.2).
+For Rademacher, the bound $\mathbf{G}(\gamma)\preceq\exp(\gamma^{2}\mathbf{C}^{2}/2)$ is proved by Tropp [2015, Lemma 4.6.3] via a scalar inequality similar to the Hoeffding’s inequality plus the transfer rule (see Section 8.2).
 Bennett [Tropp, 2012, Lemma 6.7] and Bernstein [Tropp, 2012, Lemma 6.8] follow from a similar reasoning applied to the respective series expansion.
 ∎
-C.2 A Matrix Doob’s $L^{p}$ Inequality
+9.2 A Matrix Doob’s $L^{p}$ Inequality
 Applying an operator-monotone and concave function to a matrix supermartingale, we get another matrix supermartingale. In particular,
-Lemma C.1.
+Lemma 9.1.
 Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued supermartingale on $\{\mathcal{F}_{n}\}$ and $p\in[1,\infty)$.
@@
 For example, denoting by $\|\mathbf{X}\|_{p}$ the Schatten $p$-norm of $\mathbf{X}\in\mathcal{S}_{d}$; that is, $\|\mathbf{X}\|_{p}=(\mathrm{tr}((\operatorname{abs}\mathbf{X})^{p}))^{1/p}$.
-Lemma C.2.
+Lemma 9.2.
 Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p\geqslant 1$. Then, $\{\|\mathbf{M}_{n}\|^{p}_{p}\}$ is a submartingale.
@@
 Theorem 11.2 in Klenke [2013]).
-Proposition C.3 (Doob-type Nuclear $L^{p}$ Inequality).
+Proposition 9.3 (Doob-type Nuclear $L^{p}$ Inequality).
 Let $\{\mathbf{M}_{n}\}$ be an $\mathcal{S}_{d}^{+}$-valued submartingale or $\mathcal{S}_{d}$-valued martingale, and $p>1$. Then,
-$\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{1}^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|^{p}_{1}=\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\|\mathbf{M}_{N}\|\cdot\operatorname{r}(\mathbf{M}_{N}))^{p}.$ (140)
+$\scriptsize\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|^{p}\right)\leqslant\mathbb{E}\left(\sup_{0\leqslant n\leqslant N}\|\mathbf{M}_{n}\|_{1}^{p}\right)\leqslant\left(\frac{p}{p-1}\right)^{p}\mathbb{E}\|\mathbf{M}_{N}\|^{p}_{1}=\left(\frac{p}{p-1}\right)^{p}\mathbb{E}(\|\mathbf{M}_{N}\|\cdot\operatorname{r}(\mathbf{M}_{N}))^{p}.$ (144)
 Proof.
 The second inequality is directly from applying Doob’s $L^{p}$ inequality on the nonnegative submartingale $\{\|\mathbf{M}_{n}\|_{1}\}$.
-The first $\leqslant$ and the final $=$ follow from straightward relations between the spectral and nuclear norms, and the effective rank.
+The first $\leqslant$ and the final == follow from straightward relations between the spectral and nuclear norms, and the effective rank.
 ∎
-C.3 Non-symmetric, Non-square Matrices
+9.3 Non-symmetric, Non-square Matrices
 We finally remark on the extension of our methods to non-symmetric and possibly non-square matrices.
@@
 For example, the self-normalized matrix supermartingale (Example 3.9) can be generalized as follows.
-Example C.4.
-Let $\{\mathbf{X}_{n}\}$ be $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional covariance upper bounds $\mathbf{Cov}(\mathbf{X}_{n}^{\mathsf{T}}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{1n}$, $\mathbf{Cov}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{2n}$ that are predictable.
+Example 9.4.
+Let $\{\mathbf{X}_{n}\}$ be $\mathcal{M}_{d_{1}\times d_{2}}(\mathbb{R})$-valued random matrices adapted to $\{\mathcal{F}_{n}\}$ with conditional means $\mathbb{E}(\mathbf{X}_{n}|\mathcal{F}_{n-1})=\mathbf{M}_{n}$ and conditional covariance upper bounds
+$\mathbf{Cov}(\mathbf{X}_{n}^{\mathsf{T}}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{1n}$, $\mathbf{Cov}(\mathbf{X}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{V}_{2n}$ that are predictable.
 For any predictable positive scalars $\{\gamma_{n}\}$, let
@@
 is an $\mathcal{S}_{d_{1}+d_{2}}^{+}$-valued supermartingale.
+10 Conclusion
+In this paper, we study positive semidefinite matrix supermartingales and backward submartingales.
+In particular, we present several maximal inequalities that generalize existing results in the well-known scalar case.
+We demonstrate the statistical applications of these results by noting that for data and hypotheses that inherently take matrix values, our “matrix-native” approach is beneficial and can outperform previous methods that often require “scalarization”.
+In particular, sequential testing with matrix-valued test supermartingales and e-processes allows a freely chosen rejection threshold matrix $\mathbf{A}$ subject to the constraint $\mathrm{tr}(\mathbf{A}^{-1})=\alpha$, one that can be chosen according to the prior knowledge to boost power without sacrificing the type 1 error guarantee.
+On the other hand, taking the threshold matrix $\mathbf{A}$ to be a multiple of the identity matrix, a degenerate choice that it appears, can still lead to a more powerful test in the large sample regime compared to scalarization.
+This is an empirical finding of our paper that calls for further understanding.
+Our study shall lead to more powerful and flexible sequential hypothesis testing procedures for matrix inference problems including covariance and graph testing.
+Acknowledgement
+The authors thank Joel A. Tropp, Lester Mackey, Tudor Manole, and Reihaneh Malekian for suggestions and discussions, and acknowledge support from NSF grant DMS-2310718.